### v1

# Value Gradients with Action Adaptive Search Trees in Continuous (PO)MDPs

Abstract
Solving Partially Observable Markov Decision Processes (POMDPs) in continuous state, action and observation spaces is key for autonomous planning in many real-world mobility and robotics applications. Current approaches are mostly sample based, and cannot hope to reach near-optimal solutions in reasonable time. We propose two complementary theoretical contributions. First, we formulate a novel Multiple Importance Sampling (MIS) tree for value estimation, that allows to share value information between sibling action branches. The novel MIS tree supports action updates during search time, such as gradient-based updates. Second, we propose a novel methodology to compute value gradients with online sampling based on transition likelihoods. It is applicable to MDPs, and we extend it to POMDPs via particle beliefs with the application of the propagated belief trick. The gradient estimator is computed in practice using the MIS tree with efficient Monte Carlo sampling. These two parts are combined into a new planning algorithm Action Gradient Monte Carlo Tree Search (AGMCTS). We demonstrate in a simulated environment its applicability, advantages over continuous online POMDP solvers that rely solely on sampling, and we discuss further implications.
1 Introduction
Figure 1: Action adaptive MIS tree with two levels of actions and children state nodes. Usual states for MDPs are replaced by belief states for POMDPs. The actions in the action nodes are updated as the search progresses, indicated with a darker shade, and new states are sampled from the newest actions, denoted in the same color. The MCTS selection path, highlighted in blue, chooses at action nodes the latest action, and then proceeds at random to a child state.
The Markov Decision Process (MDP) and the Partially Observable Markov Decision Process (POMDP) are flexible mathematical frameworks for modeling decision-making under uncertainty. Computing policies for (PO)MDPs that are in continuous in states, actions and observations, has mostly focused on global search algorithms based on POMCP [?], like POMCPOW and PFT-DPW [?], that explore the action space only by sampling. Several algorithms have been suggested for mixed continuous settings, in which not all the state, action and observation spaces are continuous. For continuous state and action spaces, GPS-ABT [?] and LCEOPT [?] use generalized pattern search and cross-entropy optimization respectively, both of which are gradient-free optimization methods. For continuous state and observation spaces, AdaOPS [?] suggested belief packing to combine values for similar beliefs, and LABECOP [?] introduced lazy belief extraction to increase belief diversity.
The first example of an algorithm with performance guarantees for continuous POMDPs was VOWSS [?], that combined Voronoi Optimistic Optimization [?] for action sampling with Weighted Sparse Sampling [?]. The authors also suggested a practical version of the algorithm, VOMCPOW, that combines MCTS with Double Progressive Widening (DPW). [?] presented hierarchical Voronoi-based action sampling, showing improved performance at action dimensionality of 6 to 12.
Our research is motivated by the idea that in many continuous settings, gradient-based optimization is often much faster to converge than gradient-free optimization. It might be prohibitively expensive to sample enough actions to statistically find a near-optimal one. In our approach we consider non-parametric beliefs, making it applicable to a wide range of problems, specifically not restricted by Gaussian or other parametric assumptions. Our methodology consists of two parts. First, we consider how to update actions during search time, without losing information of previous action-value estimates. Second, we derive value gradients w.r.t. actions, for the purpose of computing action gradient updates. We hypothesize that incorporating gradient optimization into continuous (PO)MDP planning will explore the action space more efficiently, especially at high dimensional settings.
A similar idea has been recently explored in the algorithm VG-UCT by [?], where the local gradient-based optimization is combined with the global-search of sampled-based MCTS, but with several drawbacks compared to our approach. Their gradient computation approach is computationally expensive due to being based on finite differences and being recursive. They assume a deterministic transition model given a noise parameter, which is problematic for deriving gradients in POMDPs. Lastly, they do not address the distribution shift that occurs from action updates, which affect the action-value estimates.
Our approach requires access to the state transition model’s likelihood. This is in contrast to previous online (PO)MDP solvers and model-free approaches that consider only use of a generative model. Yet, recent advances in physical simulators have included scalable differentiable physics simulators [??], and GPU-parallelization that have allowed generating accurate simulations with speedups in the scale of orders of magnitudes [??]. We foresee that this will allow to compute the differentiable physical transition models more easily in the near future.
1.1 Additional Related Work
Information sharing in discrete MDPs with repeating states was considered in [?]. In POMDPs, previous knowledge utilization via Multiple Importance Sampling (MIS) [?] was introduced in [?]. We share information only between sibling action branches, which eliminates a costly search through a belief database. MIS in POMDP planning for sample reuse was also considered in [?], however in the context of calculating observation likelihoods.
KR-UCT [?] use kernel regression to estimate the value of actions by sharing information between sibling action branches, and for the proposal of new actions. They consider MDPs only, and do not consider updating existing actions. In similar vein, Bayesian optimization for information sharing between different actions was explored in kCBTS [?] and BOMCP [?]. These methods are purely heuristic, and the information sharing is used only for the sampling of new actions, whereas we consider updating existing actions.
In [?], BetaZero was suggested to share information by learning offline policy and value networks over parametric beliefs. However, they consider discrete action spaces, they do not share information between action branches online, nor do they handle modeling inaccuracy.
Our gradient results can be seen as similar to the policy gradient theorem [?] or to the deterministic policy gradient theorem [?], where we calculate gradients w.r.t. actions, rather than policy parameters.
1.2 Contributions
Our work is the first to consider an action-gradient of the value function based on transition likelihoods in continuous (PO)MDPs. Specifically, our main contributions are:
  1. 1.
    We formulate a novel search tree with recursive MIS estimators, for updating actions with consistent value estimates in continuous (PO)MDPs. We show efficient update equations for the self-normalized MIS heuristic.
  2. 2.
    We provide a novel derivation for action gradients of the value function in (PO)MDPs based on transition likelihoods. This gradient computation can be computed at a single action branch at a time, unlike previous MDP value gradients. The extension to POMDPs makes this the first derivation of POMDP value gradients with belief-dependent policies to the best of our knowledge.
  3. 3.
    We derive a novel algorithm Action Gradient MCTS (AGMCTS) that combines the MIS tree with action gradients. This algorithm is applicable to both MDPs and POMDPs. We show several algorithmic improvements and heuristics to make it more efficient.
  4. 4.
    We compare AGMCTS to PFT-DPW and POMCPOW in a simulated POMDP scenario at several dimensions and different noise levels in the domain knowledge. AGMCTS had the best overall performance on a per-simulation budget.
2 Background
MDPs.
An MDP is the tuple $\langle\mathcal{S},\mathcal{A},p_{T},r,\gamma,L,b_{0}\rangle$. $\mathcal{S},\mathcal{A}$ are the state and action spaces. The transition model $p_{T}(s^{\prime}\mid s,a)$ models the probability of arriving at state $s^{\prime}$ when taking action $a\in\mathcal{A}$ at state $s\in\mathcal{S}$. The reward function $r(s,a,s^{\prime})\in\mathbb{R}$ gives the immediate reward of transitioning from state $s$ to $s^{\prime}$ by taking action $a$. We define the reward of $(s,a)$ as the expected reward over outcome states $r(s,a)\triangleq\mathbb{E}_{s^{\prime}\mid s,a}[r(s,a,s^{\prime})]$. $\gamma\in(0,1]$ is the discount factor. The MDP starts at time 00 and terminates after $L\in\mathbb{N}\cup\{\infty\}$ steps. The initial state is drawn from the distribution $s_{0}\sim b_{0}$.
A timed policy at time $i$, denoted as $\pi_{i}$, is a mapping $\pi_{i}\colon\mathcal{S}\to\mathcal{A}$. A policy $\pi$ is a collection of timed policies from the start time until the horizon: $\pi\triangleq(\pi_{t})_{t=1}^{t=L}$. The value function of a policy $\pi$ at time $t$ is the expected sum of discounted rewards until the horizon: $V_{t}^{\pi}(s_{t})\triangleq\mathbb{E}_{s_{t+1:L}\mid s_{t},\pi}[\sum_{i=t}^{L}\gamma^{i-t}r(s_{i},\pi_{i}(s_{i}),s_{i+1})]$. The action-value function is often used as an intermediate step in the estimation of the value function, and is defined as $Q_{t}^{\pi}(s,a)\triangleq\mathbb{E}_{s^{\prime}\mid s,a}[r(s,a,s^{\prime})+\gamma V_{t+1}^{\pi}(s^{\prime})]$, where state $s^{\prime}$ is at the proceeding time of $s$. The goal of planning is to compute a policy that maximizes the value function.
POMDPs.
A POMDP adds $\langle\mathcal{O},p_{O}\rangle$ to the MDP tuple. The observation space is $\mathcal{O}$, and the observation model $p_{O}(o\mid s)$ is the conditional probability of receiving an observation $o\in\mathcal{O}$ at $s\in\mathcal{S}$. A history at time $t$, $H_{t}\triangleq(b_{0},a_{0},o_{1},\dots,a_{t-1},o_{t})$, is defined as a sequence of the starting belief, followed by actions taken and observations received until $t$. Because of the partial observability, the agent has to maintain a probability distribution of the current state given past actions and observations, known as the belief. The belief space, denoted in $\mathcal{B}$, is the space of all probability distributions over $\mathcal{S}$. The belief at time $t$ is defined as the conditional distribution of the state $b_{t}(s_{t})\triangleq\mathbb{P}(s_{t}\mid H_{t})$, given the history. We denote $H_{t}^{-}\triangleq(b_{0},a_{0},o_{1},\dots,a_{t-1})$ for the same history without the last measurement, and the propagated belief $b_{t}^{-}(s_{t})\triangleq\mathbb{P}(s_{t}\mid H_{t}^{-})$. It can be shown that optimal decision-making can be made given the belief, instead of considering the entire history [?]. Since the belief fulfills the Markov property, a POMDP is an MDP on the belief space, commonly referred to as belief MDP. Policies, the value function and the action-value function, extend from MDPs to POMDPs by equivalent definitions on $b\in\mathcal{B}$ as states. As an example, $r(b,a)\triangleq\mathbb{E}_{b^{\prime}\mid b,a}[r(b,a,b^{\prime})]$. In the general case, the belief-reward function may be any function of beliefs, f.e. entropy, and this is called $\rho$-POMDP [?]. Often, the belief reward has an underlying state-reward, in the form $r(b,a,b^{\prime})=\mathbb{E}_{s,s^{\prime}\mid b,a,b^{\prime}}[r(s,a,s^{\prime})]$. We refer to this case as a state-reward POMDP.
Importance Sampling (IS).
The Monte-Carlo (MC) estimator of an expectation, denoted with $\hat{\mathbb{E}}$ or f.e. $\hat{V}_{t}^{\pi}$, is its approximation with a finite number of samples. Importance Sampling (IS) is a technique in MC estimation where a proposal distribution $q$ is used to sample from instead of $p$. The resulting estimator for $f(x)$ is $\hat{f}=\frac{1}{N}\sum_{i=1}^{N}\frac{p(x^{i})}{q(x^{i})}f(x^{i})$. This estimator is consistent if $q(x)=0\Rightarrow p(x)=0$. With a proper choice of $q$ the variance of the estimator can be reduced, or we can calculate estimates for a target $p$ which is difficult to sample directly. Commonly the sampled weights $w_{i}=p(x^{i})/q(x^{i})$ are normalized to sum to $1$, giving the self-normalized IS estimator [?, 1.3.2].
Particle Beliefs.
Often in continuous settings we cannot maintain an exact belief. Hence, particle filters are used instead to represent a belief non parametrically [?]. We denote particle beliefs with $\bar{b}$. The particle belief of $J$ particles is represented by a set of state-weight pairs $\bar{b}=\{(s^{j},w^{j})\}_{j=1}^{J}$, and is defined as the discrete distribution $\mathbb{P}(s\mid\bar{b})={\sum_{j=1}^{J}{w^{j}\cdotp\delta(s-s^{j})}}/{\sum_{j=1}^{J}{w^{j}}}$. The basic particle filtering algorithm, Sequential Importance Sampling (SIS) [?, 1.3.2], recursively updates and normalizes the importance weights of the particles based on the observation likelihood. Sequential Importance Resampling (SIR) adds resampling of state particles after the weight updates, to avoid weight degeneracy [?]. The particle belief MDP (PB-MDP) was considered in [?], where the belief space is explicitly the space of ordered particle beliefs $\bar{b}$ of $J$ particles. An ordered particle belief is the ordered set $\bar{b}=((s^{j},w^{j}))_{j=1}^{J}$.
Multiple Importance Sampling (MIS).
MIS [?] is a method to combine the advantages of several sampling methods. It is defined as $F=\sum_{i=1}^{n}\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}w_{i}(x^{i,j})\frac{f(x^{i,j})}{q_{i}(x^{i,j})}$, which is a weighted average of $n$ IS estimators. If the weights $w_{i}$ satisfy $\sum_{i=1}^{n}w_{i}(x)=1$ whenever $f(x)\neq 0$, and $w_{i}(x)=0$ whenever $q_{i}(x)=0$, then $F$ is unbiased. The choice of $w_{i}$ may depend on the samples $x^{i,j}$. The weighting strategy $w_{i}(x)=\frac{n_{i}q_{i}(x)}{\sum_{k}n_{k}q_{k}(x)}$, named the balance heuristic, has its variance provably bounded from the optimal weighting strategy.
Monte Carlo Tree Search (MCTS).
MCTS is an algorithm used to quickly explore large state spaces [?]. It iteratively repeats four steps to build a search tree that approximates the action-values, using a best-first strategy: (1) Selection Starting from the root node, descend recursively until a node to which children can be added is found; (2) Expansion A new node is added as a child, according to action expansion strategy; (3) Simulation A simulation is run from the new node according to the rollout (default) policy; (4) Backpropagation The simulation result is backpropagated through the selected nodes to update the action-values. Often UCT is used at selection to balance between exploration of new actions and exploitation of promising ones [?]. Double Progressive Widening (DPW) [?] is a technique to limit the branching factor from being infinite in continuous settings. The number of children of a node is artificially limited to $kN^{\alpha}$ for $N$ visitations, for hyperparameters $k$ and $\alpha$. The progressive widening is double due to being applied both to the branching of actions, and branching of states.
Stochastic Gradient Optimization (SGO).
SGO is a widely used technique for efficient optimization of differentiable functions [?]. Gradient optimizers tremendously rose in popularity to their applications in deep neural networks. Adam [?] is a widely used algorithm, with heuristically adaptive step size based on first and second moment estimates of the gradient. Adam has been shown to perform well with noisy gradient estimates, and requires minimal tuning to different problems.
3 Methodology
3.1 Action-Values with Importance Sampling
Figure 2: Recursive definitions of the MIS tree, for the particle belief MDP (PB-MDP) case, in which states are particle beliefs. State nodes are circles and action nodes are squares. The state $\bar{b}$ has a value estimate $\hat{V}(\bar{b})$ and visitation count $n(\bar{b})$. The action node $(\bar{b},a)$ has an immediate reward estimate $\hat{r}(\bar{b},a)$, a future value estimate $\hat{V}_{f}(\bar{b},a)$, and visitation count $n(\bar{b},a)$. The action-value estimate is defined via $\hat{Q}(\bar{b},a)\triangleq\hat{r}(\bar{b},a)+\gamma\hat{V}_{f}(\bar{b},a)$. During action update we update $(\bar{b},a)$ to $(\bar{b},a^{\prime})$. We highlight that in the PB-MDP, we also store the propagated belief $\bar{b}^{-}$. The child state node has all properties as $\bar{b}$, and stores the proposal action $a_{prop}(\bar{b}^{-})$. During action update (11) we do $\omega^{\prime}={p_{T}(\bar{b}^{-}\mid\bar{b},a^{\prime})}({p_{T}(\bar{b}^{-}\mid\bar{b},a_{prop}(\bar{b}^{-}))})^{-1}$. In this example, $a_{prop}(\bar{b}^{-})=a$.
In an MDP setting, when updating action $a$ to $a^{\prime}$, we would like to estimate the new action-value $Q_{t}^{\pi}(s,a^{\prime})$ by reusing the information that was used for estimating $Q_{t}^{\pi}(s,a)$. Otherwise, updating actions would require us to theoretically recompute the entire subtree, which would render the algorithm very inefficient. Since this process might occur several times for each action branch, in the general case each child state may have been generated from a different proposal action. Hence, we have to account for an MIS scenario in order to estimate $Q_{t}^{\pi}(s,a^{\prime})$.
Assuming that $p_{T}(s^{\prime}\mid s,a^{\prime})$ is a non-strict subset of the support of $p_{T}(s^{\prime}\mid s,a)$, we can formally write
$\textstyle Q_{t}^{\pi}(s,a^{\prime})=\mathbb{E}_{s^{\prime}\mid s,a}[\frac{p_{T}(s^{\prime}\mid s,a^{\prime})}{p_{T}(s^{\prime}\mid s,a)}(r(s,a^{\prime},s^{\prime})+\gamma V_{t+1}^{\pi}(s^{\prime}))].$ (1)
For POMDPs, the posterior belief transition likelihood is generally intractable. However, by using the law of total expectation, we introduce a dependence of the posterior belief $b^{\prime}$ on the propagated $b^{-}$. This makes only $b^{-}$ directly dependent on $a$, which allows us to introduce an importance ratio over the propagated belief rather than posterior:
$\textstyle Q_{t}^{\pi}(b,a^{\prime})=\mathbb{E}_{b^{-}\mid b,a}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}\textstyle[\frac{p_{T}(b^{-}\mid b,a^{\prime})}{p_{T}(b^{-}\mid b,a)}$
$(r(b,a^{\prime},b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime}))].$ (2)
This is known as the propagated belief trick [?]. We can see that (1) and (2) are analogous, except for the importance ratio being between propagated beliefs in the POMDP case. The propagated belief likelihood is rather simple to calculate for appropriate choice of the belief space.
Lemma 1.
Let $\bar{b}^{-}=((s^{-,j},w^{j}))_{j=1}^{J}$ be generated from $\bar{b}=((s^{j},w^{j}))_{j=1}^{J}$ by SIS or SIR. Hence, the $j$th state particle was generated via $s^{-,j}\sim p_{T}(\cdot\mid s^{j},a)$, and:
$\textstyle p_{T}(\bar{b}^{-}\mid\bar{b},a)=\prod_{i=1}^{J}p_{T}(s^{-,j}\mid s^{j},a).$ (3)
We refer to the appendix for all proofs
3.2 Action Adaptive MIS Trees
We now present our definitions of the novel MIS tree, a recursive tree structure with MIS value estimates, that allows sharing value information between sibling actions. Similarly to belief trees, the tree has levels alternating with states and action nodes, starting from the root state node. Yet in contrast, the MIS tree supports replacing an action node with a new action, and updating the value estimates of the new action and its parent state node based on previous samples. We later introduce action updates via gradient-based optimization.
For the following discussion, we present notations for recursively defining estimators at a state node $s$ in the MIS tree, or some action node $(s,a)$. The generalization to POMDPs is done by taking ordered particle belief states $\bar{b}$, and the following definitions are analogous, illustrated in Figure 2.
Denote the children actions of $s$ as $C(s)=\{a^{i}\}_{i=1}^{\lvert{C(s)}\rvert}$, and the children states of $(s,a)$ as $C(s,a)=\{s^{\prime,i}\}_{i=1}^{M}$, where $M\triangleq\lvert{C(s,a)}\rvert$. The visitation counts of state and action nodes are denoted $n(s)$ and $n(s,a)$ respectively. We denote the proposal action $a_{prop}(s)$, the action that was used to sample $s$, as possibly $a_{prop}(s^{\prime})\neq a$ for $s^{\prime}\in C(s,a)$. We assign each action node $(s,a)$ an action-value estimate $\hat{Q}(s,a)$ that estimates $Q_{t}^{\pi}(s,a)$, assuming $s$ is a state at time $t$. Compared to traditional MCTS algorithms, we also estimate the value function at each state node. We define it as a weighted mean:
$\textstyle\hat{V}(s)\triangleq\sum_{C(s)}\frac{n(s,a^{i})}{n(s)}\hat{Q}(s,a^{i}).$ (4)
The following two relationships between visitation counts are maintained in MCTS:
$\textstyle n(s)=\sum_{C(s)}n(s,a^{i}),\;\textstyle n(s,a)=\sum_{C(s,a)}n(s^{\prime,i})_{+1},$ (5)
where $n(s^{\prime,i})_{+1}\triangleq n(s^{\prime,i})+1$. The asymmetry is the result of initializing state nodes to $n(s)=0$ when assigned a rollout value, whereas action nodes are initialized to $n(s,a)=1$. This allows value estimates to be updated completely when they are visited again, and for their parent’s action-value estimate to remember the initial rollout estimate.
We split $\hat{Q}(s,a)$ into two estimators: (1) Immediate reward estimate $\hat{r}(s,a)$, for $\mathbb{E}_{s^{\prime}\mid s,a}[r(s,a,s^{\prime})]$; (2) Future value estimate $\hat{V}_{f}(s,a)$ for $\mathbb{E}_{s^{\prime}\mid s,a}[V^{\pi}(s^{\prime})]$ of the proceeding time, such that if $s$ is a state at time $t$, then $\hat{V}_{f}(s,a)$ estimates the expected $V_{t+1}^{\pi}(s^{\prime})$. Together they define $\hat{Q}(s,a)\triangleq\hat{r}(s,a)+\gamma\hat{V}_{f}(s,a)$.
Generally in MIS, each proposal distribution $q_{i}$ has $n_{i}$ samples. In our situation, we assume that each of $C(s,a)$ may have been sampled from a different action, has its own visitation counter and an associated single value estimate $V(s^{\prime,i})$. Therefore, we adapt the most naive unbiased weighting, that is to give equal weight to all samples, and weight the posterior states by the visitation counts. Hence, the naive MIS estimator for the future value is
$\textstyle\hat{V}_{f}(s,a)=(n(s,a))^{-1}\sum_{C(s,a)}\omega^{i}n(s^{\prime,i})_{+1}\hat{V}(s^{\prime,i}),$ (6)
where $\omega^{i}={p_{T}(s^{\prime,i}\mid s,a)}({p_{T}(s^{\prime,i}\mid s,a_{prop}(s^{\prime,i}))})^{-1}$. The naive MIS estimator for $\hat{r}(s,a)$ is identical, with $\hat{V}(s^{\prime,i})$ replaced by $r(s,a,s^{\prime,i})$.
In practice, this estimator has very bad performance. On the other hand, the balance heuristic requires $O(M)$ time to update [?], which can be quite expensive. We found a simple solution which is to self-normalize the weights. In regular IS, it is known to reduce the variance of the estimator at the expense of introducing bias, and asymptotically converges to the true estimate under weak assumptions [?, 1.3.2]. We define self-normalized MIS (SN-MIS) by
$\textstyle\eta\triangleq\sum_{C(s,a)}\omega^{i}n(s^{\prime,i})_{+1},$ (7)
$\textstyle\hat{V}_{f}(s,a)=\eta^{-1}\sum_{C(s,a)}\omega^{i}n(s^{\prime,i})_{+1}\hat{V}(s^{\prime,i}),$ (8)
and again the SN-MIS estimator for $\hat{r}(s,a)$ is defined by replacing $\hat{V}(s^{\prime,i})$ with $r(s,a,s^{\prime,i})$.
3.3 Updating MIS Trees in MCTS
We now discuss how to update the SN-MIS estimators in the steps of the MCTS routine. We highlight that for any other MIS weighting heuristic, a similar analysis would follow.
Currently, we do not address selection, and we assume the UCT criterion is not altered. Simulation is unaffected, as it is determined only by the rollout policy. Hence, we analyze the different subcases of backpropagation and expansion steps. We add a new step to MCTS named action update, where we update action node $(s,a)$ to $(s,a^{\prime})$ with the new action $a^{\prime}$, and we show its update equations. In general, action update can be applied at every point of MCTS.
For the recursive analysis, we assume backpropagated values for the children nodes are already updated, and at each step we only update an action node $(s,a)$ and its parent node $s$. The updated values are denoted with a prime, and we assume that we have access to the values before the update. The transition likelihood terms $w_{i}$ are usually the most expensive to calculate, hence it is important to cache them, and recalculate the numerator of $w_{i}$ only when necessary. The derivations of the update equations, alongside numerically stable forms with log-likelihoods are given in the appendix.
Action Backpropagation.
Let $s^{\prime,i}$ be an updated node. Hence, we updated $n(s^{\prime,i})$ to $n^{\prime}(s^{\prime,i})$ and $\hat{V}(s^{\prime,i})$ to $\hat{V}^{\prime}(s^{\prime,i})$. We update $n(s,a)$ by (5) and perform
$\textstyle\eta^{\prime}=\eta+\omega^{i}(n{{}^{\prime}}(s^{\prime,i})-n(s^{\prime,i})),$ (9)
$\textstyle\hat{V}_{f}^{\prime}(s,a)=(\eta^{\prime})^{-1}(\eta\hat{V}_{f}(s,a)+$
$\textstyle\omega^{i}(n{{}^{\prime}}(s^{\prime,i})\hat{V}^{\prime}(s^{\prime,i})-n(s^{\prime,i})\hat{V}(s^{\prime,i})).$ (10)
State Expansion.
This case is the same as action backpropagation, where the updated index is a new summand at index $i=M+1$. The same equations apply with $n(s^{\prime,i})=0$.
Action Update.
Let $a^{\prime}$ be the new action. Here, we have to calculate $\omega^{i}$ for all $s^{\prime,i}\in C(s,a)$ in $O(M)$ time. We update
$\textstyle(\omega^{i})^{\prime}=p_{T}(s^{\prime,i}\mid s,a^{\prime})(p_{T}(s^{\prime,i}\mid s,a_{prop}(s^{\prime,i})))^{-1},$ (11)
and recalculate (7) and (8) with the new ($\omega^{i})^{\prime}$, which are also $O(M)$ time operations. For the immediate reward, new rewards must be computed for the new triplets $(s,a^{\prime},s^{\prime,i})$, and the update is analogous.
Terminal State Backpropagation.
Let $s$ be a terminal node. Its value estimate is based only on rollouts. Hence, for a new rollout value $v^{\prime}$, we perform a running average update
$\textstyle\hat{V}^{\prime}(s)=\hat{V}(s)+\frac{n^{\prime}(s)-n(s)}{n^{\prime}(s)_{+1}}(v^{\prime}-\hat{V}(s)),$ (12)
where $n^{\prime}(s)$ is $n(s)$ plus the number of rollouts (usually 1).
Non-Terminal State Backpropagation.
Let $(s,a^{i})$ be an updated node. Hence, we updated $n(s,a^{i})$ to $n^{\prime}(s,a^{i})$ and $\hat{Q}(s,a^{i})$ to $\hat{Q}^{\prime}(s,a^{i})$. We update $n^{\prime}(s)$ by (5) and perform
$\textstyle\hat{V}^{\prime}(s)=(n^{\prime}(s))^{-1}(n(s)\hat{V}(s)+n^{\prime}(s,a^{i})\hat{Q}^{\prime}(s,a^{i})$
$\textstyle{}-n(s,a^{i})\hat{Q}(s,a^{i})).$ (13)
3.4 MDP and POMDP Value Gradient
In the previous subsections we’ve assumed that the updated action $a^{\prime}$ is arbitrary. We now present a method to compute an estimate of $\nabla_{a}Q(s,a)$, in order to compute a gradient-based update. Our derivation of $\nabla_{a}Q(s,a)$ is based on the gradient of the transition likelihood, and there are two main advantages to this approach. First, the non-recursive computation can be computed separately for each action branch in the MIS tree, which greatly reduces computational complexity. Second, which is our key observation, is that this derivation extends to POMDPs as well via the propagated belief trick, and has a computable form for the PB-MDP.
Theorem 1.
For well-behaved $p_{T}$, $r$, the MDP action gradient of the action-value function is given by the expectation
$\textstyle\nabla_{a^{\prime}}Q_{t}^{\pi}(s,a^{\prime})\textstyle=\mathbb{E}_{s^{\prime}\mid s,a}[\frac{p_{T}(s^{\prime}\mid s,a^{\prime})}{p_{T}(s^{\prime}\mid s,a)}(\nabla_{a^{\prime}}\log p_{T}(s^{\prime}\mid s,a^{\prime})$
$\textstyle\cdot(r(s,a^{\prime},s^{\prime})+\gamma V_{t+1}^{\pi}(s^{\prime}))+\nabla_{a^{\prime}}r(s,a^{\prime},s^{\prime}))].$ (14)
Additionally, for well-behaved $p_{T}$, $p_{O}$, $r$, the POMDP action gradient of the action-value function can be computed by replacing posterior belief likelihood terms with propagated belief likelihood terms, i.e. $b_{t+1}^{-}$ instead of $b_{t+1}$ in all $p_{T}$ terms.
The gradient is written such that states $s^{\prime}$ could have been sampled from an action $a$ different from $a^{\prime}$. Intuitively, it tends toward actions that both maximize immediate rewards, and maximize the likelihood of future states with high values. Importantly, because $s^{\prime}$ is an integration variable, the future value is not a function of $a^{\prime}$, and has no gradient w.r.t. it.
Theorem 1 is general, and applies to MDPs, state-reward POMDPs, and $\rho$-POMDPs. The following Theorem 2 modifies the gradient for state-reward POMDPs. The gradient for the immediate reward is based on single state samples rather than sampling full beliefs. It may give more accurate gradient estimates, albeit possibly more expensive to compute.
Theorem 2.
For a state-reward POMDP, the action gradient of the action-value function is given by:
$\textstyle\nabla_{a^{\prime}}Q_{t}^{\pi}(b,a)=\mathbb{E}_{s\mid b}\mathbb{E}_{s^{\prime}\mid s,a^{\prime}}[\nabla_{a^{\prime}}\log p_{T}(s^{\prime}\mid s,a^{\prime})\cdot r(s,a^{\prime},s^{\prime})$
$\textstyle+\nabla_{a^{\prime}}r(s,a^{\prime},s^{\prime})]\textstyle+\gamma\mathbb{E}_{b^{-}\mid b,a}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}[\frac{p_{T}(b^{-}\mid b,a^{\prime})}{p_{T}(b^{-}\mid b,a)}$
$\textstyle\nabla_{a^{\prime}}\log p_{T}(b^{-}\mid b,a^{\prime})\cdot V_{t+1}^{\pi}(b^{\prime})].$ (15)
For ordered particle beliefs, the gradient log of the propagated transition likelihood from Lemma 1 is given by:
$\textstyle\nabla_{a}\log p_{T}(\bar{b}^{-}\mid\bar{b},a)=\sum_{i=1}^{J}\nabla_{a}\log p_{T}(s^{-,j}\mid s^{j},a).$ (16)
3.5 Action Gradient Monte Carlo Tree Search
Algorithm 1 AGMCTS
procedure S IMULATE$(s,d)$
1:  if
I
S
T
ERMINAL$(s)$ OR $d=0$ then
2:     $v\leftarrow$
R
OLLOUT$(s)$Rollout until terminal state}
3:      U PDATE T ERMINAL($s,v$) {Eq. (12)}
4:     return $v$
5:  $a\leftarrow$
A
CTION
P
ROG
W
IDEN$(s)$
6:   $addSample\leftarrow$ A CTION O PT $(s,a,d)$
7:  if  $\lvert{C(sa)}\rvert\leq k_{o}N(sa)^{\alpha_{o}}$ OR $addSample$  then
8:     $s^{\prime},r\leftarrow G(s,a)$Replace $s^{\prime}$ by $(b^{\prime},b^{-})$ for POMDP}
9:     $C(sa)\leftarrow C(sa)\cup\left\{(s^{\prime},r)\right\}$
10:     v
R
OLLOUT
(s,d1)
[…]
R
OLLOUT
[…]
11:  else
12:     $s^{\prime},r\leftarrow$ sample uniformly from $C(sa)$
13:     v
S
IMULATE
(s,d1)
[…]
S
IMULATE
[…]
14:   U PDATE MIS ($s,a,s^{\prime},r$) {Eqs. (9), (10), (13)}
procedure
A
CTION
O
PT
(s,a,d)
A
CTION
O
PT
[…]
1:  $addSample\leftarrow$ FALSE
2:  for all $k=1,\dots,K_{opt}$ do
3:     $g_{a}^{q}=\hat{\nabla}_{a}\hat{Q}(s,a)$Eq. (19) or (17)}
4:     Initialize $a^{acc}\leftarrow a$ for $(s,a)$ if required
5:     aacc
O
PT
(s,aacc,gaq)
[…]
O
PT
[…]
{Adam/other optimizer}
6:     if $\lVert{a^{acc}-a}\rVert>T_{d_{a}}$ then
7:        $addSample\;|\mathbin{=}\;$
A
CTION
U
PDATE$(s,a,a^{acc})$Eqs. (11), (13); Return TRUE if all $\omega^{i}\leq T_{\omega}^{add}$.}
8:  return $addSample$
We introduce Action Gradient MCTS (AGMCTS), a novel online planning algorithm based on MCTS with DPW. It is applicable to POMDPs via ordered particle belief states, like PFT-DPW [?]. It is described in high-level in Algorithm 1, with novel components highlighted in blue. It combines the MIS estimators for the action-value function, with gradient-based action optimization, and updates actions accordingly. AGMCTS applies action update before expansion. This allows us to pass the information from to determine whether to expand the current action node.
Monte Carlo Gradient Estimation.
While computing (14), (15) can be done based on all children of the current action node, we employ the idea of stochastic gradient optimization to MC estimate the gradient. The first estimate is to sample $K_{O}^{\nabla}$ observation branches with weights $\omega^{i}$. It is possible to uniformly sample and weight by $\omega^{i}$, but we found it to be less effective. The MC approximation of (14) becomes
^$\textstyle\hat{\nabla}_{a}\hat{Q}(s,a)=\frac{1}{K_{O}^{\nabla}}\sum_{k=1}^{K_{O}^{\nabla}}\nabla_{a}\log p_{T}(s^{\prime,i_{k}}\mid s,a)$
$\textstyle\cdot(r(s,a,s^{\prime,i_{k}})+\gamma\hat{V}(s^{\prime,i_{k}}))+\nabla_{a}r(s,a,s^{\prime,i_{k}}).$ (17)
For POMDPs, we estimate (16) with $K_{b}^{\nabla}$ state particles via
$\textstyle\nabla_{a}\log p_{T}(\bar{b}^{-}\mid\bar{b},a)\approx\frac{J}{K_{b}^{\nabla}}\sum_{l=1}^{K_{b}^{\nabla}}\nabla_{a}\log p_{T}(s^{-,j_{l}}\mid s^{j_{l}},a),$ (18)
where indices $j_{l}$ are sampled uniformly from $[1,\dots,J]$. In state-reward POMDPs, we estimate the immediate reward separately. We generate $K_{r}^{\nabla}$ posterior states by sampling $s_{new}^{j}\sim p_{T}(\cdot\mid\bar{b},a)$. The MC estimate of (15) is then
^$\textstyle\hat{\nabla}_{a}\hat{Q}(\bar{b},a)=\frac{1}{K_{r}^{\nabla}}\big{(}\sum_{i=j}^{K_{r}^{\nabla}}\nabla_{a}\log p_{T}(s_{new}^{j}\mid s,a)\cdot$
$\textstyle r(s,a,s_{new}^{j})+\nabla_{a}r(s,a,s_{new}^{j})\big{)}+\big{(}\frac{\gamma}{K_{O}^{\nabla}}\sum_{k=1}^{K_{O}^{\nabla}}(\frac{J}{K_{b}^{\nabla}}\sum_{l=1}^{K_{b}^{\nabla}}$
$\textstyle\nabla_{a}\log p_{T}(s^{-,i_{k},j_{l}}\mid s^{i_{k},j_{l}},a))\textstyle\cdot\hat{V}(\bar{b}^{\prime,i_{k}})\big{)}.$ (19)
Gradient Optimization.
We’ve found throughout our experiments that the Adam optimizer, utilizing step size normalization, was crucial for consistent behavior due to the high variance of the gradient estimates caused by MC estimation and extremely high variance of the propagated belief transition probability. A slight step size decay was beneficial as well for convergence, allowing to choose larger Adam step sizes. We’ve fixed as a hyperparameter $K_{opt}$ the number of consecutive gradient optimization iterations before simulation. Together with the step size, this controls a trade-off between accuracy and computational complexity.
Lastly, since action updates are significantly more costly than computing the gradient estimates, we set a threshold action distance for action update. After each gradient iteration, we update an accumulator of the gradient steps $a_{t}^{acc}$. Only after passing a threshold distance $\lVert{a_{t}^{acc}-a_{t}}\rVert>T_{d_{a}}$, we update $a_{t}\leftarrow a_{t}^{acc}$. This limits the number of action updates, and acts like a convergence criterion.
Thresholds For Adding/Deleting State Nodes.
Although the SN-MIS estimator normalizes the contribution of each posterior node, we would like to ensure that we have relevant samples after several action updates. As a relevance heuristic, we delete child state $s^{\prime,i}$ from $C(s,a)$ if $\omega^{i}<T_{\omega}^{del}$ after action update. Additionally, if $\omega^{i}<T_{\omega}^{add}$ for all $s^{\prime,i}\in C(s,a)$, we force to sample a new posterior node, overriding the progressive widening limitation. This ensures that we do not waste computational resources on states with low contribution to the estimator $\hat{Q}(s,a)$. Due to the possibility of deleting posterior children, care must be taken when updating visitation counters since they do not monotonically increment anymore. For this reason equations (9), (10), (11), (12), (13) are written using previous and updated visitation counts.
4 Experiments
Numerical simulations were carried out to evaluate the performance of AGMCTS in a POMDP setting, compared to state-of-the-art algorithms POMCPOW and PFT-DPW, whose implementations were taken as the standard in the POMDPs.jl framework [?]. AGMCTS was implemented using POMDPs.jl as well. For gradient computations we used automatic differentiation with Enzyme.jl [?]. The analysis and comparison was based on the code of [?], and similarly we used rollout-first-then-random action sampling.
4.1 D-Continuous Light Dark
(a) Middle of scenario, $t=5$.
(b) End of scenario, $t=6$.
Figure 3: AGMCTS at 2-Continuous Light Dark. The agent’s current state is the red dot, the current belief particles in orange, the next state in black, and the next belief particles in gray, the next observation is the green dot. The goal is the blue ring centered at $(0,2.5)$, $b_{0}$ is the red ring centered at $(0,0)$, and the beacon is the green star at $(2.5,0)$. The reward function for the posterior state is the blue-red heatmap drawn in the background. The shaded area is where the observation noise is $\sigma_{O}>5$. The chosen action branch of AGMCTS is drawn as the blue arrows, getting darker in hue with each action update. We can see that the optimized actions point closer to the goal from the current belief’s mean.
This is a continuous extrapolation of the well-known Light-Dark POMDP. In this domain, $\mathcal{S}=\mathcal{O}=\mathbb{R}^{D}$. The agent starts at a random position on the sphere centered at the origin $S^{D-1}_{0.5}$. The goal position is located at $s_{g}=(\boldsymbol{0}_{D-1},2.5)$. A beacon is located to the side at $s_{b}=(2.5,\boldsymbol{0}_{D-1})$. $T=0.2$ is the goal tolerance, and the POMDP terminates if $\lVert{s-s_{g}}\rVert<T$. The horizon is $L=6$ and $\gamma=0.99$. The action space is the ball $B_{1.5}(\boldsymbol{0}_{D})$. The transition model is a simple Gaussian with added noise: $s^{\prime}\sim\mathcal{N}(s+a,(0.025)^{2}\mathbf{I})$. The observations are the agent’s relative position to the beacon, with a Gaussian added noise that increases with distance: $o\sim\mathcal{N}(s-s_{b},(\sigma_{O}(\lVert{s-s_{b}}\rVert))^{2}\mathbf{I})$, with noise function $\sigma_{O}(x)=\min\{15,0.01(x+x^{8})\}$. The reward function is $r(s,a,s^{\prime})=10\cdot g_{1}(d)-2\cdot g_{2}(d)-0.02d^{2}$, where $d\mathbin{=}\lVert{s^{\prime}-s_{g}}\rVert$. $g_{1}$ is a Gaussian centered at 00 with width $\frac{T}{2}$, $g_{2}$ is a Gaussian centered at $5T$ with width $T$. This reward function only depends on the posterior state. The agent needs to carefully target the goal as the reward decreases quickly from the goal, and advancing at random towards it will incur a penalty due to the moat $g_{2}$.
To simulate more difficult domains, we include a noisy rollout policy. The rollout heads directly to $s_{g}$, but the computed action has an added noise sampled from $\mathcal{N}(\mathbf{0},\sigma_{r}^{2}\mathbf{I})$.
4.2 Performance Evaluation
[…]
D $\sigma_{r}$ POMCPOW PFT-DPW AGMCTS
2 0.1 $\boldsymbol{6.62}\pm 0.07$ $5.71\pm 0.08$ $\boldsymbol{6.67}\pm 0.06$
2 0.2 $\boldsymbol{6.23}\pm 0.07$ $4.74,0.08$ $5.96\pm 0.08$
2 0.3 $\boldsymbol{5.61}\pm 0.08$ $4.01\pm 0.09$ $\boldsymbol{5.49}\pm 0.09$
3 0.1 $3.42\pm 0.08$ $3.63\pm 0.08$ $\boldsymbol{5.47}\pm 0.07$
3 0.2 $1.87\pm 0.08$ $1.79\pm 0.08$ $\boldsymbol{3.82}\pm 0.1$
3 0.3 $1.06\pm 0.07$ $0.29\pm 0.06$ $\boldsymbol{1.58}\pm 0.1$
4 0.1 $2.58\pm 0.07$ $2.36\pm 0.07$ $\boldsymbol{4.26},0.08$
4 0.2 $0.53\pm 0.05$ $0.2\pm 0.05$ $\boldsymbol{1.53}\pm 0.09$
4 0.3 $\boldsymbol{-0.36}\pm 0.03$ $-0.86\pm 0.02$ $-1.13\pm 0.04$
Table 1: Performance results for the different algorithms at D-Continuous Light Dark for $D=2,3,4$ dimensions. AGMCTS and PFT-DPW were run for 500 simulations, and POMCPOW for an equivalent time to PFT-DPW. The rollout policy’s standard deviation was $\sigma_{r}=0.1$ during hyperparameter tuning, and then the algorithms were also tested for $\sigma_{r}=0.2,0.3$. The reported results are mean returns over 1000 scenarios $\pm$ standard error. AGMCTS achieved the best performance overall in the different domains.
5 Conclusions
We introduce a novel methodology for considering dynamic action branches in MCTS in continuous (PO)MDP settings. The MIS tree allows action updates while keeping action-value estimates consistent. The new gradient derivation allows computing the action-value gradient in continuous POMDPs, and is supported by the MIS tree for its practical estimation. Our novel planning algorithm AGMCTS incorporates these two novelties together, and with additional algorithmic heuristics, is shown to achieve improved performance on a per-simulation budget. Our experiments suggest that gradient-based action updates may be more effective at higher dimensions, and where domain knowledge is less accurate.
Practical Considerations.
AGMCTS requires a transition likelihood model, both for MIS updates and for the gradient computation. It is application-dependent whether the computational cost of the gradient computation is worth the improvement in performance, compared to using more simulations. In our experiments, AGMCTS had significantly slower planning sessions compared to PFT-DPW. Each algorithm was run in a single thread, yet particle belief algorithms such as AGMCTS and PFT-DPW are highly parallelizable. Furthermore, AGMCTS could be combined with more sophisticated heuristics, like Voronoi-based action sampling and Bayesian optimization methods, which might boost performance. Finally, it is important to note that many algorithmic choices can be tailored depending on the application, like the MIS heuristic and the gradient optimizer, which may significantly affect run times.
Future Work.
The gradient computation and reuse that we’ve shown for POMDPs are based on the propagated belief, which requires full belief trajectories that might be computationally demanding in some domains. We plan to extend our methods efficiently for state trajectories as well. For the action selection we’ve used the standard UCT, however it is not obvious that it remains theoretically correct for the MIS estimator. It is likely that the exploration bonus varies at action updates, and further investigation is required. Finally, AGMCTS requires significant hyperparameter tuning, including DPW parameters and parameters related to action optimization. It is future research to find algorithms that are more general across different domains.
Acknowledgments
This research was partially supported by the Israel Science Foundation (ISF).
Appendix A Proofs and Derivations
A.1 Lemma 1
Lemma 1.
Let $\bar{b}^{-}=((s^{-,j},w^{j}))_{j=1}^{J}$ be generated from $\bar{b}=((s^{j},w^{j}))_{j=1}^{J}$ by SIS or SIR. Hence, the $j$th state particle was generated via $s^{-,j}\sim p_{T}(\cdot\mid s^{j},a)$, and the weight was copied, i.e. $w^{-,j}=w^{j}$. The ordered particle belief likelihood is then:
$\textstyle p_{T}(\bar{b}^{-}\mid\bar{b},a)=\prod_{i=1}^{J}p_{T}(s^{-,j}\mid s^{j},a).$ (20)
Proof.
In the SIS and SIR algorithms, each particle is sampled $s^{-,j}\sim p_{T}(\cdot\mid s^{j},a)$ independently for each $j=1,\ldots,J$. Therefore:
$p(\bar{b}^{-}\mid\bar{b},a)$ (21)
$=p(((s^{-,j},w^{-,j}))_{j=1}^{J}\mid((s^{l},w^{l}))_{l=1}^{J},a),$ (22)
due to the beliefs being ordered, and the independent sampling, we can factorize as:
$=\prod_{i=1}^{J}p(((s^{-,j},w^{-,j}))\mid((s^{l},w^{l}))_{l=1}^{J},a)$ (23)
$=\prod_{i=1}^{J}p(((s^{-,j},w^{-,j}))\mid(s^{j},w^{j}),a).$ (24)
Finally, since $w^{-,j}=w^{j}$, we can eliminate the weights to obtain
$p(\bar{b}^{-}\mid\bar{b},a)=\prod_{i=1}^{J}p_{T}(s^{-,j}\mid s^{j},a),$ (25)
and we denote $p_{T}(\bar{b}^{-}\mid\bar{b},a)\triangleq p(\bar{b}^{-}\mid\bar{b},a)$. ∎
A.2 Derivations of MIS Tree Update Equations
Here we show the derivation of the update equation, and also show their numerically stable forms using log likelihoods for probability values. We denote the regular and the weighted log-sum-exp function as $LSE$, differentiated by having one or two vector inputs, and they are defined as
$LSE(\boldsymbol{x}_{1,\ldots,n})\triangleq\log\left(\sum_{i=1}^{n}\exp(\boldsymbol{x}_{i})\right),$ (26)
$LSE(\boldsymbol{x}_{1,\ldots,n},\boldsymbol{a}_{1,\ldots,n})\triangleq\log\left(\sum_{i=1}^{n}\exp(\boldsymbol{x}_{i})\cdotp\boldsymbol{a}_{i}\right).$ (27)
To handle negative weights, (27) can be defined to compute the log of absolute value and sign separately of the log argument, such that the sign is multiplied back after exponentiation if required. We also denote a vector with brackets []delimited-[][$\cdot$][ ⋅ ].
We remind the visitation counters relationships:
$n(s)=\sum_{C(s)}n(s,a^{i}),\;n(s,a)=\sum_{C(s,a)}n(s^{\prime,i})_{+1},$ (28)
the state value estimator:
$\hat{V}(s)\triangleq\sum_{C(s)}\frac{n(s,a^{i})}{n(s)}\hat{Q}(s,a^{i}),$ (29)
the self normalized MIS estimator:
$\eta\triangleq\sum_{C(s,a)}\omega^{i}n(s^{\prime,i})_{+1},$ (30)
$\hat{V}_{f}(s,a)=\eta^{-1}\sum_{C(s,a)}\omega^{i}n(s^{\prime,i})_{+1}\hat{V}(s^{\prime,i}),$ (31)
and we also remind that we denoted $M\triangleq\lvert{C(s,a)}\rvert$.
We define the following vectors for each a given action node $(s,a)$ or $(\bar{b},a)$:
  1. 1.
    Child state visitation counters.
    MDP:
    $\boldsymbol{n}\triangleq(n(s^{\prime,i})_{+1})_{i=1}^{M},$ (32)
    PB-MDP:
    $\boldsymbol{n}\triangleq(n(\bar{b}^{\prime,i})_{+1})_{i=1}^{M}.$ (33)
  2. 2.
    Target transition likelihoods.
    MDP:
    $\boldsymbol{p}=(p_{T}(s^{\prime,i}\mid s,a))_{i=1}^{M},$ (34)
    PB-MDP:
    $\boldsymbol{p}=(p_{T}(\bar{b}^{-,i}\mid\bar{b},a))_{i=1}^{M}.$ (35)
  3. 3.
    Proposal transition likelihoods.
    MDP:
    $\boldsymbol{q}=(p_{T}(s^{\prime,i}\mid s,a_{prop}(s^{\prime,i})))_{i=1}^{M},$ (36)
    PB-MDP:
    $\boldsymbol{q}=(p_{T}(\bar{b}^{-,i}\mid\bar{b},a_{prop}(\bar{b}^{-,i})))_{i=1}^{M}.$ (37)
  4. 4.
    Immediate rewards.
    MDP:
    $\boldsymbol{r}=(r(s,a,s^{\prime,i}))_{i=1}^{M}$ (38)
    PB-MDP:
    $\boldsymbol{r}=(r(\bar{b},a,\bar{b}^{\prime,i}))_{i=1}^{M}.$ (39)
  5. 5.
    Future value estimates.
    MDP:
    $\boldsymbol{v}=(\hat{V}(s^{\prime,i}))_{i=1}^{M}$ (40)
    PB-MDP:
    $\boldsymbol{v}=(\hat{V}(\bar{b}^{\prime,i}))_{i=1}^{M}$ (41)
From here on the PB-MDP is the same up to different notations as the MDP case, and we will show only the latter. Using these notations the immediate reward and future value estimates can be compactly written:
$\eta=\boldsymbol{1}^{T}(\boldsymbol{n}\odot\boldsymbol{p}\oslash\boldsymbol{q}),$ (42)
$\hat{r}(s,a)=\eta^{-1}\boldsymbol{1}^{T}(\boldsymbol{n}\odot\boldsymbol{p}\odot\boldsymbol{r}\oslash\boldsymbol{q}),$ (43)
$\hat{V}_{f}(s,a)=\eta^{-1}\boldsymbol{1}^{T}(\boldsymbol{n}\odot\boldsymbol{p}\odot\boldsymbol{v}\oslash\boldsymbol{q}),$ (44)
[…]
We store the following values:
  1. 1.
    $l(\eta)$.
  2. 2.
    Log likelihoods $l(\boldsymbol{p})$ and $l(\boldsymbol{q})$.
  3. 3.
    $\boldsymbol{n}^{\prime}$ and previous visitation counters $\boldsymbol{n}$.
  4. 4.
    $\boldsymbol{v}^{\prime}$ and previous future value estimates $\boldsymbol{v}$.
The equations for $\hat{r}(s,a)$ are analogous to $\hat{V}_{f}(s,a)$, and we will not show their derivation explicitly.
Action Backpropagation.
Let $s^{\prime,j}$ be an updated node. Hence, we updated $\boldsymbol{n}_{j}$ to $\boldsymbol{n}^{\prime}_{j}$ and $\boldsymbol{v}_{j}$ to $\boldsymbol{v}^{\prime}_{j}$. We update $n(s,a)$ by (28) and perform
$\eta^{\prime}=\eta+\frac{\boldsymbol{p}_{j}}{\boldsymbol{q}_{j}}\left(\boldsymbol{n}_{j}^{\prime}-\boldsymbol{n}_{j}\right),$ (45)
$\hat{V}_{f}^{\prime}(s,a)=\frac{1}{\eta^{\prime}}\left(\eta\hat{V}_{f}(s,a)+\frac{\boldsymbol{p}_{j}}{\boldsymbol{q}_{j}}(\boldsymbol{n}_{j}^{\prime}\boldsymbol{v}_{j}^{\prime}-\boldsymbol{n}_{j}\boldsymbol{v}_{j})\right).$ (46)
Therefore, the log-form update equations are:
$l(\eta^{\prime})=LSE\left([l(\eta),l(\boldsymbol{p}_{j})-l(\boldsymbol{q}_{j})+l(\boldsymbol{n}_{j}^{\prime}-\boldsymbol{n}_{j})]\right)$ (47)
$\hat{V}_{f}^{\prime}(s,a)=\exp\Big{(}LSE\big{(}[l(\eta),l(\boldsymbol{p}_{j})-l(\boldsymbol{q}_{j})],$
$[\hat{V}_{f}(s,a),\boldsymbol{n}_{j}^{\prime}\boldsymbol{v}_{j}^{\prime}-\boldsymbol{n}_{j}\boldsymbol{v}_{j}]\big{)}-l(\eta^{\prime})\Big{)}.$ (48)
Proof.
We start with the update for $\eta^{\prime}$:
$\eta^{\prime}-\eta$ (49)
$=\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}^{\prime}\boldsymbol{p}_{i}}{\boldsymbol{q}_{i}}-\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}}{\boldsymbol{q}_{i}}$ (50)
$=\frac{\boldsymbol{p}_{j}}{\boldsymbol{q}_{j}}(\boldsymbol{n}_{j}^{\prime}-\boldsymbol{n}_{j}),$ (51)
due to all terms being equal except for $\boldsymbol{n}_{j}^{\prime}\neq\boldsymbol{n}_{j}$. Similarly, for $\hat{V}_{f}(s,a)$:
$\hat{V}_{f}^{\prime}(s,a)-\frac{\eta}{\eta^{\prime}}\hat{V}_{f}(s,a)$ (52)
$=\frac{1}{\eta^{\prime}}\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}^{\prime}\boldsymbol{p}_{i}\boldsymbol{v}_{i}^{\prime}}{\boldsymbol{q}_{i}}-\frac{\eta}{\eta^{\prime}}\frac{1}{\eta}\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}\boldsymbol{v}_{i}}{\boldsymbol{q}_{i}}$ (53)
$=\frac{1}{\eta^{\prime}}\left(\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}^{\prime}\boldsymbol{p}_{i}\boldsymbol{v}_{i}^{\prime}}{\boldsymbol{q}_{i}}-\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}\boldsymbol{v}_{i}}{\boldsymbol{q}_{i}}\right)$ (54)
$=\frac{1}{\eta^{\prime}}\frac{\boldsymbol{p}_{j}}{\boldsymbol{q}_{j}}(\boldsymbol{n}_{j}^{\prime}\boldsymbol{v}_{j}^{\prime}-\boldsymbol{n}_{j}\boldsymbol{v}_{j}),$ (55)
due to all terms being equal except for $\boldsymbol{v}^{\prime}_{j}\neq\boldsymbol{v}_{j}$, $\boldsymbol{n}^{\prime}_{j}\neq\boldsymbol{n}_{j}$. ∎
State Expansion.
This case is the same as action backpropagation, where the updated index is a new summand at index $j=M+1$. The same equations apply with $n(s^{\prime,j})=0$.
Proof.
We append new terms to all vectors: $\boldsymbol{n}_{M+1}$, $\boldsymbol{p}_{M+1}$, $\boldsymbol{v}_{M+1}$, $\boldsymbol{q}_{M+1}$. Update to $\eta$:
$\eta^{\prime}-\eta$ (56)
$=\sum_{i=1}^{M+1}\frac{\boldsymbol{n}_{i}^{\prime}\boldsymbol{p}^{\prime}_{i}}{\boldsymbol{q}^{\prime}_{i}}-\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}}{\boldsymbol{q}_{i}}$ (57)
$=\frac{\boldsymbol{n}_{M+1}^{\prime}\boldsymbol{p}^{\prime}_{M+1}}{\boldsymbol{q}^{\prime}_{M+1}},$ (58)
due to all terms being equal except for index $M+1$. This corresponds to (51) with $\boldsymbol{n}_{M+1}=0$, after initializing $\boldsymbol{p}^{\prime}_{M+1}=\boldsymbol{p}_{M+1}$ and $\boldsymbol{q}^{\prime}_{M+1}=\boldsymbol{q}_{M+1}$ Similarly, for $\hat{V}_{f}(s,a)$:
$\hat{V}_{f}^{\prime}(s,a)-\frac{\eta}{\eta^{\prime}}\hat{V}_{f}(s,a)$ (59)
$=\frac{1}{\eta^{\prime}}\sum_{i=1}^{M+1}\frac{\boldsymbol{n}_{i}^{\prime}\boldsymbol{p}_{i}\boldsymbol{v}_{i}^{\prime}}{\boldsymbol{q}_{i}}-\frac{\eta}{\eta^{\prime}}\frac{1}{\eta}\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}\boldsymbol{v}_{i}}{\boldsymbol{q}_{i}}$ (60)
$=\frac{1}{\eta^{\prime}}\left(\sum_{i=1}^{M+1}\frac{\boldsymbol{n}_{i}^{\prime}\boldsymbol{p}_{i}\boldsymbol{v}_{i}^{\prime}}{\boldsymbol{q}_{i}}-\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}\boldsymbol{v}_{i}}{\boldsymbol{q}_{i}}\right)$ (61)
$=\frac{1}{\eta^{\prime}}\frac{\boldsymbol{p}^{\prime}_{M+1}}{\boldsymbol{q}^{\prime}_{M+1}}\boldsymbol{n}_{M+1}^{\prime}\boldsymbol{v}_{M+1}^{\prime},$ (62)
which again corresponds to (55) with $\boldsymbol{n}_{M+1}=0$. ∎
Action Update.
Let $a^{\prime}$ be the new action. Here, we have to calculate $\omega^{i}$ for all $s^{\prime,i}\in C(s,a)$ in $O(M)$ time. We update
$(\omega^{i})^{\prime}=p_{T}(s^{\prime,i}\mid s,a^{\prime})(p_{T}(s^{\prime,i}\mid s,a(s^{\prime,i})))^{-1},$ (63)
and recalculate (30) and (31) with the new ($\omega^{i})^{\prime}$, which are also $O(M)$ time operations. For the immediate reward, new rewards must be computed for the new triplets $(s,a^{\prime},s^{\prime,i})$, and the update is analogous.
Therefore, the log-form equations are:
$l(\eta^{\prime})=LSE(l(\boldsymbol{n})\oplus l(\boldsymbol{p}^{\prime})\ominus l(\boldsymbol{q}))$ (64)
$\hat{V}_{f}^{\prime}(s,a)=\exp\left(LSE(l(\boldsymbol{p}^{\prime})\ominus l(\boldsymbol{q}),\boldsymbol{n}\odot\boldsymbol{v})-l(\eta^{\prime})\right)$ (65)
Proof.
This update corresponds to computing a new target likelihood vector $\boldsymbol{p}$. Therefore, the update equations follow from the definitions:
$\eta^{\prime}=\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}^{\prime}}{\boldsymbol{q}_{i}},$ (66)
$\hat{V}_{f}^{\prime}(s,a^{\prime})=\frac{1}{\eta^{\prime}}\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}^{\prime}\boldsymbol{v}_{i}}{\boldsymbol{q}_{i}}.$ (67)
Terminal State Backpropagation.
Let $s$ be a terminal node. Its value estimate is based only on rollouts. Hence, for a new rollout value $v^{\prime}$, we perform a running average update
$\hat{V}^{\prime}(s)=\hat{V}(s)+\frac{n^{\prime}(s)-n(s)}{n^{\prime}(s)_{+1}}(v^{\prime}-\hat{V}(s)),$ (68)
where $n^{\prime}(s)$ is $n(s)$ plus the number of rollouts (usually 1).
Proof.
This is a simple running average update for the mean estimator:
$\hat{V}(s)=\frac{1}{n(s)_{+1}}\sum_{i=1}^{n(s)_{+1}}v^{i},$ (69)
for all $v^{i}$ rollout values from the terminal node $s$. ∎
Non-Terminal State Backpropagation.
Let $(s,a^{j})$ be an updated node. Hence, we updated $n(s,a^{j})$ to $n^{\prime}(s,a^{j})$ and $\hat{Q}(s,a^{j})$ to $\hat{Q}^{\prime}(s,a^{j})$. We update $n^{\prime}(s)$ by (28) and perform
$\hat{V}^{\prime}(s)=(n^{\prime}(s))^{-1}\Big{(}n(s)\hat{V}(s)+n^{\prime}(s,a^{j})\hat{Q}^{\prime}(s,a^{j})$
${}-n(s,a^{j})\hat{Q}(s,a^{j})\Big{)}.$ (71)
Proof.
Following the definition of the state value estimator (29), the updated estimator is:
$\hat{V}^{\prime}(s)=\sum_{C(s)}\frac{n^{\prime}(s,a^{i})}{n^{\prime}(s)}\hat{Q}^{\prime}(s,a^{i}).$ (72)
Therefore:
$\hat{V}^{\prime}(s)-\frac{n(s)}{n^{\prime}(s)}\hat{V}(s)$ (73)
$=\sum_{C(s)}\frac{n^{\prime}(s,a^{i})}{n^{\prime}(s)}\hat{Q}^{\prime}(s,a^{i})-\frac{n(s)}{n^{\prime}(s)}\sum_{C(s)}\frac{n(s,a^{i})}{n(s)}\hat{Q}(s,a^{i})$ (74)
$=\frac{1}{n^{\prime}(s)}\sum_{C(s)}(n^{\prime}(s,a^{i})\hat{Q}^{\prime}(s,a^{i})-n(s,a^{i})\hat{Q}(s,a^{i})).$ (75)
All terms are equal except for the updated node $j$. Hence, we obtain
$\hat{V}^{\prime}(s)-\frac{n(s)}{n^{\prime}(s)}\hat{V}(s)$
$=\frac{1}{n^{\prime}(s)}\left(n^{\prime}(s,a^{j})\hat{Q}^{\prime}(s,a^{j})-n(s,a^{j})\hat{Q}(s,a^{j})\right).$ (76)
A.3 Theorem 1
We first highlight that the required ”well-behaved” conditions for the theorem are for changing the order of derivatives and integrals by the Leibniz integral rule. A sufficient condition is that the integrand and all its partial derivatives are continuous in all variables, which is satisfied when both $p_{T},r$ satisfy these conditions. Stronger versions of this theorem exist, such as only requiring that the partial derivatives exist almost everywhere. As this is not the focus of our discussion, we leave that for future work. Additionally, as stated in the main paper, it must be satisfied that $p_{T}(s^{\prime}\mid s,a)=0$ implies $p_{T}(s^{\prime}\mid s,a^{\prime})=0$.
Theorem 1.
For well-behaved $p_{T}$, $r$, the MDP action gradient of the action-value function is given by the expectation
$\nabla_{a^{\prime}}Q_{t}^{\pi}(s,a^{\prime})=\mathbb{E}_{s^{\prime}\mid s,a}\bigg{[}\frac{p_{T}(s^{\prime}\mid s,a^{\prime})}{p_{T}(s^{\prime}\mid s,a)}$
$(\nabla_{a^{\prime}}\log p_{T}(s^{\prime}\mid s,a^{\prime})$
$\cdot(r(s,a^{\prime},s^{\prime})+\gamma V_{t+1}^{\pi}(s^{\prime}))+\nabla_{a^{\prime}}r(s,a^{\prime},s^{\prime}))\bigg{]}.$ (77)
Additionally, for well-behaved $p_{T}$, $p_{O}$, $r$, the POMDP action gradient of the action-value function can be computed by replacing posterior belief likelihood terms with propagated belief likelihood terms, i.e. $b_{t+1}^{-}$ instead of $b_{t+1}$ in all $p_{T}$ terms.
Proof.
We start with the MDP case.
$\nabla_{a^{\prime}}Q_{t}^{\pi}(s,a^{\prime})$ (78)
$=\nabla_{a^{\prime}}\mathbb{E}_{s{{}^{\prime}}\mid s,a}\bigg{[}\frac{p(s{{}^{\prime}}\mid s,a{{}^{\prime}})}{p(s{{}^{\prime}}\mid s,a)}(r(s,a{{}^{\prime}},s{{}^{\prime}})+\gamma V_{t+1}^{\pi}(s{{}^{\prime}}))\bigg{]},$ (79)
Applying the Leibniz integral rule:
$=\mathbb{E}_{s{{}^{\prime}}\mid s,a}\bigg{[}\nabla_{a^{\prime}}\frac{p(s{{}^{\prime}}\mid s,a{{}^{\prime}})}{p(s{{}^{\prime}}\mid s,a)}\cdot(r(s,a{{}^{\prime}},s{{}^{\prime}})+\gamma V_{t+1}^{\pi}(s{{}^{\prime}}))$
$+\frac{p(s{{}^{\prime}}\mid s,a{{}^{\prime}})}{p(s{{}^{\prime}}\mid s,a)}\cdot\nabla_{a^{\prime}}(r(s,a{{}^{\prime}},s{{}^{\prime}})+\gamma V_{t+1}^{\pi}(s{{}^{\prime}}))\bigg{]}.$ (80)
For the first term, we use the log-gradient substitution:
$\nabla_{a^{\prime}}p(s{{}^{\prime}}\mid s,a{{}^{\prime}})=p(s{{}^{\prime}}\mid s,a{{}^{\prime}})\cdot\nabla_{a^{\prime}}\log p(s{{}^{\prime}}\mid s,a{{}^{\prime}}),$ (81)
and for the second term, since $V_{t+1}^{\pi}(s^{\prime})$ is not a function of $a^{\prime}$, we obtain:
$\nabla_{a^{\prime}}(r(s,a{{}^{\prime}},s{{}^{\prime}})+\gamma V_{t+1}^{\pi}(s{{}^{\prime}}))=\nabla_{a^{\prime}}r(s,a{{}^{\prime}},s{{}^{\prime}}).$ (82)
By gathering terms we obtain the result (77).
We now show the POMDP case. We assume that the same conditions are satisfied for $p_{T},p_{O},r$. We first show the propagated belief trick.
$Q_{t}^{\pi}(b,a)=\mathbb{E}_{b^{\prime}\mid b,a}\left[r(b,a,b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime})\right]$ (83)
$=\int_{b^{\prime}}p(b^{\prime}\mid b,a)(r(b,a,b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime}))\mathop{}\!\mathrm{d}b^{\prime},$ (84)
and we continue by marginalizing over the propagated belief $b^{-}$ and the observation $o$:
$=\int_{b^{-}}\int_{o}\int_{b^{\prime}}p(b^{-},o,b^{\prime}\mid b,a)$
$(r(b,a,b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime}))\mathop{}\!\mathrm{d}b^{\prime}\mathop{}\!\mathrm{d}o\mathop{}\!\mathrm{d}b^{-}.$ (85)
By its definition, the propagated belief $b^{-}$ is the generated belief by conditioning on the history without the last observation. Therefore, using the chain rule and removing variables that are conditionally independent, we obtain:
[…]
$p(b^{-}\mid b,a)(r(b,a,b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime}))\mathop{}\!\mathrm{d}b^{\prime}\mathop{}\!\mathrm{d}o\mathop{}\!\mathrm{d}b^{-}$ (86)
$=\int_{b^{-}}\int_{o}\int_{b^{\prime}}p(b^{\prime}\mid o,b^{-})p(o\mid b^{-})p(b^{-}\mid b,a)$
$(r(b,a,b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime}))\mathop{}\!\mathrm{d}b^{\prime}\mathop{}\!\mathrm{d}o\mathop{}\!\mathrm{d}b^{-}$ (87)
$=\mathbb{E}_{b^{-}\mid b,a}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}\left[r(b,a,b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime})\right].$ (88)
We introduce an importance ratio over the propagated belief to calculate the derivative for proposal action $a$ and target action $a^{\prime}$:
$Q_{t}^{\pi}(b,a^{\prime})=\mathbb{E}_{b^{-}\mid b,a}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}\bigg{[}\frac{p_{T}(b^{-}\mid b,a^{\prime})}{p_{T}(b^{-}\mid b,a)}$
$(r(b,a^{\prime},b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime}))\bigg{]},$ (89)
where we denoted $p_{T}(b^{-}\mid b,a^{\prime})\triangleq p(b^{-}\mid b,a^{\prime})$ as the transition model for the belief MDP. Similarly to the derivation of the MDP case, we can apply the Leibniz integral rule to obtain the result:
$\nabla_{a^{\prime}}Q_{t}^{\pi}(b,a^{\prime})=\mathbb{E}_{b^{-}\mid b,a}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}\bigg{[}\frac{p_{T}(b^{-}\mid b,a^{\prime})}{p_{T}(b^{-}\mid b,a)}$
$(\nabla_{a^{\prime}}\log p_{T}(b^{-}\mid b,a^{\prime})$
$\cdot(r(b,a^{\prime},b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime}))+\nabla_{a^{\prime}}r(b,a^{\prime},b^{\prime}))\bigg{]}.$ (90)
A.4 Theorem 2
We highlight that as an extension of Theorem 1, it requires the same conditions. Additionally, for the simplification of the immediate reward expectation, we make use of Fubini’s theorem for changing integration orders. Sufficient conditions are that there exists some $R\in\mathbb{R}$ such that $\sup_{s,a,s^{\prime}}\lvert{r(s,a,s^{\prime})}\rvert<R$, and that $r,p_{T},p_{O}$ are all continuous.
Theorem 2.
For a state-reward POMDP, the action gradient of the action-value function is given by:
$\nabla_{a^{\prime}}Q_{t}^{\pi}(b,a)=\mathbb{E}_{s\mid b}\mathbb{E}_{s^{\prime}\mid s,a^{\prime}}[\nabla_{a^{\prime}}\log p_{T}(s^{\prime}\mid s,a^{\prime})\cdot r(s,a^{\prime},s^{\prime})$
$+\nabla_{a^{\prime}}r(s,a^{\prime},s^{\prime})]+\gamma\mathbb{E}_{b^{-}\mid b,a}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}\bigg{[}\frac{p_{T}(b^{-}\mid b,a^{\prime})}{p_{T}(b^{-}\mid b,a)}$
$\nabla_{a^{\prime}}\log p_{T}(b^{-}\mid b,a^{\prime})\cdot V_{t+1}^{\pi}(b^{\prime})\bigg{]}.$ (91)
Proof.
The central part of this proof is the decomposition of the belief action-value function to an expectation over state trajectories of immediate rewards, and an expectation over belief trajectories of future values. Afterwards, we apply the Leibniz integral rule similarly to Theorem 1.
We start with the belief action-value function:
$Q_{t}^{\pi}(b,a)=\mathbb{E}_{b^{\prime}\mid b,a}\left[r(b,a,b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime})\right]$ (92)
$\mathbb{E}_{b^{\prime}\mid b,a}\left[r(b,a,b^{\prime})\right]+\gamma\mathbb{E}_{b^{\prime}\mid b,a}\left[V_{t+1}^{\pi}(b^{\prime})\right].$ (93)
We expand the first expectation, using the chain rule and then Fubini’s theorem to cancel the expectation over the observation and posterior belief:
$\mathbb{E}_{b^{\prime}\mid b,a}\left[r(b,a,b^{\prime})\right]$ (94)
$=\mathbb{E}_{b^{-}\mid b,a}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}\left[\mathbb{E}_{s,s^{\prime}\mid b,a,b^{\prime}}[r(s,a,s^{\prime})]\right]$ (95)
$=\int_{b^{-}}\int_{o}\int_{b^{\prime}}\int_{s}\int_{s^{\prime}}p(s,s^{\prime},o,b^{-},b^{\prime}\mid b,a)$
$r(s,a,s^{\prime})\mathop{}\!\mathrm{d}s^{\prime}\mathop{}\!\mathrm{d}s\mathop{}\!\mathrm{d}b^{\prime}\mathop{}\!\mathrm{d}o\mathop{}\!\mathrm{d}b^{-}$ (96)
$=\int_{b^{-}}\int_{o}\int_{b^{\prime}}\int_{s}\int_{s^{\prime}}p(o,b^{\prime},b^{-}\mid s,s^{\prime},b,a)p(s,s^{\prime}\mid b,a)$
$r(s,a,s^{\prime})\mathop{}\!\mathrm{d}s^{\prime}\mathop{}\!\mathrm{d}s\mathop{}\!\mathrm{d}b^{\prime}\mathop{}\!\mathrm{d}o\mathop{}\!\mathrm{d}b^{-}$ (97)
$=\int_{s}\int_{s^{\prime}}p(s,s^{\prime}\mid b,a)r(s,a,s^{\prime})$
(∫$\cancel{\left(\int_{b^{-}}\int_{o}\int_{b^{\prime}}p(o,b^{\prime},b^{-}\mid s,s^{\prime},b,a)\mathop{}\!\mathrm{d}b^{\prime}\mathop{}\!\mathrm{d}o\mathop{}\!\mathrm{d}b^{-}\right)}\mathop{}\!\mathrm{d}s^{\prime}\mathop{}\!\mathrm{d}s$ (98)
[…]
$=\mathbb{E}_{s\mid b}\mathbb{E}_{s^{\prime}\mid s,a}[r(s,a,s^{\prime})].$ (100)
Now, we compute $\nabla_{a^{\prime}}$ for the terms obtained. The gradient of the immediate reward expectation, when the proposal action is $a$ and the target action is $a^{\prime}$, is given by (skipping some steps similar to Theorem 1):
$\nabla_{a^{\prime}}\mathbb{E}_{s\mid b}\mathbb{E}_{s^{\prime}\mid s,a}\bigg{[}\frac{p_{T}(s^{\prime}\mid s,a^{\prime})}{p_{T}(s^{\prime}\mid s,a)}r(s,a^{\prime},s^{\prime})\bigg{]}$ (101)
$=\mathbb{E}_{s\mid b}\mathbb{E}_{s^{\prime}\mid s,a}\bigg{[}\frac{p_{T}(s^{\prime}\mid s,a^{\prime})}{p_{T}(s^{\prime}\mid s,a)}\big{(}\nabla_{a^{\prime}}\log p_{T}(s^{\prime}\mid s,a^{\prime})$
$\cdot r(s,a^{\prime},s^{\prime})+\nabla_{a^{\prime}}r(s,a^{\prime},s^{\prime})\big{)}\bigg{]},$ (102)
and the importance ratio cancels out for $a=a^{\prime}$.
Next, the gradient of the future value expectation is simply:
$\nabla_{a^{\prime}}\mathbb{E}_{b^{\prime}\mid b,a}\left[V_{t+1}^{\pi}(b^{\prime})\right]$ (103)
$=\nabla_{a^{\prime}}\mathbb{E}_{b^{-}\mid b,a}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}\bigg{[}\frac{p_{T}(b^{-}\mid b,a^{\prime})}{p_{T}(b^{-}\mid b,a)}V_{t+1}^{\pi}(b^{\prime})\bigg{]}$ (104)
$=\mathbb{E}_{b^{-}\mid b,a}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}\bigg{[}\frac{p_{T}(b^{-}\mid b,a^{\prime})}{p_{T}(b^{-}\mid b,a)}$
$\nabla_{a^{\prime}}\log p_{T}(b^{-}\mid b,a^{\prime})\cdot V_{t+1}^{\pi}(b^{\prime})\bigg{]}.$ (105)
Combining the two results, we obtain the final expression (91). ∎
Appendix B Experiments
B.1 AGMCTS Implementation Details
Optimizer
For the Adam optimizer, we used the hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$. The learning rate $\eta_{Adam}$ was determined by cross entropy optimization (CE), described later, in conjunction with the optimization over the other algorithm hyperparameters.
In our implementation we used an exponential decay on the step size after the Adam step size. The calculation for the update of the accumulated action is done as:
δAdam=
A
DAM
(a,gaq)
[…]
A
DAM
[…]
$\lambda=\max\{$0.999$^{T},$0.1$\}a^{acc}=a+\lambda\delta_{Adam},$ (107)
where $T$ is the number of gradient optimizations performed at this particular action node.
B.2 D-Continuous Light Dark
In this domain, $\mathcal{X}=\mathcal{Z}=\mathbb{R}^{D}$. The agent starts at a random position on the sphere of radius $r_{0}$ centered at the origin: $S^{D-1}_{r_{0}}$. The goal position is located at $x_{g}=(\boldsymbol{0}_{D-1},r_{g})$. A beacon is located to the side at $x_{b}=(r_{b},\boldsymbol{0}_{D-1})$. The constants we took are $r_{a}=1.5$, $r_{g}=r_{b}=2.5$, $r_{0}=0.5$. $T=0.2$ is the goal tolerance, and the POMDP terminates if $\lVert{x-x_{g}}\rVert<T$, or after $L=6$ time steps. We set $\gamma=0.99$.
The action space is $B_{r_{a}}(\boldsymbol{0}_{D})$. The transition model is a simple gaussian with added noise: $x^{\prime}\sim\mathcal{N}(x+a,\sigma_{T}^{2}\mathbf{I})$ with $\sigma_{T}=0.025$.
The observations are the agent’s relative position to the beacon, with a gaussian added noise that indreases with distance: $z\sim\mathcal{N}(x-x_{b},(\sigma_{Z}(\lVert{x-x_{b}}\rVert))^{2}\mathbf{I})$, with noise function $\sigma_{Z}(x)=\min\{\sigma_{Z}^{\max},k_{\sigma_{Z}}(x+x^{\alpha_{\sigma_{Z}}})\}$, for $\sigma_{Z}^{\max}=15$, $k_{\sigma_{Z}}=0.01$, $\alpha_{\sigma_{Z}}=8$.
The reward function is $r(x,a,x^{\prime})=R_{goal}\exp(-\frac{1}{2}{(\frac{d}{0.5T})}^{2})-R_{moat}\exp(-\frac{1}{2}{(\frac{d-5T}{T})}^{2})-R_{dist}d^{2}$, where $d\mathbin{=}\lVert{x-x_{g}}\rVert$. We took constants $R_{goal}\mathbin{=}10$, $R_{moat}\mathbin{=}2$, $R_{dist}\mathbin{=}0.02$. This reward function only depends on the posterior state. The agent needs to carefully target the goal as the reward decreases quickly from the center, and advancing at random towards the goal will incur a penalty due to the moat.
To simulate more difficult domains, we include a noisy rollout policy. The rollout heads directly to $x_{g}$, but the computed action has an added noise of $\mathcal{N}(\mathbf{0},\sigma_{r}^{2}\mathbf{I})$. In our experiments we used $\sigma_{r}=0.1$ for parameter tuning, and additionally evaluated for $\sigma_{r}=0.2,0.3$.
The implementation of D-Continuous Light Dark was written using POMDPs.jl [?]. The gradient function $\nabla_{a}r$ was equal 00 in our setting, and $\nabla_{a}\log p_{T}$ was calculated using Enzyme.jl [?]. Enzyme.jl provided extremely efficient gradients, achieving the execution times almost as close to the original functions. We did not care to optimize for making primal and dual calls, which might have allowed to save some computation time. However, it seems that it would have had a relatively minor effect as the heaviest computation was calculating the belief transition importance ratio, which due to it being a ratio of two products, cannot be MC approximated without introducing bias. Therefore, we iterated over all belief particles for its computation. Yet, in more complex problem where the dynamics are more complex, it might prove beneficial to compute gradients in the same pass as the transition likelihood wherever it can be done.
B.3 Evaluations
Parameter Notation Explanation
$c$ UCT exploration bonus
$k_{a}$ Action progressive widening (factor)
$\alpha_{a}$ Action progressive widening (power)
$k_{o}$ Observation progressive widening (factor)
$\alpha_{o}$ Observation progressive widening (power)
$\eta_{Adam}$ Adam step size (i.e. learning rate)
$T_{d_{a}}$ Action update distance threshold
Table 2: Hyperparameters optimized by CE in the evaluations of each algorithm for the D-Continuous Light Dark domain, for $D$ dimensions.
Parameter D POMCPOW PFT-DPW AGMCTS
$c$ 2 0.983 1.689 4.026
$k_{a}$ 2 0.350 7.332 8.346
$\alpha_{a}$ 2 0.834 0.473 0.515
$k_{o}$ 2 0.215 10.49 12.03
[…]
$c$ 3 1.024 2.429 5.212
$k_{a}$ 3 0.485 7.309 8.075
$\alpha_{a}$ 3 0.582 0.326 0.471
$k_{o}$ 3 0.744 11.27 15.20
$\alpha_{o}$ 3 0.226 0.195 0.317
[…]
$c$ 4 1.259 1.111 2.625
$k_{a}$ 4 0.360 9.309 8.043
$\alpha_{a}$ 4 0.559 0.343 0.495
$k_{o}$ 4 1.023 10.48 17.21
$\alpha_{o}$ 4 0.278 0.109 0.460
[…]
Table 3: Hyperparameters used in the evaluations of each algorithm for the D-Continuous Light Dark domain, for $D$ dimensions, rounded to $3$ significant digits after the decimal point.
We evaluated the algorithms on the D-Continuous Light Dark domain for $D=2,3,4$ dimensions. PFT-DPW and AGMCTS were given a budget of 500 simulations. The particle count during planning was set to be dependent on the dimension:
$J_{D=2}=256,\quad J_{D=3}=512,$
$J_{D=4}=1024.$ (108)
The particle filter used for inference between planning steps had a particle count $J^{\text{PF}}$ given by:
$J^{\text{PF}}_{D=2}=2048,\quad J^{\text{PF}}_{D=3}=4096,$
$J^{\text{PF}}_{D=4}=8192.$ (109)
POMCPOW was given a simulation budget according to the formula
$n_{D}^{sims,POMCPOW}=500\cdot$0.08$\cdot J_{D},$ (110)
which was empirically found to produce similar runtimes to PFT-DPW in the chosen domain.
In order to save computation time and not compute a partially observable rollout, the rollout computation for a new particle belief was done by drawing $s_{0}$ and 10 additional particles from the belief. The rollout actions were computed based on $s_{0}$, but the return was the mean of returns from the 10 particles. This mechanism allows to approximate a partially observable rollout without receiving any future observations. For POMCPOW, since every new node is expanded initially only with a single particle, the rollout was computed with that particle. This is a disadvantage for POMCPOW as it made the rollout returns less meaningful, however it seemd to have achieved better results than PFT-DPW despite that.
The algorithm hyperparameters were first determined by cross entropy optimization (CE) [?], largely following [?], using Gaussian distributions over the continuous DPW hyperparameters:
  1. 1.
    UCT exploration bonus $c$.
  2. 2.
    Action progressive widening parameters $k_{a}$ and $\alpha_{a}$.
  3. 3.
    Observation progressive widening parameters $k_{o}$ and $\alpha_{o}$.
For AGMCTS, we also optimized over the Adam learning rate $\eta_{Adam}$, and the action update distance threshold $T_{d_{a}}$. We searched for hyperparameters for each dimension while using rollout policy noise $\sigma_{r}=0.1$, and the results for $\sigma_{r}=0.2,0.3$ were obtained using the same hyperparameters of the same dimension. We used 150 parameter samples, 40 simulations for each parameter sample to determine its mean return, used 30 elite samples to fit the new distribution, with smoothing the new distribution’s parameters:
$\mu_{\text{new}}=\alpha_{\mu}\mu_{\text{new}}+(1-\alpha_{\mu})\mu_{\text{old}},$ (111)
$\Sigma_{\text{new}}=\alpha_{\Sigma}\Sigma_{\text{new}}+(1-\alpha_{\Sigma})\Sigma_{\text{old}},$ (112)
where we chose $\alpha_{\mu}=0.8,\alpha_{\Sigma}=0.5$.
We performed 50 CE iterations for each algorithm in each dimension. The starting mean for $D=2$ was the same for all algorithms, and for $D=3,4$ we used the mean of the best hyperparameters found for $D-1$. The hyperparameters chosen were those from the CE iteration achieving the maximal mean of mean returns over the 30 elite samples. The parameters used for each algorithm are summarized in Table 3. After choosing the parameters, we reported the mean return over 1000 scenarios, with the standard error of the mean. Table 4 summarizes the results with the individual mean runtimes for each algorithm in each scenario.
D $\sigma_{r}$ POMCPOW time (s) PFT-DPW time (s) AGMCTS time (s)
2 0.1 $\boldsymbol{6.62}\pm 0.07$ 0.114 $5.71\pm 0.08$ 0.126 $\boldsymbol{6.67}\pm 0.06$ 1.665
2 0.2 $\boldsymbol{6.23}\pm 0.07$ 0.122 $4.74,0.08$ 0.129 $5.96\pm 0.08$ 1.730
2 0.3 $\boldsymbol{5.61}\pm 0.08$ 0.129 $4.01\pm 0.09$ 0.129 $\boldsymbol{5.49}\pm 0.09$ 1.838
3 0.1 $3.42\pm 0.08$ 0.216 $3.63\pm 0.08$ 0.299 $\boldsymbol{5.47}\pm 0.07$ 2.442
3 0.2 $1.87\pm 0.08$ 0.240 $1.79\pm 0.08$ 0.296 $\boldsymbol{3.82}\pm 0.1$ 2.416
3 0.3 $1.06\pm 0.07$ 0.253 $0.29\pm 0.06$ 0.309 $\boldsymbol{1.58}\pm 0.1$ 2.389
4 0.1 $2.58\pm 0.07$ 0.548 $2.36\pm 0.07$ 0.683 $\boldsymbol{4.26},0.08$ 10.576
4 0.2 $0.53\pm 0.05$ 0.640 $0.2\pm 0.05$ 0.707 $\boldsymbol{1.53}\pm 0.09$ 8.303
4 0.3 $\boldsymbol{-0.36}\pm 0.03$ 0.660 $-0.86\pm 0.02$ 0.736 $-1.13\pm 0.04$ 5.778
Table 4: Performance results for the different algorithms at D-Continuous Light Dark for $D=2,3,4$ dimensions. AGMCTS and PFT-DPW were run for 500 simulations with particle counts given by (108), and POMCPOW for an approximately equivalent time to PFT-DPW according to (110). The rollout policy’s standard deviation was $\sigma_{r}=0.1$ during hyperparameter tuning, and then the algorithms were also tested for $\sigma_{r}=0.2,0.3$. The reported results are mean returns over 1000 scenarios $\pm$ standard error. AGMCTS achieved the best performance in most scenarios.

### v3

@@
-Value Gradients with Action Adaptive Search Trees in Continuous (PO)MDPs
+Action-Gradient Monte Carlo Tree Search for Non-Parametric Continuous (PO)MDPs
 Abstract
-Solving Partially Observable Markov Decision Processes (POMDPs) in continuous state, action and observation spaces is key for autonomous planning in many real-world mobility and robotics applications.
-Current approaches are mostly sample based, and cannot hope to reach near-optimal solutions in reasonable time.
-We propose two complementary theoretical contributions.
-First, we formulate a novel Multiple Importance Sampling (MIS) tree for value estimation, that allows to share value information between sibling action branches.
-The novel MIS tree supports action updates during search time, such as gradient-based updates.
-Second, we propose a novel methodology to compute value gradients with online sampling based on transition likelihoods.
-It is applicable to MDPs, and we extend it to POMDPs via particle beliefs with the application of the propagated belief trick.
-The gradient estimator is computed in practice using the MIS tree with efficient Monte Carlo sampling.
-These two parts are combined into a new planning algorithm Action Gradient Monte Carlo Tree Search (AGMCTS).
-We demonstrate in a simulated environment its applicability, advantages over continuous online POMDP solvers that rely solely on sampling, and we discuss further implications.
+Autonomous systems that operate in continuous state, action, and observation spaces require planning and reasoning under uncertainty.
+Existing online planning methods for such POMDPs are almost exclusively sample-based, yet they forego the power of high-dimensional gradient optimization as combining it into Monte Carlo Tree Search (MCTS) has proved difficult, especially in non-parametric settings.
+We close this gap with three contributions.
+First, we derive a novel action-gradient theorem for both MDPs and POMDPs in terms of transition likelihoods, making gradient information accessible during tree search.
+Second, we introduce the Multiple Importance Sampling (MIS) tree, that re-uses samples for changing action branches, yielding consistent value estimates that enable in-search gradient steps.
+Third, we derive exact transition probability computation via the Area Formula for smooth generative models common in physical domains, a result of independent interest.
+These elements combine into Action-Gradient Monte Carlo Tree Search (AGMCTS), the first planner to blend non-parametric particle search with online gradient refinement in POMDPs.
+Across several challenging continuous MDP and POMDP benchmarks, AGMCTS outperforms widely-used sample-only solvers in solution quality.
 1 Introduction
-Figure 1: Action adaptive MIS tree with two levels of actions and children state nodes.
-Usual states for MDPs are replaced by belief states for POMDPs.
-The actions in the action nodes are updated as the search progresses, indicated with a darker shade, and new states are sampled from the newest actions, denoted in the same color.
-The MCTS selection path, highlighted in blue, chooses at action nodes the latest action, and then proceeds at random to a child state.
-The Markov Decision Process (MDP) and the Partially Observable Markov Decision Process (POMDP) are flexible mathematical frameworks for modeling decision-making under uncertainty.
-Computing policies for (PO)MDPs that are in continuous in states, actions and observations, has mostly focused on global search algorithms based on POMCP [?], like POMCPOW and PFT-DPW [?], that explore the action space only by sampling.
-Several algorithms have been suggested for mixed continuous settings, in which not all the state, action and observation spaces are continuous.
-For continuous state and action spaces, GPS-ABT [?] and LCEOPT [?] use generalized pattern search and cross-entropy optimization respectively, both of which are gradient-free optimization methods.
-For continuous state and observation spaces, AdaOPS [?] suggested belief packing to combine values for similar beliefs, and LABECOP [?] introduced lazy belief extraction to increase belief diversity.
-The first example of an algorithm with performance guarantees for continuous POMDPs was VOWSS [?], that combined Voronoi Optimistic Optimization [?] for action sampling with Weighted Sparse Sampling [?].
-The authors also suggested a practical version of the algorithm, VOMCPOW, that combines MCTS with Double Progressive Widening (DPW).
-[?] presented hierarchical Voronoi-based action sampling, showing improved performance at action dimensionality of 6 to 12.
-Our research is motivated by the idea that in many continuous settings, gradient-based optimization is often much faster to converge than gradient-free optimization.
+Planning under uncertainty is a fundamental problem in artificial intelligence, and doing so efficiently holds key for integrating autonomous agents in our daily lives.
+Specifically in robotics and physical systems, the planning problem is often continuous in states, actions and observations.
+The Markov Decision Process (MDP) and the Partially Observable Markov Decision Process (POMDP) are flexible mathematical frameworks for modeling long-term decision-making under uncertainty, in which policies are computed to maximize long-term goals.
+Efficiently solving these problems however is not trivial, with specific research on the continuous case being fairly recent.
+Online planners mostly focused on exploring the search space by gradually increasing a sampling budget in a technique known as Double Progressive Widening (DPW) [1], with notable POMDP planners being POMCPOW and PFT-DPW [2].
+This blind approach is often inefficient, as it requires evaluating many actions to find a near-optimal one.
+To address this, recent work in continuous domains has explored black-box optimization methods that guide sampling.
+Techniques such as Voronoi optimistic optimization [3, 4, 5], kernel regression [6] and Bayesian optimization [7, 8] have all been proposed to bias action selection towards promising regions.
+However, they do not leverage the internal structure of the transition or observation model, missing the opportunity for model-aware introspection to improve planning efficiency.
+Our research is motivated by the idea that in many continuous settings, gradient-informed optimization is often much faster to converge than gradient-free methods.
 It might be prohibitively expensive to sample enough actions to statistically find a near-optimal one.
-In our approach we consider non-parametric beliefs, making it applicable to a wide range of problems, specifically not restricted by Gaussian or other parametric assumptions.
-Our methodology consists of two parts.
-First, we consider how to update actions during search time, without losing information of previous action-value estimates.
-Second, we derive value gradients w.r.t.
-actions, for the purpose of computing action gradient updates.
-We hypothesize that incorporating gradient optimization into continuous (PO)MDP planning will explore the action space more efficiently, especially at high dimensional settings.
-A similar idea has been recently explored in the algorithm VG-UCT by [?], where the local gradient-based optimization is combined with the global-search of sampled-based MCTS, but with several drawbacks compared to our approach.
-Their gradient computation approach is computationally expensive due to being based on finite differences and being recursive.
-They assume a deterministic transition model given a noise parameter, which is problematic for deriving gradients in POMDPs.
-Lastly, they do not address the distribution shift that occurs from action updates, which affect the action-value estimates.
-Our approach requires access to the state transition model’s likelihood.
-This is in contrast to previous online (PO)MDP solvers and model-free approaches that consider only use of a generative model.
-Yet, recent advances in physical simulators have included scalable differentiable physics simulators [??], and GPU-parallelization that have allowed generating accurate simulations with speedups in the scale of orders of magnitudes [??].
-We foresee that this will allow to compute the differentiable physical transition models more easily in the near future.
-1.1 Additional Related Work
-Information sharing in discrete MDPs with repeating states was considered in [?].
-In POMDPs, previous knowledge utilization via Multiple Importance Sampling (MIS) [?] was introduced in [?].
-We share information only between sibling action branches, which eliminates a costly search through a belief database.
-MIS in POMDP planning for sample reuse was also considered in [?], however in the context of calculating observation likelihoods.
-KR-UCT [?] use kernel regression to estimate the value of actions by sharing information between sibling action branches, and for the proposal of new actions.
-They consider MDPs only, and do not consider updating existing actions.
-In similar vein, Bayesian optimization for information sharing between different actions was explored in kCBTS [?] and BOMCP [?].
-These methods are purely heuristic, and the information sharing is used only for the sampling of new actions, whereas we consider updating existing actions.
-In [?], BetaZero was suggested to share information by learning offline policy and value networks over parametric beliefs.
-However, they consider discrete action spaces, they do not share information between action branches online, nor do they handle modeling inaccuracy.
-Our gradient results can be seen as similar to the policy gradient theorem [?] or to the deterministic policy gradient theorem [?], where we calculate gradients w.r.t.
-actions, rather than policy parameters.
-1.2 Contributions
-Our work is the first to consider an action-gradient of the value function based on transition likelihoods in continuous (PO)MDPs. Specifically, our main contributions are:
-  1. 1.
-    We formulate a novel search tree with recursive MIS estimators, for updating actions with consistent value estimates in continuous (PO)MDPs.
-We show efficient update equations for the self-normalized MIS heuristic.
-  2. 2.
-    We provide a novel derivation for action gradients of the value function in (PO)MDPs based on transition likelihoods.
-This gradient computation can be computed at a single action branch at a time, unlike previous MDP value gradients.
-The extension to POMDPs makes this the first derivation of POMDP value gradients with belief-dependent policies to the best of our knowledge.
-  3. 3.
-    We derive a novel algorithm Action Gradient MCTS (AGMCTS) that combines the MIS tree with action gradients.
-This algorithm is applicable to both MDPs and POMDPs.
-We show several algorithmic improvements and heuristics to make it more efficient.
-  4. 4.
-    We compare AGMCTS to PFT-DPW and POMCPOW in a simulated POMDP scenario at several dimensions and different noise levels in the domain knowledge.
-AGMCTS had the best overall performance on a per-simulation budget.
+Yet, computing the gradient of the action-value function is not trivial.
+Recently Lee et al.
+[9] explored gradient steps integrated in MCTS for continuous MDPs.
+Their analysis was limited to Lipschitz-continuous transition models and rewards, they did not analyze the value estimate drift that occurs from action updates, and most importantly their derivation does not extend to POMDPs trivially.
+Our gradient results can be seen as similar to the policy gradient theorem [10], however we calculate score gradients w.r.t.
+actions, rather than policy parameters, and in POMDPs.
+Policy gradients in POMDPs have been limited to memory-less policies [11] or to finite action spaces [12], yet we assume belief-dependent policies.
+Leurent and Maillard [13] considered information sharing in discrete MDPs for the specific case of repeating states.
+Novitsky et al.
+[14] introduced previous knowledge utilization in POMDPs via Multiple Importance Sampling (MIS) [15].
+A key difference is that we share information only between sibling action branches by introducing the novel MIS tree, which eliminates a costly search through a belief database.
+MIS in POMDP planning for sample reuse was also considered in [16], however in the context of calculating observation likelihoods.
+Finally, AlphaZero and BetaZero [17, 18] leverage value and policy networks to facilitate offline information sharing.
+However, these methods are currently constrained to parametric beliefs, discrete action spaces, and lack mechanisms for online information sharing between action branches.
+Our findings complement offline learning approaches, as MCTS-based online planning can act as a policy improvement operator [19].
+In this paper, we propose a novel online planning algorithm Action-Gradient Monte Carlo Tree Search (AGMCTS), that combines local gradient optimization iterations with global MCTS search for efficient optimization in continuous (PO)MDPs, for combining the advantages of both.
+Our contributions are theoretical, algorithmic and experimental.
+In section 3, we start by deriving a score action-gradient of the value function in MDPs via importance sampling, and we extend it to POMDPs.
+In particular, it is calculable in particle-belief MDPs and can be estimated via samples.
+We continue in Section 4 to describe the formulation of the novel MIS tree data structure.
+It allows updating actions branches during search time without losing information of previous action-value estimates.
+The MIS tree supports numerous gradient updates by making efficient and incremental computations.
+Our novel planning algorithm AGMCTS is presented in Section 5, alongside several approximation and heuristic techniques that improve performance and make the algorithm practical for online use.
+We finish our computational contributions in Section 6, showing that transition densities can be computed via the Area Formula for smooth generative models, allowing efficient computation of the action-gradient by using modern automatic differentiation.
+We conduct extensive experiments in Section 7, comparing AGMCTS to DPW (MDP), and to PFT-DPW and POMCPOW (POMDP) in several simulated scenarios.
+AGMCTS obtains the overall best performance on a per-simulation budget across most domains, and we discuss conclusions and future directions in Section 8.
+The proofs for all theorems and lemmas are reported in Appendix A, further details on AGMCTS are in Appendix B, and all details on the experimental setup and methods are found in Appendix C.
 2 Background
-MDPs.
 An MDP is the tuple $\langle\mathcal{S},\mathcal{A},p_{T},r,\gamma,L,b_{0}\rangle$.
-$\mathcal{S},\mathcal{A}$ are the state and action spaces.
+$\mathcal{S}\subseteq\mathbb{R}^{n_{s}},\mathcal{A}\subseteq\mathbb{R}^{n_{a}}$ are the state and action spaces.
 The transition model $p_{T}(s^{\prime}\mid s,a)$ models the probability of arriving at state $s^{\prime}$ when taking action $a\in\mathcal{A}$ at state $s\in\mathcal{S}$.
-The reward function $r(s,a,s^{\prime})\in\mathbb{R}$ gives the immediate reward of transitioning from state $s$ to $s^{\prime}$ by taking action $a$.
-We define the reward of $(s,a)$ as the expected reward over outcome states $r(s,a)\triangleq\mathbb{E}_{s^{\prime}\mid s,a}[r(s,a,s^{\prime})]$.
+The reward function $r(s,a,s^{\prime})\in\mathbb{R}$ gives the immediate reward of transitioning from state $s$ to $s^{\prime}$ by taking action $a$, and the expected reward $r(s,a)\triangleq\mathbb{E}_{s^{\prime}\mid s,a}[r(s,a,s^{\prime})]$.
 $\gamma\in(0,1]$ is the discount factor.
@@
 The initial state is drawn from the distribution $s_{0}\sim b_{0}$.
-A timed policy at time $i$, denoted as $\pi_{i}$, is a mapping $\pi_{i}\colon\mathcal{S}\to\mathcal{A}$.
-A policy $\pi$ is a collection of timed policies from the start time until the horizon: $\pi\triangleq(\pi_{t})_{t=1}^{t=L}$.
+A policy $\pi$ is a mapping from states to actions, and in general may be time-dependent.
+In our analysis we assume deterministic policies.
 The value function of a policy $\pi$ at time $t$ is the expected sum of discounted rewards until the horizon: $V_{t}^{\pi}(s_{t})\triangleq\mathbb{E}_{s_{t+1:L}\mid s_{t},\pi}[\sum_{i=t}^{L}\gamma^{i-t}r(s_{i},\pi_{i}(s_{i}),s_{i+1})]$.
-The action-value function is often used as an intermediate step in the estimation of the value function, and is defined as $Q_{t}^{\pi}(s,a)\triangleq\mathbb{E}_{s^{\prime}\mid s,a}[r(s,a,s^{\prime})+\gamma V_{t+1}^{\pi}(s^{\prime})]$, where state $s^{\prime}$ is at the proceeding time of $s$.
+The action-value function is defined as $Q_{t}^{\pi}(s,a)\triangleq\mathbb{E}_{s^{\prime}\mid s,a}[r(s,a,s^{\prime})+\gamma V_{t+1}^{\pi}(s^{\prime})]$, where state $s^{\prime}$ is at the proceeding time of $s$.
 The goal of planning is to compute a policy that maximizes the value function.
-POMDPs.
 A POMDP adds $\langle\mathcal{O},p_{O}\rangle$ to the MDP tuple.
-The observation space is $\mathcal{O}$, and the observation model $p_{O}(o\mid s)$ is the conditional probability of receiving an observation $o\in\mathcal{O}$ at $s\in\mathcal{S}$.
+The observation space is $\mathcal{O}\subseteq\mathbb{R}^{n_{o}}$, and the observation model $p_{O}(o\mid s)$ is the conditional probability of receiving an observation $o\in\mathcal{O}$ at $s\in\mathcal{S}$.
 A history at time $t$, $H_{t}\triangleq(b_{0},a_{0},o_{1},\dots,a_{t-1},o_{t})$, is defined as a sequence of the starting belief, followed by actions taken and observations received until $t$.
 Because of the partial observability, the agent has to maintain a probability distribution of the current state given past actions and observations, known as the belief.
-The belief space, denoted in $\mathcal{B}$, is the space of all probability distributions over $\mathcal{S}$.
-The belief at time $t$ is defined as the conditional distribution of the state $b_{t}(s_{t})\triangleq\mathbb{P}(s_{t}\mid H_{t})$, given the history.
+The belief at time $t$ is $b_{t}(s_{t})\triangleq\mathbb{P}(s_{t}\mid H_{t})$.
 We denote $H_{t}^{-}\triangleq(b_{0},a_{0},o_{1},\dots,a_{t-1})$ for the same history without the last measurement, and the propagated belief $b_{t}^{-}(s_{t})\triangleq\mathbb{P}(s_{t}\mid H_{t}^{-})$.
-It can be shown that optimal decision-making can be made given the belief, instead of considering the entire history [?].
-Since the belief fulfills the Markov property, a POMDP is an MDP on the belief space, commonly referred to as belief MDP.
-Policies, the value function and the action-value function, extend from MDPs to POMDPs by equivalent definitions on $b\in\mathcal{B}$ as states.
-As an example, $r(b,a)\triangleq\mathbb{E}_{b^{\prime}\mid b,a}[r(b,a,b^{\prime})]$.
-In the general case, the belief-reward function may be any function of beliefs, f.e.
-entropy, and this is called $\rho$-POMDP [?].
-Often, the belief reward has an underlying state-reward, in the form $r(b,a,b^{\prime})=\mathbb{E}_{s,s^{\prime}\mid b,a,b^{\prime}}[r(s,a,s^{\prime})]$.
-We refer to this case as a state-reward POMDP.
-Importance Sampling (IS).
-The Monte-Carlo (MC) estimator of an expectation, denoted with $\hat{\mathbb{E}}$ or f.e.
-$\hat{V}_{t}^{\pi}$, is its approximation with a finite number of samples.
-Importance Sampling (IS) is a technique in MC estimation where a proposal distribution $q$ is used to sample from instead of $p$.
-The resulting estimator for $f(x)$ is $\hat{f}=\frac{1}{N}\sum_{i=1}^{N}\frac{p(x^{i})}{q(x^{i})}f(x^{i})$.
-This estimator is consistent if $q(x)=0\Rightarrow p(x)=0$.
-With a proper choice of $q$ the variance of the estimator can be reduced, or we can calculate estimates for a target $p$ which is difficult to sample directly.
-Commonly the sampled weights $w_{i}=p(x^{i})/q(x^{i})$ are normalized to sum to $1$, giving the self-normalized IS estimator [?, 1.3.2].
-Particle Beliefs.
-Often in continuous settings we cannot maintain an exact belief.
-Hence, particle filters are used instead to represent a belief non parametrically [?].
+It can be shown that optimal decision-making can be made given the belief, instead of considering the entire history [20].
+Since the belief fulfills the Markov property, a POMDP is an MDP on the belief space, referred to as belief MDP.
+Policies, the value function and the action-value function, extend from MDPs to POMDPs by equivalent definitions on beliefs as states.
+Importance Sampling (IS) is a technique in Monte-Carlo (MC) estimation where a proposal distribution $q$ is used to generate samples of the target $p$, and the resulting estimator for $g(x)=\mathbb{E}_{x\sim p}[f(x)]$ is $\hat{g}_{\text{IS}}=N^{-1}\sum_{i=1}^{N}f(x^{i})p(x^{i})/q(x^{i})$.
+$\hat{g}$ is unbiased if $q(x)=0$ implies $p(x)=0$.
+The choice of $q$ affects the variance of the estimator, and IS can be used to compute estimates even if samples cannot be generated from $p$.
+Commonly the sampled weights $w_{i}=p(x^{i})/q(x^{i})$ are normalized to sum to $1$, giving the self-normalized IS estimator [21, 1.3.2].
+It is biased but under weak assumptions is asymptotically consistent.
+MIS [15] is a method to combine the advantages of several sampling methods.
+The MIS estimate is a weighted average of $n$ IS estimators $\hat{g}_{\text{MIS}}=\sum_{i=1}^{n}n_{i}^{-1}\sum_{j=1}^{n_{i}}w_{i}(x^{i,j})f(x^{i,j})/q_{i}(x^{i,j})$, with $n_{i}$ samples each.
+If the weights $w_{i}$ satisfy requirements: (MIS-I) $\sum_{i=1}^{n}w_{i}(x)=1$ whenever $f(x)\neq 0$; (MIS-II) $w_{i}(x)=0$ whenever $q_{i}(x)=0$, then $\hat{g}_{\text{MIS}}$ is unbiased.
+The strength of MIS is that $w_{i}$ may depend on the samples $x^{i,j}$.
+Many heuristic weighting strategies can be employed, however the balance heuristic $w_{i}(x)=n_{i}q_{i}(x)/\sum_{k}n_{k}q_{k}(x)$, that weights according to a mixture of all $q_{i}$, has its variance provably bounded from the optimal weighting strategy.
+Particle filters are often used to represent a belief non-parametrically [22].
 We denote particle beliefs with $\bar{b}$.
-The particle belief of $J$ particles is represented by a set of state-weight pairs $\bar{b}=\{(s^{j},w^{j})\}_{j=1}^{J}$, and is defined as the discrete distribution $\mathbb{P}(s\mid\bar{b})={\sum_{j=1}^{J}{w^{j}\cdotp\delta(s-s^{j})}}/{\sum_{j=1}^{J}{w^{j}}}$.
-The basic particle filtering algorithm, Sequential Importance Sampling (SIS) [?, 1.3.2], recursively updates and normalizes the importance weights of the particles based on the observation likelihood.
-Sequential Importance Resampling (SIR) adds resampling of state particles after the weight updates, to avoid weight degeneracy [?].
-The particle belief MDP (PB-MDP) was considered in [?], where the belief space is explicitly the space of ordered particle beliefs $\bar{b}$ of $J$ particles.
-An ordered particle belief is the ordered set $\bar{b}=((s^{j},w^{j}))_{j=1}^{J}$.
-Multiple Importance Sampling (MIS).
-MIS [?] is a method to combine the advantages of several sampling methods.
-It is defined as $F=\sum_{i=1}^{n}\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}w_{i}(x^{i,j})\frac{f(x^{i,j})}{q_{i}(x^{i,j})}$, which is a weighted average of $n$ IS estimators.
-If the weights $w_{i}$ satisfy $\sum_{i=1}^{n}w_{i}(x)=1$ whenever $f(x)\neq 0$, and $w_{i}(x)=0$ whenever $q_{i}(x)=0$, then $F$ is unbiased.
-The choice of $w_{i}$ may depend on the samples $x^{i,j}$.
-The weighting strategy $w_{i}(x)=\frac{n_{i}q_{i}(x)}{\sum_{k}n_{k}q_{k}(x)}$, named the balance heuristic, has its variance provably bounded from the optimal weighting strategy.
-Monte Carlo Tree Search (MCTS).
-MCTS is an algorithm used to quickly explore large state spaces [?].
+The particle belief of $J$ particles is represented by a set of state-weight pairs $\bar{b}=\{(s^{j},\lambda^{j})\}_{j=1}^{J}$, and is defined as the discrete distribution $\mathbb{P}(s\mid\bar{b})={\sum_{j=1}^{J}{\lambda^{j}\cdotp\delta(s-s^{j})}}/{\sum_{j=1}^{J}{\lambda^{j}}}$.
+For computational simplicity, we assume throughout the paper that the bootstrap filter is used to update particle beliefs [21, 1.3.2].
+This filter advances sampled particles using the transition model, reweights them based on the observation likelihood, and resamples to avoid weight degeneracy [23].
+Additionally, we assume that the belief space is explicitly that of ordered particle beliefs, where beliefs are ordered sets $\bar{b}=((s^{j},\lambda^{j}))_{j=1}^{J}$.
+Our results and in particular Theorem 1 can be generalized to more sophisticated particle filtering methods, and we leave that for future work.
+Lastly, MCTS is an algorithm used to quickly explore large state spaces [24].
 It iteratively repeats four steps to build a search tree that approximates the action-values, using a best-first strategy: (1) Selection Starting from the root node, descend recursively until a node to which children can be added is found; (2) Expansion A new node is added as a child, according to action expansion strategy; (3) Simulation A simulation is run from the new node according to the rollout (default) policy; (4) Backpropagation The simulation result is backpropagated through the selected nodes to update the action-values.
-Often UCT is used at selection to balance between exploration of new actions and exploitation of promising ones [?].
-Double Progressive Widening (DPW) [?] is a technique to limit the branching factor from being infinite in continuous settings.
-The number of children of a node is artificially limited to $kN^{\alpha}$ for $N$ visitations, for hyperparameters $k$ and $\alpha$.
-The progressive widening is double due to being applied both to the branching of actions, and branching of states.
-Stochastic Gradient Optimization (SGO).
-SGO is a widely used technique for efficient optimization of differentiable functions [?].
-Gradient optimizers tremendously rose in popularity to their applications in deep neural networks.
-Adam [?] is a widely used algorithm, with heuristically adaptive step size based on first and second moment estimates of the gradient.
-Adam has been shown to perform well with noisy gradient estimates, and requires minimal tuning to different problems.
-3 Methodology
-3.1 Action-Values with Importance Sampling
-Figure 2: Recursive definitions of the MIS tree, for the particle belief MDP (PB-MDP) case, in which states are particle beliefs.
-State nodes are circles and action nodes are squares.
-The state $\bar{b}$ has a value estimate $\hat{V}(\bar{b})$ and visitation count $n(\bar{b})$.
-The action node $(\bar{b},a)$ has an immediate reward estimate $\hat{r}(\bar{b},a)$, a future value estimate $\hat{V}_{f}(\bar{b},a)$, and visitation count $n(\bar{b},a)$.
-The action-value estimate is defined via $\hat{Q}(\bar{b},a)\triangleq\hat{r}(\bar{b},a)+\gamma\hat{V}_{f}(\bar{b},a)$.
-During action update we update $(\bar{b},a)$ to $(\bar{b},a^{\prime})$.
-We highlight that in the PB-MDP, we also store the propagated belief $\bar{b}^{-}$.
-The child state node has all properties as $\bar{b}$, and stores the proposal action $a_{prop}(\bar{b}^{-})$.
-During action update (11) we do $\omega^{\prime}={p_{T}(\bar{b}^{-}\mid\bar{b},a^{\prime})}({p_{T}(\bar{b}^{-}\mid\bar{b},a_{prop}(\bar{b}^{-}))})^{-1}$.
-In this example, $a_{prop}(\bar{b}^{-})=a$.
-In an MDP setting, when updating action $a$ to $a^{\prime}$, we would like to estimate the new action-value $Q_{t}^{\pi}(s,a^{\prime})$ by reusing the information that was used for estimating $Q_{t}^{\pi}(s,a)$.
-Otherwise, updating actions would require us to theoretically recompute the entire subtree, which would render the algorithm very inefficient.
-Since this process might occur several times for each action branch, in the general case each child state may have been generated from a different proposal action.
-Hence, we have to account for an MIS scenario in order to estimate $Q_{t}^{\pi}(s,a^{\prime})$.
-Assuming that $p_{T}(s^{\prime}\mid s,a^{\prime})$ is a non-strict subset of the support of $p_{T}(s^{\prime}\mid s,a)$, we can formally write
-$\textstyle Q_{t}^{\pi}(s,a^{\prime})=\mathbb{E}_{s^{\prime}\mid s,a}[\frac{p_{T}(s^{\prime}\mid s,a^{\prime})}{p_{T}(s^{\prime}\mid s,a)}(r(s,a^{\prime},s^{\prime})+\gamma V_{t+1}^{\pi}(s^{\prime}))].$ (1)
-For POMDPs, the posterior belief transition likelihood is generally intractable.
-However, by using the law of total expectation, we introduce a dependence of the posterior belief $b^{\prime}$ on the propagated $b^{-}$.
-This makes only $b^{-}$ directly dependent on $a$, which allows us to introduce an importance ratio over the propagated belief rather than posterior:
-$\textstyle Q_{t}^{\pi}(b,a^{\prime})=\mathbb{E}_{b^{-}\mid b,a}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}\textstyle[\frac{p_{T}(b^{-}\mid b,a^{\prime})}{p_{T}(b^{-}\mid b,a)}$
-$(r(b,a^{\prime},b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime}))].$ (2)
-This is known as the propagated belief trick [?].
-We can see that (1) and (2) are analogous, except for the importance ratio being between propagated beliefs in the POMDP case.
-The propagated belief likelihood is rather simple to calculate for appropriate choice of the belief space.
-Lemma 1.
-Let $\bar{b}^{-}=((s^{-,j},w^{j}))_{j=1}^{J}$ be generated from $\bar{b}=((s^{j},w^{j}))_{j=1}^{J}$ by SIS or SIR.
-Hence, the $j$th state particle was generated via $s^{-,j}\sim p_{T}(\cdot\mid s^{j},a)$, and:
-$\textstyle p_{T}(\bar{b}^{-}\mid\bar{b},a)=\prod_{i=1}^{J}p_{T}(s^{-,j}\mid s^{j},a).$ (3)
-We refer to the appendix for all proofs
-3.2 Action Adaptive MIS Trees
-We now present our definitions of the novel MIS tree, a recursive tree structure with MIS value estimates, that allows sharing value information between sibling actions.
-Similarly to belief trees, the tree has levels alternating with states and action nodes, starting from the root state node.
-Yet in contrast, the MIS tree supports replacing an action node with a new action, and updating the value estimates of the new action and its parent state node based on previous samples.
-We later introduce action updates via gradient-based optimization.
-For the following discussion, we present notations for recursively defining estimators at a state node $s$ in the MIS tree, or some action node $(s,a)$.
-The generalization to POMDPs is done by taking ordered particle belief states $\bar{b}$, and the following definitions are analogous, illustrated in Figure 2.
-Denote the children actions of $s$ as $C(s)=\{a^{i}\}_{i=1}^{\lvert{C(s)}\rvert}$, and the children states of $(s,a)$ as $C(s,a)=\{s^{\prime,i}\}_{i=1}^{M}$, where $M\triangleq\lvert{C(s,a)}\rvert$.
-The visitation counts of state and action nodes are denoted $n(s)$ and $n(s,a)$ respectively.
-We denote the proposal action $a_{prop}(s)$, the action that was used to sample $s$, as possibly $a_{prop}(s^{\prime})\neq a$ for $s^{\prime}\in C(s,a)$.
-We assign each action node $(s,a)$ an action-value estimate $\hat{Q}(s,a)$ that estimates $Q_{t}^{\pi}(s,a)$, assuming $s$ is a state at time $t$.
-Compared to traditional MCTS algorithms, we also estimate the value function at each state node.
-We define it as a weighted mean:
-$\textstyle\hat{V}(s)\triangleq\sum_{C(s)}\frac{n(s,a^{i})}{n(s)}\hat{Q}(s,a^{i}).$ (4)
-The following two relationships between visitation counts are maintained in MCTS:
-$\textstyle n(s)=\sum_{C(s)}n(s,a^{i}),\;\textstyle n(s,a)=\sum_{C(s,a)}n(s^{\prime,i})_{+1},$ (5)
-where $n(s^{\prime,i})_{+1}\triangleq n(s^{\prime,i})+1$.
-The asymmetry is the result of initializing state nodes to $n(s)=0$ when assigned a rollout value, whereas action nodes are initialized to $n(s,a)=1$.
-This allows value estimates to be updated completely when they are visited again, and for their parent’s action-value estimate to remember the initial rollout estimate.
-We split $\hat{Q}(s,a)$ into two estimators: (1) Immediate reward estimate $\hat{r}(s,a)$, for $\mathbb{E}_{s^{\prime}\mid s,a}[r(s,a,s^{\prime})]$; (2) Future value estimate $\hat{V}_{f}(s,a)$ for $\mathbb{E}_{s^{\prime}\mid s,a}[V^{\pi}(s^{\prime})]$ of the proceeding time, such that if $s$ is a state at time $t$, then $\hat{V}_{f}(s,a)$ estimates the expected $V_{t+1}^{\pi}(s^{\prime})$.
-Together they define $\hat{Q}(s,a)\triangleq\hat{r}(s,a)+\gamma\hat{V}_{f}(s,a)$.
-Generally in MIS, each proposal distribution $q_{i}$ has $n_{i}$ samples.
-In our situation, we assume that each of $C(s,a)$ may have been sampled from a different action, has its own visitation counter and an associated single value estimate $V(s^{\prime,i})$.
-Therefore, we adapt the most naive unbiased weighting, that is to give equal weight to all samples, and weight the posterior states by the visitation counts.
-Hence, the naive MIS estimator for the future value is
-$\textstyle\hat{V}_{f}(s,a)=(n(s,a))^{-1}\sum_{C(s,a)}\omega^{i}n(s^{\prime,i})_{+1}\hat{V}(s^{\prime,i}),$ (6)
-where $\omega^{i}={p_{T}(s^{\prime,i}\mid s,a)}({p_{T}(s^{\prime,i}\mid s,a_{prop}(s^{\prime,i}))})^{-1}$.
-The naive MIS estimator for $\hat{r}(s,a)$ is identical, with $\hat{V}(s^{\prime,i})$ replaced by $r(s,a,s^{\prime,i})$.
-In practice, this estimator has very bad performance.
-On the other hand, the balance heuristic requires $O(M)$ time to update [?], which can be quite expensive.
-We found a simple solution which is to self-normalize the weights.
-In regular IS, it is known to reduce the variance of the estimator at the expense of introducing bias, and asymptotically converges to the true estimate under weak assumptions [?, 1.3.2].
-We define self-normalized MIS (SN-MIS) by
-$\textstyle\eta\triangleq\sum_{C(s,a)}\omega^{i}n(s^{\prime,i})_{+1},$ (7)
-$\textstyle\hat{V}_{f}(s,a)=\eta^{-1}\sum_{C(s,a)}\omega^{i}n(s^{\prime,i})_{+1}\hat{V}(s^{\prime,i}),$ (8)
-and again the SN-MIS estimator for $\hat{r}(s,a)$ is defined by replacing $\hat{V}(s^{\prime,i})$ with $r(s,a,s^{\prime,i})$.
-3.3 Updating MIS Trees in MCTS
-We now discuss how to update the SN-MIS estimators in the steps of the MCTS routine. We highlight that for any other MIS weighting heuristic, a similar analysis would follow.
-Currently, we do not address selection, and we assume the UCT criterion is not altered.
-Simulation is unaffected, as it is determined only by the rollout policy.
-Hence, we analyze the different subcases of backpropagation and expansion steps.
-We add a new step to MCTS named action update, where we update action node $(s,a)$ to $(s,a^{\prime})$ with the new action $a^{\prime}$, and we show its update equations.
-In general, action update can be applied at every point of MCTS.
-For the recursive analysis, we assume backpropagated values for the children nodes are already updated, and at each step we only update an action node $(s,a)$ and its parent node $s$.
-The updated values are denoted with a prime, and we assume that we have access to the values before the update.
-The transition likelihood terms $w_{i}$ are usually the most expensive to calculate, hence it is important to cache them, and recalculate the numerator of $w_{i}$ only when necessary.
-The derivations of the update equations, alongside numerically stable forms with log-likelihoods are given in the appendix.
-Action Backpropagation.
-Let $s^{\prime,i}$ be an updated node.
-Hence, we updated $n(s^{\prime,i})$ to $n^{\prime}(s^{\prime,i})$ and $\hat{V}(s^{\prime,i})$ to $\hat{V}^{\prime}(s^{\prime,i})$.
-We update $n(s,a)$ by (5) and perform
-$\textstyle\eta^{\prime}=\eta+\omega^{i}(n{{}^{\prime}}(s^{\prime,i})-n(s^{\prime,i})),$ (9)
-$\textstyle\hat{V}_{f}^{\prime}(s,a)=(\eta^{\prime})^{-1}(\eta\hat{V}_{f}(s,a)+$
-$\textstyle\omega^{i}(n{{}^{\prime}}(s^{\prime,i})\hat{V}^{\prime}(s^{\prime,i})-n(s^{\prime,i})\hat{V}(s^{\prime,i})).$ (10)
-State Expansion.
-This case is the same as action backpropagation, where the updated index is a new summand at index $i=M+1$. The same equations apply with $n(s^{\prime,i})=0$.
-Action Update.
-Let $a^{\prime}$ be the new action. Here, we have to calculate $\omega^{i}$ for all $s^{\prime,i}\in C(s,a)$ in $O(M)$ time. We update
-$\textstyle(\omega^{i})^{\prime}=p_{T}(s^{\prime,i}\mid s,a^{\prime})(p_{T}(s^{\prime,i}\mid s,a_{prop}(s^{\prime,i})))^{-1},$ (11)
-and recalculate (7) and (8) with the new ($\omega^{i})^{\prime}$, which are also $O(M)$ time operations.
-For the immediate reward, new rewards must be computed for the new triplets $(s,a^{\prime},s^{\prime,i})$, and the update is analogous.
-Terminal State Backpropagation.
-Let $s$ be a terminal node. Its value estimate is based only on rollouts. Hence, for a new rollout value $v^{\prime}$, we perform a running average update
-$\textstyle\hat{V}^{\prime}(s)=\hat{V}(s)+\frac{n^{\prime}(s)-n(s)}{n^{\prime}(s)_{+1}}(v^{\prime}-\hat{V}(s)),$ (12)
-where $n^{\prime}(s)$ is $n(s)$ plus the number of rollouts (usually 1).
-Non-Terminal State Backpropagation.
-Let $(s,a^{i})$ be an updated node. Hence, we updated $n(s,a^{i})$ to $n^{\prime}(s,a^{i})$ and $\hat{Q}(s,a^{i})$ to $\hat{Q}^{\prime}(s,a^{i})$. We update $n^{\prime}(s)$ by (5) and perform
-$\textstyle\hat{V}^{\prime}(s)=(n^{\prime}(s))^{-1}(n(s)\hat{V}(s)+n^{\prime}(s,a^{i})\hat{Q}^{\prime}(s,a^{i})$
-$\textstyle{}-n(s,a^{i})\hat{Q}(s,a^{i})).$ (13)
-3.4 MDP and POMDP Value Gradient
-In the previous subsections we’ve assumed that the updated action $a^{\prime}$ is arbitrary.
-We now present a method to compute an estimate of $\nabla_{a}Q(s,a)$, in order to compute a gradient-based update.
-Our derivation of $\nabla_{a}Q(s,a)$ is based on the gradient of the transition likelihood, and there are two main advantages to this approach.
-First, the non-recursive computation can be computed separately for each action branch in the MIS tree, which greatly reduces computational complexity.
-Second, which is our key observation, is that this derivation extends to POMDPs as well via the propagated belief trick, and has a computable form for the PB-MDP.
+Often UCT is used at selection to balance between exploration of new actions and exploitation of promising ones [25].
+Double Progressive Widening (DPW) [1] is a technique to limit the branching factor from being infinite in continuous settings.
+The number of children of a node is artificially limited to $kN^{\alpha}$ for $N$ visitations, for fixed hyperparameters $k$ and $10<\alpha<$ 1.
+3 Action-Value Gradients via Importance Sampling
+In a tree search setting, when updating action $a$ to $a^{\prime}$, we would like to estimate the new action-value $Q_{t}^{\pi}(s,a^{\prime})$ by reusing the information that was used for estimating $Q_{t}^{\pi}(s,a)$.
+Otherwise, updating actions would require us to theoretically recompute the entire subtree, which would render a practical algorithm very inefficient.
+Since this process might occur several times for each action branch, generally each child state may have been generated from a different proposal action.
+Hence, we have to account for an MIS scenario in order to calculate $Q$.
+For simplicity of equations, we present the IS-accounted $Q$ function with proposal $q(s^{\prime}\mid s,a)$ of a single action, and we later combine several of these estimates via MIS in Section 4.
+The MDP and POMDP action-value functions satisfy
+$\textstyle Q_{t}^{\pi}(s,a^{\prime})=\mathbb{E}_{s^{\prime}\mid q}[\omega_{p_{T}/q}(s^{\prime})(r(s,a^{\prime},s^{\prime})+\gamma V_{t+1}^{\pi}(s^{\prime}))],$ (1)
+$\textstyle Q_{t}^{\pi}(b,a^{\prime})=\mathbb{E}_{b^{-},o,b^{\prime}\mid q}\textstyle[\omega_{p_{T}/q}(b^{-})(r(b,a^{\prime},b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime}))],$ (2)
+where $\omega_{p_{T}/q}(s^{\prime})\triangleq p_{T}(s^{\prime}\mid s,a^{\prime})/q(s^{\prime}\mid s,a)$, and it is defined similarly for the POMDP case.
+For the importance weight to be well-defined, we require that $q(s^{\prime}\mid s,a)=0\Rightarrow p_{T}(s^{\prime}\mid s,a)=0$.
+A sufficient conditions is that all actions share the same support: $\text{Supp}[p_{T}(\cdot\mid s,a)]$ is independent of $a$.
+This is typical in continuous domains with full-support noise models, for instance additive Gaussian noise with action-dependent mean but fixed covariance, where the support remains $\mathbb{R}^{n_{s}}$ for every action.
+However, as Theorem 2 later shows, this is not strictly necessary.
+In POMDPs, the posterior belief transition likelihood is generally intractable.
+The trick is that by using the law of total expectation, we introduce a dependence of the posterior belief $b^{\prime}$ on the propagated $b^{-}$.
+The distribution of $b^{-}$ is non-degenerate in non-parametric settings.
+This makes only $b^{-}$ directly dependent on $a$, allowing to introduce an importance ratio over the propagated belief rather than posterior.
+This is known as the propagated belief trick [14], resulting in analogous equations.
+The importance of this derivation is that it allows us to derive a practical action-gradient for POMDPs via particle-belief approximations.
+We now introduce our main results of score action-gradients in MDPs and POMDPs.
+To the best of our knowledge, it is the first gradient theorem of the action-value function in POMDPs with belief-dependent policies.
 Theorem 1.
-For well-behaved $p_{T}$, $r$, the MDP action gradient of the action-value function is given by the expectation
-$\textstyle\nabla_{a^{\prime}}Q_{t}^{\pi}(s,a^{\prime})\textstyle=\mathbb{E}_{s^{\prime}\mid s,a}[\frac{p_{T}(s^{\prime}\mid s,a^{\prime})}{p_{T}(s^{\prime}\mid s,a)}(\nabla_{a^{\prime}}\log p_{T}(s^{\prime}\mid s,a^{\prime})$
-$\textstyle\cdot(r(s,a^{\prime},s^{\prime})+\gamma V_{t+1}^{\pi}(s^{\prime}))+\nabla_{a^{\prime}}r(s,a^{\prime},s^{\prime}))].$ (14)
-Additionally, for well-behaved $p_{T}$, $p_{O}$, $r$, the POMDP action gradient of the action-value function can be computed by replacing posterior belief likelihood terms with propagated belief likelihood terms, i.e.
-$b_{t+1}^{-}$ instead of $b_{t+1}$ in all $p_{T}$ terms.
+Assume $q=0\Rightarrow p_{T}=0$ and bounded $r$, $V$.
+If $\lVert{\nabla_{a}r}\rVert$, $\lVert{\nabla_{a}p_{T}}\rVert$ have bounding integrable envelopes w.r.t.
+the MDP/POMDP measures, then the following action-gradient identity results from the dominated-convergence form of Leibniz integral rule [26, Theorem 2.27]:
+$\textstyle\nabla_{a^{\prime}}Q_{t}^{\pi}(s,a^{\prime})\textstyle=\mathbb{E}_{s^{\prime}\mid q}[\omega_{p_{T}/q}(s^{\prime})[\nabla_{a^{\prime}}\log p_{T}(s^{\prime}\mid s,a^{\prime})(r(s,a^{\prime},s^{\prime})+\gamma V_{t+1}^{\pi}(s^{\prime}))+\nabla_{a^{\prime}}r]],$ (3)
+$\textstyle\nabla_{a^{\prime}}Q_{t}^{\pi}(b,a^{\prime})=\mathbb{E}_{b^{-},o,b^{\prime}\mid q}\textstyle[\omega_{p_{T}/q}(b^{-})[\nabla_{a^{\prime}}\log p_{T}(b^{-}\mid b,a^{\prime})(r(b,a^{\prime},b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime}))+\nabla_{a^{\prime}}r]].$ (4)
+The proof is similar to the policy gradient theorem [10].
+The single-envelope assumption is deliberately weak.
+Consequently, $r$, $p_{T}$, and $p_{O}$ may be piece-wise differentiable with countably many discontinuities, which is useful when modeling saturated actuators, contact dynamics, or clipped rewards.
 The gradient is written such that states $s^{\prime}$ could have been sampled from an action $a$ different from $a^{\prime}$.
@@
 it.
-Theorem 1 is general, and applies to MDPs, state-reward POMDPs, and $\rho$-POMDPs.
-The following Theorem 2 modifies the gradient for state-reward POMDPs.
-The gradient for the immediate reward is based on single state samples rather than sampling full beliefs.
+There are two main advantages for an MCTS algorithm to compute score gradients compared to recursive gradient estimates similar to Lee et al.
+[9].
+First, the non-recursive computation can be computed separately for each action branch in the MIS tree, which greatly reduces computational complexity.
+Second, which is our key observation, is that this derivation extends to POMDPs as well via the propagated belief trick.
+The propagated belief transition density can be simple to calculate for appropriate choices of the belief space.
+For the specific case of ordered particle beliefs updated with the bootstrap filter, we get the following result.
+Lemma 1.
+Let $\bar{b}^{-}=((s^{-,j},\lambda^{j}))_{j=1}^{J}$ be generated from $\bar{b}=((s^{j},\lambda^{j}))_{j=1}^{J}$ by the bootstrap filter.
+Hence, the $j$’th state particle was generated via $s^{-,j}\sim p_{T}(\cdot\mid s^{j},a)$, and it holds that
+$\textstyle p_{T}(\bar{b}^{-}\mid\bar{b},a)=\prod_{i=1}^{J}p_{T}(s^{-,j}\mid s^{j},a),$ (5)
+$\textstyle\nabla_{a}\log p_{T}(\bar{b}^{-}\mid\bar{b},a)=\sum_{i=1}^{J}\nabla_{a}\log p_{T}(s^{-,j}\mid s^{j},a).$ (6)
+4 Action-Adaptive MCTS via MIS Trees
+Figure 1: The different update operations in the MIS tree from left to right: (1) Action update, a novel tree operation, in which a new action is connected to previous children states with updated action-value; (2) Action backpropagation; (3) State backpropagation.
+We now present our definitions of the novel MIS tree, a recursive tree structure with MIS value estimates, that allows sharing value information between sibling actions.
+Similarly to a belief tree, it has levels alternating with states and action nodes, starting from the root state node.
+Yet in contrast, the MIS tree supports replacing an action node with a new action, while keeping integrity of the value estimate based on existing samples.
+We call this new operation action update.
+We present notations for recursively defining the MIS tree via a state node $s$ or an action node $(s,a)$.
+The generalization to POMDPs is done by taking ordered particle belief states $\bar{b}$ instead of states $s$.
+Denote the children actions of $s$ as $C(s)=\{a^{i}\}_{i}$, and the children states of $(s,a)$ as $C(s,a)=\{s^{\prime,i}\}_{i}$.
+The visitation counts of state and action nodes are denoted $n(s)$ and $n(s,a)$ respectively.
+We denote the proposal action $a_{prop}(s)$, the action that was used to sample $s$, which is the highlight of the MIS tree as possibly $a_{prop}(s^{\prime})\neq a$ for $s^{\prime}\in C(s,a)$.
+We assign each action node $(s,a)$ an action-value estimate $\hat{Q}(s,a)$.
+Compared to traditional MCTS algorithms, we also estimate the value function at each state node.
+We define it as a weighted mean, based on invariants that the visitation counts satisfy during MCTS:
+$\textstyle\hat{V}(s)\triangleq n(s)^{-1}\sum_{a\in C(s)}n(s,a)\hat{Q}(s,a),$ (7)
+$\textstyle n(s)=\sum_{a\in C(s)}n(s,a),\quad\textstyle n(s,a)=\sum_{s^{\prime}\in C(s,a)}n(s^{\prime})_{+1},$ (8)
+[…]
+$\textstyle\hat{V}_{f}(s,a)\triangleq\sum_{s^{\prime}\in C(s,a)}w_{i}\hat{V}(s^{\prime})\cdot p_{T}(s^{\prime}\mid s,a)/p_{T}(s^{\prime}\mid s,a_{prop}(s^{\prime})),$ (9)
+The MIS estimator for $\hat{r}(s,a)$ is obtained by replacing $\hat{V}(s^{\prime})$ with $r(s,a,s^{\prime})$.
+Theorem 2.
+If the weighting strategy $w_{i}$ satisfies assumption (MIS-I) and (MIS-II), then the MIS tree yields unbiased estimates of the value and action-value functions for a given tree structure.
+Choosing a good weighting strategy is therefore important.
+Novitsky et al.
+[14] showed that the weights of the balance heuristic can be recursively updated in $O(\lvert{C(s,a)}\rvert)$, however even that could get expensive when performing many updates, as updates require re-evaluating transition probabilities and the reward function.
+We employ a simple solution which is to self-normalize the weights, and we define self-normalized MIS (SN-MIS) by
+$\textstyle\hat{V}_{f}(s,a)=\eta(s,a)^{-1}\sum_{s^{\prime}\in C(s,a)}\omega_{a}(s^{\prime})n(s^{\prime})_{+1}\hat{V}(s^{\prime}),\quad\eta(s,a)\triangleq\sum_{s^{\prime}\in C(s,a)}\omega_{a}(s^{\prime})n(s^{\prime})_{+1},$ (10)
+where $\omega_{a}(s^{\prime})\triangleq p_{T}(s^{\prime}\mid s,a)/p_{T}(s^{\prime}\mid s,a_{prop}(s^{\prime}))$, and similarly the SN-MIS estimator for $\hat{r}(s,a)$ is defined by replacing $\hat{V}(s^{\prime})$ with $r(s,a,s^{\prime})$.
+This estimator produces biased results, however the weights by construction are bounded in $[0,1]$ and sum to $1$ which limits the variance.
+It currently remains unproven whether SN-MIS is consistent.
+The major advantage of SN-MIS is that its update operations can be computed in $O(1)$ time, except for action update which is $O(\lvert{C(s,a)}\rvert)$.
+We now show the update equations of SN-MIS for different operations of the MIS tree, depicted in Figure 1.
+We highlight that for any other MIS weighting, a similar analysis would follow, and similar recursive updates may be obtained.
+Currently, we do not address selection, and we assume the UCT criterion is not altered.
+Simulation is unaffected, as it is determined only by the rollout policy.
+Hence, we analyze the different subcases of expansion and backpropagation steps.
+We add a new step to MCTS named action update, where we update action node $(s,a)$ to $(s,a^{\prime})$ with the new action $a^{\prime}$, which in general may be applied at any stage of the MCTS routine.
+The full derivations of the update equations, alongside numerically stable forms with log-likelihoods are given in Appendix A.
+Action update. Let $a^{\prime}$ be the new action. Hence, we calculate $\omega_{a^{\prime}}(s^{\prime})$ for all $s^{\prime}\in C(s,a)$ via:
+$\textstyle\omega_{a^{\prime}}(s^{\prime})=p_{T}(s^{\prime}\mid s,a^{\prime})/p_{T}(s^{\prime}\mid s,a_{prop}(s^{\prime})),$ (11)
+and recalculate (10) with the new $\omega_{a^{\prime}}(s^{\prime})$ also in $O(\lvert{C(s,a)}\rvert)$ time.
+For the immediate reward estimate, we evaluate $r(s,a^{\prime},s^{\prime})$ for the new triplets, and the update is analogous.
+In POMDPs, the explicit computation of the new transition likelihood $p_{T}(s^{\prime}\mid s,a^{\prime})$ via (5) requires $J$ evaluations of the transition density.
+Being a highly non-linear function, estimates based on a subset of particles can result in very large biases.
+However, for small changes in $a$ we can make the linear approximation
+$\textstyle\log\omega_{a^{\prime}}(s^{\prime})-\log\omega_{a}(s^{\prime})\approx(\nabla_{a}\log\omega_{a}(s^{\prime}))^{T}\delta a=(\nabla_{a}\log p_{T}(\bar{b}^{-}\mid\bar{b},a))^{T}\delta a,$ (12)
+and due to (6) being a sum, it is candidate for unbiased MC estimation with a subset of samples.
+This linearization holds when $\delta a$ is sufficiently small, a good practice in gradient-based optimization, and can be computed efficiently by caching $\nabla_{a}\log p_{T}(\bar{b}^{-}\mid\bar{b},a)$ from computing (4).
+Action backpropagation.
+Let $s^{\prime}$ be an updated node.
+Hence, we updated $n(s^{\prime})$ to $n^{\prime}(s^{\prime})$ and $\hat{V}(s^{\prime})$ to $\hat{V}^{\prime}(s^{\prime})$.
+We update $n(s,a)$ by (8) and perform
+$\textstyle\eta^{\prime}(s,a)=\eta(s,a)+\omega_{a}(s^{\prime})(n{{}^{\prime}}(s^{\prime})-n(s^{\prime})),$ (13)
+$\textstyle\hat{V}_{f}^{\prime}(s,a)=(\eta^{\prime}(s,a))^{-1}(\eta(s,a)\hat{V}_{f}(s,a)+\textstyle\omega_{a}(s^{\prime})(n{{}^{\prime}}(s^{\prime})\hat{V}^{\prime}(s^{\prime})-n(s^{\prime})\hat{V}(s^{\prime})).$ (14)
+Assuming a general $n^{\prime}(s^{\prime})$ supports cases where we delete branches in the tree. In case state expansion occurred, the same equations apply by initializing $n(s^{\prime})=0$.
+State backpropagation.
+Let $s$ be a state node.
+If $s$ is at the maximum tree depth, its value estimate is based only on rollouts.
+We perform a running average update with the new rollout value $v^{\prime}$,
+$\textstyle\hat{V}^{\prime}(s)=\hat{V}(s)+(n^{\prime}(s)_{+1})^{-1}(n^{\prime}(s)-n(s))(v^{\prime}-\hat{V}(s)),$ (15)
+where $n^{\prime}(s)$ is $n(s)$ plus the number of rollouts (usually 1, but we keep it general).
+If $s$ is not a terminal state, its value is updated based on the new action-value estimate of its updated child.
+Let $(s,a)$ be its updated child, with updated $n^{\prime}(s,a)$ and $\hat{Q}^{\prime}(s,a)$.
+We update $n^{\prime}(s)$ by (8) and perform
+$\textstyle\hat{V}^{\prime}(s)=(n^{\prime}(s))^{-1}(n(s)\hat{V}(s)+n^{\prime}(s,a)\hat{Q}^{\prime}(s,a)-n(s,a)\hat{Q}(s,a)).$ (16)
+5 Action-Gradient Monte Carlo Tree Search (AGMCTS)
+Algorithm 1 AGMCTS
+procedure S IMULATE$(s,d)$
+1:  if $d=0$ then
+2:     $v\leftarrow$
+R
+OLLOUT$(s)$Rollout until terminal state}
+3:      U PDATE T ERMINAL($s,v$) {Eq. (15)}
+4:     return $v$
+5:  $a\leftarrow$
+A
+CTION
+P
+ROG
+W
+IDEN$(s)$
+6:   $addSample\leftarrow$ A CTION O PT $(s,a,d)$
+7:  if  $\lvert{C(sa)}\rvert\leq k_{o}N(sa)^{\alpha_{o}}$ OR $addSample$  then
+8:     $s^{\prime},r\leftarrow G(s,a)$$(b^{\prime},b^{-})$ for POMDP}
+9:     $C(sa)\leftarrow C(sa)\cup\left\{(s^{\prime},r)\right\}$
+10:     v
+R
+OLLOUT
+(s,d1)
+[…]
+R
+OLLOUT
+[…]
+[…]
+11:  else
+12:     $s^{\prime},r\leftarrow$ sample uniformly from $C(sa)$
+13:     v
+S
+IMULATE
+(s,d1)
+[…]
+S
+IMULATE
+[…]
+[…]
+14:   U PDATE MIS ($s,a,s^{\prime},r$) {Eqs. (14), (16)}
+procedure
+A
+CTION
+O
+PT
+(s,a,d)
+A
+CTION
+O
+PT
+[…]
+[…]
+1:  $addSample\leftarrow$ FALSE
+2:  for all $k=1,\dots,K_{opt}$ do
+3:     $g_{a}^{q}=\hat{\nabla}_{a}\hat{Q}(s,a)$Eqs. (3) or (4)}
+4:     a
+O
+PT
+(s,a,gaq)
+[…]
+O
+PT
+[…]
+[…]
+{Adam/other}
+5:      $addSample\;|\mathbin{=}$ A CTION U PDATE$(s,a,a^{\prime})$Eqs. (11), (12), (16); Return TRUE if all $\omega^{i}\leq T_{\omega}^{add}$.}
+6:  return $addSample$
+We introduce Action-Gradient MCTS (AGMCTS), a novel online planning algorithm that builds on MCTS with DPW.
+Like PFT-DPW [2], AGMCTS is applicable to POMDPs via ordered particle-belief states.
+A high-level description is provided in Algorithm 1, with novel components highlighted in blue.
+AGMCTS uses the MIS tree for value estimation with gradient-based action optimization, steering value estimates and future sampling accordingly.
+The algorithm supports applying action update either before or after expansion; we present the former variant, which allows gradients to influence whether the current action node should be expanded.
+To make AGMCTS practical, we introduce several approximations and heuristics, which are discussed in detail in Appendix B and briefly outlined here.
+Monte Carlo Gradient Estimation.
+While computing (3) can be done based on all children of the current action node, we borrow from stochastic gradient optimization the idea to use an MC estimate of the gradient.
+When computing exact importance weight updates (11), we sample $K_{O}^{\nabla}$ observation branches with weights $\omega_{a}(s^{\prime})$.
+It is possible to uniformly sample and weight by $\omega_{a}(s^{\prime})$, but we found it to be less effective.
+Alternatively, when using the linear approximation (12), the gradient $\nabla_{a}\log p_{T}(s^{\prime}\mid s,a)$ has to be calculated for each $s^{\prime}$, so we iterate over all the children $C(s,a)$.
+In POMDPs, we estimate (6) with $K_{b}^{\nabla}$ state particles via $\nabla_{a}\log p_{T}(\bar{b}^{-}\mid\bar{b},a)\approx J(K_{b}^{\nabla})^{-1}\sum_{l=1}^{K_{b}^{\nabla}}\nabla_{a}\log p_{T}(s^{-,j_{l}}\mid s^{j_{l}},a),$ where indices $j_{l}$ are sampled uniformly from $[1,\dots,J]$.
+Gradient Optimization.
+We’ve found throughout our experiments that the Adam optimizer [27], employing step size normalization, was crucial for consistent behavior due to the high variance of the gradient estimates.
+We’ve fixed a hyperparameter $K_{opt}$, the number of consecutive gradient step iterations before simulation.
+Together with the step size, this controls a trade-off between accuracy and computational complexity.
+Moreover, to control runtime and add stability to the optimization, we’ve added thresholds on minimum and maximum norm of action updates in certain domains.
+Thresholds For Adding/Deleting State Nodes.
+As a relevance heuristic, we delete child state $s^{\prime}$ from $C(s,a)$ if $\omega_{a}(s^{\prime})<T_{\omega}^{del}$ after action update.
+This ensures that we do not waste computational resources on states with low contribution to the estimator $\hat{Q}(s,a)$.
+Additionally, if $\omega_{a}(s^{\prime})<T_{\omega}^{add}$ for all $s^{\prime}\in C(s,a)$, we force sampling a new posterior node, to ensure a good state sample exists.
+6 Probability Density of Smooth Generative Models via the Area Formula
+Our derivation thus far was based on the assumption that the transition density $p_{T}(s^{\prime}\mid s,a)$ and its log-gradient $\nabla_{a}\log p_{T}(s^{\prime}\mid s,a)$ are computable.
+Yet in most problem settings, the transition density is not explicit, rather it is given in the form of a generative model $s^{\prime}=f(s,a,\nu)$ with $\nu\in\mathbb{R}^{n_{\nu}}$, where $\nu\sim p_{\nu}(\cdot\mid s,a)$ is a noise variable drawn from some known parametric distribution.
+If $n_{\nu}=n_{s}$ and $\nu\mapsto f(s,a,\nu)$ is bijective and differentiable, the change-of-variables formula states
+$\textstyle p(s^{\prime}\mid s,a)=p_{\nu}(\nu^{*}\mid s,a)\left|{D_{\nu}f(s,a,\nu^{*})}\right|^{-1}\quad\text{with}\quad s^{\prime}=f(s,a,\nu^{*}),$ (17)
+where $D_{\nu}f$ is the Jacobian matrix of $f$ w.r.t.
+$\nu$.
+Unfortunately in many settings $n_{\nu}<n_{s}$ and the map $\nu\mapsto f(s,a,\nu)$ is not a bijection but rather an embedding into a lower-dimensional manifold in $\mathbb{R}^{n_{s}}$.
+In this situation, the density $p(s^{\prime}\mid s,a)$ is not defined with respect to the usual Lebesgue measure on $\mathbb{R}^{n_{s}}$, rather with respect to the induced measure on the manifold.
+In this case, the Area Formula allows us to compute the resulting density.
+Negro [28] recently proved in Theorem 4.1 a closed-form expression of the Area Formula for locally Lipschitz maps.
+If $f$ is locally Lipschitz, and it holds that $\operatorname{rank}(D_{\nu}f)=n_{\nu}$ a.e., then the induced probability measure over $s^{\prime}$ is absolutely continuous w.r.t.
+the Hausdorff measure $\mathcal{H}^{n_{\nu}}$ on $\mathbb{R}^{n_{s}}$, and its Radon-Nykodym derivative satisfies
+$\textstyle p(s^{\prime}\mid s,a)=\sum_{\nu^{*}}p_{\nu}(\nu^{*}\mid s,a){(J_{n_{\nu}}f(x,a,\nu^{*}))}^{-1}\quad\forall\nu^{*}:s^{\prime}=f(s,a,\nu^{*}),$ (18)
+or 00 when there are no such $\nu^{*}$, where the $n_{\nu}$-dimensional Jacobian is given by the Cauchy-Binet formula: $J_{n_{\nu}}f(x,a,\nu)=\sqrt{\det((D_{\nu}f)^{t}\cdot(D_{\nu}f))}$.
+This mandates that we’d be able to: (1) Obtain all solutions $\{\nu^{*}:f(s,a,\nu^{*})=s^{\prime}\}$; (2) Compute the Jacobian matrix $D_{\nu}f$.
+Often in robotics and physical domains, $f$ is a physical simulator defined by continuous-time dynamics integrated in time, with either initial conditions that are perturbed by $\nu$ or with noise injected to the output.
+Differentiable physical simulators is a growing research field [29, 30, 31], and under certain conditions both cases can meet requirements (1) and (2).
+Input noise to simulator.
+Let the input noise be described by the function $h(s,a,\nu)=(\tilde{s},\tilde{a})$.
+The simulator then produces a successor state via $s^{\prime}=f(\tilde{s},\tilde{a})$.
+If for a new action $a^{\prime}$ there exists a unique $\nu^{*}$ satisfying $h(s,a^{\prime},\nu^{*})=(\tilde{s},\tilde{a})$, i.e.
+such that the perturbed input remains unchanged, then we can evaluate the sensitivity of the output state with respect to the noise: $D_{\nu}f(s,a^{\prime},\nu^{*})=D_{(\tilde{s},\tilde{a})}f\cdot\left.\frac{\partial h}{\partial\nu}\right|_{s,a^{\prime},\nu^{*}}$.
+Moreover, the Jacobian matrix $D_{(\tilde{s},\tilde{a})}f$ can be cached from the original sample $(s,a,\nu,s^{\prime})$, allowing to compute the density without additional evaluations of the simulator.
+Noise in simulator output.
+Let the simulator output be $x=f(s,a)$, and the next state be given by $s^{\prime}=g(x,\nu)$.
+If we can find all $\nu^{*}$ such that $s^{\prime}=g(f(s,a^{\prime}),\nu^{*})$, then we compute the Jacobian via $D_{\nu}f(s,a^{\prime},\nu^{*})=\frac{\partial g}{\partial\nu}|_{x=f(s,a^{\prime}),\nu^{*}}$ with a single forward simulation $f(s,a^{\prime})$.
+In both cases, once we have a computable function $p(s^{\prime}\mid s,a)$, we can compute the log-gradient $\nabla_{a}\log p(s^{\prime}\mid s,a)$ via the chain rule and automatically differentiating the functions of interest.
+We detail our different computational approaches for the scenarios from Section 7 in Section C.
+7 Experiments
+Numerical simulations were carried out to evaluate the performance of AGMCTS.
+In MDP scenarios we compared it to standard MCTS with DPW, and in POMDPs we compared it to the widely used PFT-DPW and POMCPOW, whose implementations were taken from the POMDPs.jl framework [32].
+AGMCTS was implemented using POMDPs.jl as well.
+Automatic differentiation was mostly done via Enzyme.jl [33], and we used ForwardDiff.jl overloaded with SciMLSensitivity.jl for ODE solution sensitivity [34, 35].
+The analysis and comparison were based on the code of [4].
+The full details of each scenario are in the appendix, and we highlight here important aspects of each.
+ND-Continuous Light-Dark POMDP.
+This is a continuous-form extrapolation of the well-known Light-Dark POMDP.
+It is designed to showcase: (1) Reward function that varies continuously with posterior state and requires very accurate action choice; (2) Initial belief that is non-unimodal (spherical); (3) Conditional observation density that is very non-uniform throughout $\mathcal{S}$.
+Figure 2: 2D-Continuous Light-Dark POMDP.
+The agent (red dot) starts randomly on the red circle, localizes near the beacon (green star) and then aims for the goal (small blue circle).
+The reward function is indicated by the background heatmap.
+The chosen action’s optimization trajectory of AGMCTS is plotted in darkening shades of blue arrows.
+In this domain, $\mathcal{S}=\mathcal{O}=\mathbb{R}^{N}$.
+The agent starts at a random position on the sphere centered at the origin $S^{N-1}_{0.5}$.
+The goal position is located to the top of the origin, and a beacon is located to the right.
+The POMDP terminates if the agent enters the goal zone within a tolerance $T=0.2$ or within $L=6$ steps.
+Actions are limited to $\lVert{a}\rVert\leq 1.5$.
+The transition model is linear with additive Gaussian noise $\mathcal{N}(0,(0.025)^{2}\mathbf{I})$.
+The observations are normally distributed around the agent’s relative position to the beacon, with covariance that increases rapidly with distance.
+The reward function depends only on the posterior state, has high reward that decreases with distance from the goal, a penalty moat around it to encourage mindful actions, and a general global quadratic distance penalty.
+The rollout’s actions head directly to the goal, but for increased difficulty have additive Gaussian noise in each component.
+We benchmark in dimensions $2,3,4$.
+Mountain Car and Hill Car MDP/POMDP.
+These domains are based on the well-known Mountain Car [36] and Car on the Hill (Hill Car for short) [37] problems.
+The goal is to drive a car up a hill by accelerating and braking, while avoiding over-speeding and leaving the left boundary.
+We make both problems significantly harder by terminating an episode early and with a large penalty if the car is over-speeding, and defining a continuous action space $\mathcal{A}=[-a_{\max},a_{\max}]$.
+Mountain Car has the difficulty of very long planning horizons of up to 200 steps, while Hill Car has a medium horizon of up to 30 steps, but its generative transition model is given by integrating the ODE of the car’s dynamics.
+In both problems, additive noise is applied to the action, which is then clamped to $\mathcal{A}$ before being used as input to the simulator, demonstrating a transition distribution that is a mixture of a continuous and discrete components.
+Lunar Lander MDP/POMDP.
+This domain is taken exactly as in [8, 4], where the goal is to land a spacecraft in a target zone safely.
+The vehicle state is $(x,y,\theta,\dot{x},\dot{y},\omega)$, and observations are noisy measurements of the angular rate, horizontal speed and range from the ground.
+The transition model is given by additive Gaussian noise to the output of a deterministic simulator.
+The action space is very large, and failure rewards are sparse, making gradients at times not as informative in this domain.
+7.1 Performance Evaluation
+[…]
+Figure 3: Performance trends of AGMCTS, DPW, PFT-DPW and POMCPOW in the light-dark, mountain/hill car, and lunar lander domains.
+The returns are averaged over 1000 seeds, with shaded area showing ±$\pm 3$ standard errors.
+The x-axis shows the simulations budget per planning session.
+8 Conclusions
+We introduced a novel methodology for computing action gradients in POMDPs, alongside dynamic updates of action branches within MCTS for online planning.
+We evaluated our novel planner AGMCTS across several continuous MDP and POMDP domains, demonstrating both its practicality and effectiveness.
+To support this, we presented a tractable method for computing transition densities and their log-gradients for smooth generative models based on the Area Formula, a technique broadly applicable to continuous MDPs and POMDPs.
+We hope this work paves the way for further research at the intersection of gradient-based optimization and online planning under uncertainty, advancing alongside future developments in both offline and online reinforcement learning.
+Practical Considerations.
+AGMCTS requires more computation time per simulation.
+This depends on many factors, including transition density computation overhead, number of simulations per planning session, tree shape and heuristics.
+In our experiments, each algorithm was run in a single thread, yet particle-belief algorithms such as AGMCTS and PFT-DPW are highly parallelizable.
+It is application-dependent whether the computational cost and modeling requirements of gradient computation are worth the improvement in performance, compared to using more simulations.
+Furthermore, AGMCTS could be combined with more sophisticated heuristics, like Voronoi-based action sampling and Bayesian optimization methods, which might further boost performance.
+Future Work.
+The gradient computation and reuse that we’ve shown for POMDPs are based on the propagated belief, which require full belief trajectories that might be computationally demanding in some domains.
+We plan to extend our methods efficiently for state trajectories as well.
+Finally, AGMCTS requires significant hyperparameter tuning, including DPW parameters and parameters related to action optimization.
+Future work includes algorithms that generalize more effectively.
+Acknowledgments
+This research was partially supported by the Israel Science Foundation (ISF).
+Appendix A Proofs and Derivations
+A.1 Theorem 1
+We first highlight that the required "well-behaved" conditions for the theorem are for changing the order of derivatives and integrals by the Leibniz integral rule.
+Sufficient conditions are characterized by Folland [26, Theorem 2.27]:
+Theorem 2.27.
+Suppose that $f:X\times[a,b]\to\mathbb{C}(-\infty<a<b<\infty)$ and that $f(\cdot,t):X\to\mathbb{C}$ is integrable for each $t\in[a,b]$. Let $F(t)=\int_{X}f(x,t)\mathop{}\!\mathrm{d}\mu(x)$.
+  1. 1.
+    Suppose that there exists $g\in L^{1}(\mu)$ such that $\lvert{f(x,t)}\rvert\leq g(x)$ for all $x,t$.
+If $\lim_{t\to t_{0}}f(x,t)=f(x,t_{0})$ for every x, then $lim_{t\to t_{0}}F(t)=F(t_{0})$; in particular, if $f(x,\cdot)$ is continuous for each x, then F is continuous.
+  2. 2.
+    Suppose that $\partial f/\partial t$ exists and there is a $g\in L^{1}(\mu)$ such that $\lvert{\partial f/\partial t(x,t)}\rvert\leq g(x)$ for all $x,t$.
+Then $F$ is differentiable and $F^{\prime}=\int(\partial f/\partial t)(x,t)\mathop{}\!\mathrm{d}\mu(x)$.
+More general versions of the theorem have been proven [38].
+As this is not the focus of our work, we leave the formalization of Theorem 1 general.
+We note that Lebesgue integrability may hold for a very broad class of functions, including functions with a countable set of discontinuities, and densities of common distributions such as regular and clipped Gaussians.
+This is important for many MDP and POMDP problem formulations, in which sparse rewards and discontinuities are often used.
+Additionally, as stated in the main paper, it must be satisfied that $q(s^{\prime}\mid s,a)=0$ implies $p_{T}(s^{\prime}\mid s,a^{\prime})=0$ in the MDP case, and similarly $q(b^{-}\mid b,a)=0$ implies $p_{T}(b^{-}\mid b,a^{\prime})=0$ for the POMDP case.
+Theorem 1.
+Assume $q=0\Rightarrow p_{T}=0$ and bounded $r$, $V$.
+If $\lVert{\nabla_{a}r}\rVert$, $\lVert{\nabla_{a}p_{T}}\rVert$ have bounding integrable envelopes w.r.t.
+the MDP/POMDP measures, then the following action-gradient identity results from the dominated-convergence form of Leibniz integral rule [26, Theorem 2.27]:
+$\nabla_{a^{\prime}}Q_{t}^{\pi}(s,a^{\prime})=\mathbb{E}_{s^{\prime}\mid q}[\omega_{p_{T}/q}(s^{\prime})[\nabla_{a^{\prime}}\log p_{T}(s^{\prime}\mid s,a^{\prime})(r(s,a^{\prime},s^{\prime})+\gamma V_{t+1}^{\pi}(s^{\prime}))+\nabla_{a^{\prime}}r]],$ (19)
+$\nabla_{a^{\prime}}Q_{t}^{\pi}(b,a^{\prime})=\mathbb{E}_{b^{-},o,b^{\prime}\mid q}[\omega_{p_{T}/q}(b^{-})[\nabla_{a^{\prime}}\log p_{T}(b^{-}\mid b,a^{\prime})(r(b,a^{\prime},b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime}))+\nabla_{a^{\prime}}r]].$ (20)
+Proof.
+We remind the definition of the importance weights
+$\omega_{p_{T}/q}(s^{\prime})=\frac{p_{T}(s{{}^{\prime}}\mid s,a{{}^{\prime}})}{q(s{{}^{\prime}}\mid s,a)},\quad\omega_{p_{T}/q}(b^{-})=\frac{p_{T}(b^{-}\mid s,a{{}^{\prime}})}{q(b^{-}\mid s,a)}.$ (21)
+We start with the MDP case, for which the result is straight-forward:
+$\nabla_{a^{\prime}}Q_{t}^{\pi}(s,a^{\prime})=\nabla_{a^{\prime}}\mathbb{E}_{s^{\prime}\mid q}\left[\omega_{p_{T}/q}\left(s^{\prime}\right)(r\left(s,a^{\prime},s^{\prime}\right)+\gamma V_{t+1}^{\pi}\left(s^{\prime}\right))\right]$ (22)
+$=\nabla_{a^{\prime}}\int_{s^{\prime}}q(s^{\prime}\mid s,a)\frac{p_{T}(s^{\prime}\mid s,a^{\prime})}{q(s^{\prime}\mid s,a)}(r\left(s,a^{\prime},s^{\prime}\right)+\gamma V_{t+1}^{\pi}\left(s^{\prime}\right))\mathop{}\!\mathrm{d}s^{\prime}$ (23)
+$=\nabla_{a^{\prime}}\int_{s^{\prime}}p_{T}(s^{\prime}\mid s,a^{\prime})(r\left(s,a^{\prime},s^{\prime}\right)+\gamma V_{t+1}^{\pi}\left(s^{\prime}\right))\mathop{}\!\mathrm{d}s^{\prime}.$ (24)
+From the assumption that $r$ and $V$ are bounded, their sum is also bounded by $M>0$:
+$\left|{r\left(s,a^{\prime},s^{\prime}\right)+\gamma V_{t+1}^{\pi}\left(s^{\prime}\right)}\right|\leq M$ (25)
+We have that for each $a^{\prime}_{i}$ (a coordinate of $a^{\prime}\in\mathcal{A}$), the partial derivative satisfies
+$\frac{\partial}{\partial a^{\prime}_{i}}\left(p_{T}(s^{\prime}\mid s,a^{\prime})(r\left(s,a^{\prime},s^{\prime}\right)+\gamma V_{t+1}^{\pi}\left(s^{\prime}\right))\right)=$ (26)
+$\frac{\partial}{\partial a^{\prime}_{i}}p_{T}(s^{\prime}\mid s,a^{\prime})(r\left(s,a^{\prime},s^{\prime}\right)+\gamma V_{t+1}^{\pi}\left(s^{\prime}\right))+p_{T}(s^{\prime}\mid s,a^{\prime})(\frac{\partial}{\partial a^{\prime}_{i}}r\left(s,a^{\prime},s^{\prime}\right))=$ (27)
+$p_{T}(s^{\prime}\mid s,a^{\prime})\frac{\partial}{\partial a^{\prime}_{i}}\log p_{T}(s^{\prime}\mid s,a^{\prime})(r\left(s,a^{\prime},s^{\prime}\right)+\gamma V_{t+1}^{\pi}\left(s^{\prime}\right))+p_{T}(s^{\prime}\mid s,a^{\prime})(\frac{\partial}{\partial a^{\prime}_{i}}r\left(s,a^{\prime},s^{\prime}\right)),$ (28)
+where importantly, $\frac{\partial}{\partial a^{\prime}_{i}}V_{t+1}^{\pi}\left(s^{\prime}\right)=0$ due to $s^{\prime}$ being an integration variable that is independent of $a^{\prime}$.
+By our assumptions, it follows that there exist integrable $g_{1}(s^{\prime})$, $g_{2}(s^{\prime})$ w.r.t.
+$p_{T}$ for which $\frac{\partial}{\partial a^{\prime}_{i}}\log p_{T}(s^{\prime}\mid s,a^{\prime})\leq g_{1}(s^{\prime})$ and $r\left(s,a^{\prime},s^{\prime}\right)\leq g_{2}(s^{\prime})$ for all $s$, $a^{\prime}$, $s^{\prime}$.
+Therefore, there exists an integrable envelope $g(s^{\prime})\triangleq Mg_{1}(s^{\prime})+g_{2}(s^{\prime})$ for (26):
+$\left|{p_{T}(s^{\prime}\mid s,a^{\prime})\frac{\partial}{\partial a^{\prime}_{i}}\log p_{T}(s^{\prime}\mid s,a^{\prime})(r\left(s,a^{\prime},s^{\prime}\right)+\gamma V_{t+1}^{\pi}\left(s^{\prime}\right))+p_{T}(s^{\prime}\mid s,a^{\prime})(\frac{\partial}{\partial a^{\prime}_{i}}r\left(s,a^{\prime},s^{\prime}\right))}\right|$ (29)
+$\leq p_{T}(s^{\prime}\mid s,a^{\prime})\left(Mg_{1}(s^{\prime})+g_{2}(s^{\prime})\right)$ (30)
+$=p_{T}(s^{\prime}\mid s,a^{\prime})g(s^{\prime}).$ (31)
+Summation and scalar multiplication preserve integrability, and hence $g(s^{\prime})$ is integrable w.r.t.
+$p_{T}$ [26].
+Therefore, from Theorem 2.27 (clause 2), we can apply the Leibniz integral rule and write in vector form:
+$=\nabla_{a^{\prime}}\int_{s^{\prime}}p_{T}(s^{\prime}\mid s,a^{\prime})(r\left(s,a^{\prime},s^{\prime}\right)+\gamma V_{t+1}^{\pi}\left(s^{\prime}\right))\mathop{}\!\mathrm{d}s^{\prime}=$ (32)
+$\int_{s^{\prime}}\nabla_{a^{\prime}}p_{T}(s^{\prime}\mid s,a^{\prime})(r\left(s,a^{\prime},s^{\prime}\right)+\gamma V_{t+1}^{\pi}\left(s^{\prime}\right))+p_{T}(s^{\prime}\mid s,a^{\prime})(\nabla_{a^{\prime}}r\left(s,a^{\prime},s^{\prime}\right))\mathop{}\!\mathrm{d}s^{\prime},$ (33)
+For the first term, we use the log-gradient substitution:
+$\nabla_{a^{\prime}}p_{T}(s{{}^{\prime}}\mid s,a{{}^{\prime}})=p_{T}(s{{}^{\prime}}\mid s,a{{}^{\prime}})\cdot\nabla_{a^{\prime}}\log p_{T}(s{{}^{\prime}}\mid s,a{{}^{\prime}}),$ (34)
+by gathering terms and bringing back the importance ratio, we obtain the required result (1):
+$\int_{s^{\prime}}\nabla_{a^{\prime}}p_{T}(s^{\prime}\mid s,a^{\prime})(r\left(s,a^{\prime},s^{\prime}\right)+\gamma V_{t+1}^{\pi}\left(s^{\prime}\right))+p_{T}(s^{\prime}\mid s,a^{\prime})(\nabla_{a^{\prime}}r\left(s,a^{\prime},s^{\prime}\right)\mathop{}\!\mathrm{d}s^{\prime}$ (35)
+$=\int_{s^{\prime}}p_{T}(s{{}^{\prime}}\mid s,a{{}^{\prime}})\cdot\nabla_{a^{\prime}}\log p_{T}(s{{}^{\prime}}\mid s,a{{}^{\prime}})(r\left(s,a^{\prime},s^{\prime}\right)+\gamma V_{t+1}^{\pi}\left(s^{\prime}\right))+$
+$p_{T}(s^{\prime}\mid s,a^{\prime})(\nabla_{a^{\prime}}r\left(s,a^{\prime},s^{\prime}\right))\mathop{}\!\mathrm{d}s^{\prime}$ (36)
+$=\int_{s^{\prime}}p_{T}(s{{}^{\prime}}\mid s,a{{}^{\prime}})\left(\nabla_{a^{\prime}}\log p_{T}(s^{\prime}\mid s,a^{\prime})(r\left(s,a^{\prime},s^{\prime}\right)+\gamma V_{t+1}^{\pi}(s^{\prime}))+\nabla_{a^{\prime}}r(s,a^{\prime},s^{\prime})\right)\mathop{}\!\mathrm{d}s^{\prime}$ (37)
+$=\int_{s^{\prime}}q(s^{\prime}\mid s,a)\frac{p_{T}(s{{}^{\prime}}\mid s,a{{}^{\prime}})}{q(s^{\prime}\mid s,a)}$
+$\left(\nabla_{a^{\prime}}\log p_{T}(s^{\prime}\mid s,a^{\prime})(r\left(s,a^{\prime},s^{\prime}\right)+\gamma V_{t+1}^{\pi}(s^{\prime}))+\nabla_{a^{\prime}}r(s,a^{\prime},s^{\prime})\right)\mathop{}\!\mathrm{d}s^{\prime}$ (38)
+$=\mathbb{E}_{s^{\prime}\mid q}\left[\omega_{p_{T}/q}(s^{\prime})[\nabla_{a^{\prime}}\log p_{T}(s^{\prime}\mid s,a^{\prime})(r\left(s,a^{\prime},s^{\prime}\right)+\gamma V_{t+1}^{\pi}(s^{\prime}))+\nabla_{a^{\prime}}r(s,a^{\prime},s^{\prime})]\right].$ (39)
+We now show the POMDP case, which is the main novelty of this theorem. We assume that the same conditions are satisfied for $p_{T},r$ as in the MDP case.
+We begin by showing the propagated belief trick:
+$Q_{t}^{\pi}(b,a)=\mathbb{E}_{b^{\prime}\mid b,a}\left[r(b,a,b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime})\right]$ (40)
+$=\int_{b^{\prime}}p(b^{\prime}\mid b,a)(r(b,a,b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime}))\mathop{}\!\mathrm{d}b^{\prime},$ (41)
+and we continue by marginalizing over the propagated belief $b^{-}$ and the observation $o$:
+$=\int_{b^{-}}\int_{o}\int_{b^{\prime}}p(b^{-},o,b^{\prime}\mid b,a)(r(b,a,b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime}))\mathop{}\!\mathrm{d}b^{\prime}\mathop{}\!\mathrm{d}o\mathop{}\!\mathrm{d}b^{-}.$ (42)
+By its definition, the propagated belief $b^{-}$ is the generated belief by conditioning on the history without the last observation.
+Therefore, by using the chain rule and removing variables that are conditionally independent, we obtain:
+[…]
+$p(b^{-}\mid b,a)(r(b,a,b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime}))\mathop{}\!\mathrm{d}b^{\prime}\mathop{}\!\mathrm{d}o\mathop{}\!\mathrm{d}b^{-}$ (43)
+$=\int_{b^{-}}\int_{o}\int_{b^{\prime}}p(b^{\prime}\mid o,b^{-})p(o\mid b^{-})p(b^{-}\mid b,a)(r(b,a,b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime}))\mathop{}\!\mathrm{d}b^{\prime}\mathop{}\!\mathrm{d}o\mathop{}\!\mathrm{d}b^{-}$ (44)
+$=\mathbb{E}_{b^{-}\mid b,a}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}\left[r(b,a,b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime})\right].$ (45)
+We introduce an importance ratio over the propagated belief to calculate the derivative for proposal action $a$ and target action $a^{\prime}$:
+$Q_{t}^{\pi}(b,a^{\prime})=\mathbb{E}_{b^{-}\mid q}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}\bigg{[}\frac{p_{T}(b^{-}\mid b,a^{\prime})}{q(b^{-}\mid b,a)}(r(b,a^{\prime},b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime}))\bigg{]},$ (46)
+where we denoted $p_{T}(b^{-}\mid b,a^{\prime})\triangleq p(b^{-}\mid b,a^{\prime})$ as the transition model of the propagated belief.
+Similarly to the derivation of the MDP case, we can apply the Leibniz integral rule to obtain the result:
+$\nabla_{a^{\prime}}Q_{t}^{\pi}(b,a^{\prime})=\mathbb{E}_{b^{-}\mid q}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}\bigg{[}\frac{p_{T}(b^{-}\mid b,a^{\prime})}{q(b^{-}\mid b,a)}$
+$(\nabla_{a^{\prime}}\log p_{T}(b^{-}\mid b,a^{\prime})\cdot(r(b,a^{\prime},b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime}))+\nabla_{a^{\prime}}r(b,a^{\prime},b^{\prime}))\bigg{]}.$ (47)
+A.2 Theorem 1 Extended (Immediate Reward Gradient)
+This theorem was not presented in the main paper due to space limitations.
+It is an alternate form of Theorem 1 for MDPs and state-reward POMDPs, and it was used in the gradient calculation of some of the experimental domains, hence it is brought here for completeness.
+Theorem 1 Extended (Immediate Reward Gradient) modifies the immediate reward gradient for MDPs and state-reward POMDPs, by basing it on state samples from the updated action rather than scoring previous value estimates.
 It may give more accurate gradient estimates, albeit possibly more expensive to compute.
+We refer to a state-reward POMDP when the reward function can be expressed as $r(b,a,b^{\prime})=\mathbb{E}_{s,s^{\prime}\mid b,a,b^{\prime}}[r(s,a,s^{\prime})]$.
+We highlight that as an extension of Theorem 1, it requires the same conditions.
+Additionally, for the simplification of the immediate reward expectation, we make use of Fubini’s theorem for changing integration orders.
+Sufficient conditions are that there exists some $R\in\mathbb{R}$ such that $\sup_{s,a,s^{\prime}}\lvert{r(s,a,s^{\prime})}\rvert<R$, and that $r$, $p_{T}$, $p_{O}$ are all a.e.
+continuous and bounded.
+Theorem 1 Extended (Immediate Reward Gradient).
+Under the conditions of Theorem 1, the action gradient of the MDP action-value function is given by:
+$\nabla_{a^{\prime}}Q_{t}^{\pi}(s,a^{\prime})=\mathbb{E}_{s^{\prime}\mid s,a}[\nabla_{a^{\prime}}\log p_{T}(s^{\prime}\mid s,a^{\prime})\cdot r(s,a^{\prime},s^{\prime})+\nabla_{a^{\prime}}r(s,a^{\prime},s^{\prime})]$
+$+\gamma\mathbb{E}_{s^{\prime}\mid q}\left[\frac{p_{T}(s^{\prime}\mid s,a^{\prime})}{q(s^{\prime}\mid s,a)}\nabla_{a^{\prime}}\log p_{T}(s^{\prime}\mid s,a^{\prime})\cdot V_{t+1}^{\pi}(s^{\prime})\right].$ (48)
+For a state-reward POMDP, the action gradient of the action-value function is given by:
+$\nabla_{a^{\prime}}Q_{t}^{\pi}(b,a^{\prime})=\mathbb{E}_{s\mid b}\mathbb{E}_{s^{\prime}\mid s,a^{\prime}}[\nabla_{a^{\prime}}\log p_{T}(s^{\prime}\mid s,a^{\prime})\cdot r(s,a^{\prime},s^{\prime})+\nabla_{a^{\prime}}r(s,a^{\prime},s^{\prime})]$
+$+\gamma\mathbb{E}_{b^{-}\mid q}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}\bigg{[}\frac{p_{T}(b^{-}\mid b,a^{\prime})}{q(b^{-}\mid b,a)}\nabla_{a^{\prime}}\log p_{T}(b^{-}\mid b,a^{\prime})\cdot V_{t+1}^{\pi}(b^{\prime})\bigg{]}.$ (49)
+Proof.
+The central part of this proof is the decomposition of the belief action-value function to an expectation over state trajectories of immediate rewards, and an expectation over belief trajectories of future values.
+Afterwards, we apply the Leibniz integral rule similarly to Theorem 1.
+We show the POMDP case, as the MDP case is rather trivial.
+We start with the belief action-value function:
+$Q_{t}^{\pi}(b,a)=\mathbb{E}_{b^{\prime}\mid b,a}\left[r(b,a,b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime})\right]$ (50)
+$=\mathbb{E}_{b^{\prime}\mid b,a}\left[r(b,a,b^{\prime})\right]+\gamma\mathbb{E}_{b^{\prime}\mid b,a}\left[V_{t+1}^{\pi}(b^{\prime})\right].$ (51)
+We expand the first expectation, using the chain rule and then Fubini’s theorem to cancel the expectation over the observation and posterior belief:
+$\mathbb{E}_{b^{\prime}\mid b,a}\left[r(b,a,b^{\prime})\right]=\mathbb{E}_{b^{-}\mid b,a}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}\left[\mathbb{E}_{s,s^{\prime}\mid b,a,b^{\prime}}[r(s,a,s^{\prime})]\right]$ (52)
+$=\int_{b^{-}}\int_{o}\int_{b^{\prime}}\int_{s}\int_{s^{\prime}}p(s,s^{\prime},o,b^{-},b^{\prime}\mid b,a)r(s,a,s^{\prime})\mathop{}\!\mathrm{d}s^{\prime}\mathop{}\!\mathrm{d}s\mathop{}\!\mathrm{d}b^{\prime}\mathop{}\!\mathrm{d}o\mathop{}\!\mathrm{d}b^{-}$ (53)
+$=\int_{b^{-}}\int_{o}\int_{b^{\prime}}\int_{s}\int_{s^{\prime}}p(o,b^{\prime},b^{-}\mid s,s^{\prime},b,a)p(s,s^{\prime}\mid b,a)r(s,a,s^{\prime})\mathop{}\!\mathrm{d}s^{\prime}\mathop{}\!\mathrm{d}s\mathop{}\!\mathrm{d}b^{\prime}\mathop{}\!\mathrm{d}o\mathop{}\!\mathrm{d}b^{-}$ (54)
+[…]
+[…]
+$=\mathbb{E}_{s\mid b}\mathbb{E}_{s^{\prime}\mid s,a}[r(s,a,s^{\prime})].$ (57)
+Now, we compute $\nabla_{a^{\prime}}$ for the terms obtained.
+The gradient of the immediate reward expectation, when the proposal action is $a$ and the target action is $a^{\prime}$, is given by (skipping some steps similar to Theorem 1):
+$\nabla_{a^{\prime}}\mathbb{E}_{s\mid b}\mathbb{E}_{s^{\prime}\mid s,a}\left[r(s,a^{\prime},s^{\prime})\right]$ (58)
+$=\mathbb{E}_{s\mid b}\mathbb{E}_{s^{\prime}\mid s,a}\left[\left(\nabla_{a^{\prime}}\log p_{T}(s^{\prime}\mid s,a^{\prime})\cdot r(s,a^{\prime},s^{\prime})+\nabla_{a^{\prime}}r(s,a^{\prime},s^{\prime})\right)\right].$ (59)
+Next, the gradient of the future value expectation is simply:
+$\nabla_{a^{\prime}}\mathbb{E}_{b^{\prime}\mid b,a}\left[V_{t+1}^{\pi}(b^{\prime})\right]$ (60)
+$=\nabla_{a^{\prime}}\mathbb{E}_{b^{-}\mid q}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}\left[\frac{p_{T}(b^{-}\mid b,a^{\prime})}{q(b^{-}\mid b,a)}V_{t+1}^{\pi}(b^{\prime})\right]$ (61)
+$=\mathbb{E}_{b^{-}\mid q}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}\left[\frac{p_{T}(b^{-}\mid b,a^{\prime})}{q(b^{-}\mid b,a)}\nabla_{a^{\prime}}\log p_{T}(b^{-}\mid b,a^{\prime})\cdot V_{t+1}^{\pi}(b^{\prime})\right].$ (62)
+Combining the two results, we obtain the final expression (49). ∎
+A.3 Lemma 1
+Lemma 1.
+Let $\bar{b}^{-}=((s^{-,j},\lambda^{j}))_{j=1}^{J}$ be generated from $\bar{b}=((s^{j},\lambda^{j}))_{j=1}^{J}$ by the bootstrap filter.
+Hence, the $j$’th state particle was generated via $s^{-,j}\sim p_{T}(\cdot\mid s^{j},a)$, and it holds that
+$p_{T}(\bar{b}^{-}\mid\bar{b},a)=\prod_{i=1}^{J}p_{T}(s^{-,j}\mid s^{j},a),$ (63)
+$\nabla_{a}\log p_{T}(\bar{b}^{-}\mid\bar{b},a)=\sum_{i=1}^{J}\nabla_{a}\log p_{T}(s^{-,j}\mid s^{j},a).$ (64)
+Proof.
+In the bootstrap filter algorithm, each particle is sampled $s^{-,j}\sim p_{T}(\cdot\mid s^{j},a)$ independently for each $j=1,\ldots,J$. Therefore:
+$p(\bar{b}^{-}\mid\bar{b},a)=p(((s^{-,j},\lambda^{-,j}))_{j=1}^{J}\mid((s^{l},\lambda^{l}))_{l=1}^{J},a),$ (65)
+due to the beliefs being ordered, and the independent sampling, we can factorize as:
+$=\prod_{i=1}^{J}p(((s^{-,j},\lambda^{-,j}))\mid((s^{l},\lambda^{l}))_{l=1}^{J},a)$ (66)
+$=\prod_{i=1}^{J}p(((s^{-,j},\lambda^{-,j}))\mid(s^{j},\lambda^{j}),a).$ (67)
+Finally, since $\lambda^{-,j}=\lambda^{j}$, we can eliminate the weights to obtain
+$p(\bar{b}^{-}\mid\bar{b},a)=\prod_{i=1}^{J}p_{T}(s^{-,j}\mid s^{j},a),$ (68)
+and we denote $p_{T}(\bar{b}^{-}\mid\bar{b},a)\triangleq p(\bar{b}^{-}\mid\bar{b},a)$. ∎
+A.4 Theorem 2
+We first start giving the formal statement of Lemma 9.1 by Veach and Guibas [15]:
+Lemma 9.1.
+Let $F$ be any estimator of the form
+$F=\sum_{i=1}^{n}\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}w_{i}(X_{i,j})\frac{f(X_{i,j})}{p_{i}(X_{i,j})},$
+where $n_{i}\geq 1$ for all $i$, and the weighting functions $w_{i}$ satisfy conditions
+[…]
+[…]
+The proof is simple and is satisfied by the linearity of the expectation.
+An important corollary brought by Veach and Guibas [15] is that these conditions imply that at any point where $f(x)\neq 0$, at least one of the $p_{i}(x)$ must be positive.
+Thus, it is not necessary for every $p_{i}$ to sample the whole domain, and it is allowable for some $p_{i}$ to be specialized sampling techniques that concentrate on specific regions of the integrand.
+We remind that our target function to estimate is given by the equations:
+$Q_{t}^{\pi}(s,a^{\prime})=\mathbb{E}_{s^{\prime}\mid q}\left[\omega_{p_{T}/q}(s^{\prime})(r(s,a^{\prime},s^{\prime})+\gamma V_{t+1}^{\pi}(s^{\prime}))\right]$ (69)
+$=\mathbb{E}_{s^{\prime}\mid q}\left[\frac{p_{T}(s{{}^{\prime}}\mid s,a{{}^{\prime}})}{q(s{{}^{\prime}}\mid s,a)}(r(s,a^{\prime},s^{\prime})+\gamma V_{t+1}^{\pi}(s^{\prime}))\right].$ (70)
+Hence, for each child observation branch in $C(s,a)$, the target function to estimate is the same $p_{T}(s{{}^{\prime}}\mid s,a{{}^{\prime}})(r(s,a^{\prime},s^{\prime})+\gamma V_{t+1}^{\pi}(s^{\prime}))$, with a sampling technique that is the proposal transition distribution based on the action that was used to generate this state child, i.e.
+$q(s^{\prime}\mid s,a)=p_{T}(s^{\prime}\mid s,a_{prop}(s^{\prime}))$.
+We define the following MIS estimators:
+$\hat{V}_{f}(s,a)\triangleq\sum_{s^{\prime}\in C(s,a)}w_{i}\frac{p_{T}(s^{\prime}\mid s,a)}{p_{T}(s^{\prime}\mid s,a_{prop}(s^{\prime}))}\hat{V}(s^{\prime}),$ (71)
+$\hat{r}(s,a)\triangleq\sum_{s^{\prime}\in C(s,a)}w_{i}\frac{p_{T}(s^{\prime}\mid s,a)}{p_{T}(s^{\prime}\mid s,a_{prop}(s^{\prime}))}r(s,a,s^{\prime}),$ (72)
+for weights $w_{i}$ that may in general depend on the $s$, $a$, $s^{\prime}$ and $a^{\prime}$. Moreover, we defined the value estimate of each state node recursively as:
+$\hat{V}(s)\triangleq\sum_{a\in C(s)}\frac{n(s,a)}{n(s)}\hat{Q}(s,a),$ (73)
+unless $s^{\prime}$ is at the maximum tree depth, in which case its value estimate is the mean of the rollout values:
+$\hat{V}(s)=\frac{1}{n(s)_{+1}}\sum_{i=1}^{n(s)_{+1}}v^{i},$ (74)
+for all $v^{i}$ rollout values from the terminal node $s$.
 Theorem 2.
-For a state-reward POMDP, the action gradient of the action-value function is given by:
-$\textstyle\nabla_{a^{\prime}}Q_{t}^{\pi}(b,a)=\mathbb{E}_{s\mid b}\mathbb{E}_{s^{\prime}\mid s,a^{\prime}}[\nabla_{a^{\prime}}\log p_{T}(s^{\prime}\mid s,a^{\prime})\cdot r(s,a^{\prime},s^{\prime})$
-$\textstyle+\nabla_{a^{\prime}}r(s,a^{\prime},s^{\prime})]\textstyle+\gamma\mathbb{E}_{b^{-}\mid b,a}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}[\frac{p_{T}(b^{-}\mid b,a^{\prime})}{p_{T}(b^{-}\mid b,a)}$
-$\textstyle\nabla_{a^{\prime}}\log p_{T}(b^{-}\mid b,a^{\prime})\cdot V_{t+1}^{\pi}(b^{\prime})].$ (15)
-For ordered particle beliefs, the gradient log of the propagated transition likelihood from Lemma 1 is given by:
-$\textstyle\nabla_{a}\log p_{T}(\bar{b}^{-}\mid\bar{b},a)=\sum_{i=1}^{J}\nabla_{a}\log p_{T}(s^{-,j}\mid s^{j},a).$ (16)
-3.5 Action Gradient Monte Carlo Tree Search
-Algorithm 1 AGMCTS
+If the weighting strategy $w_{i}$ satisfies assumption (MIS-I) and (MIS-II), then the MIS tree yields unbiased estimates of the value and action-value functions for a given tree structure.
+Proof.
+The proof goes by a mutual induction over $\hat{V}$ and $\hat{Q}$, proving that a node’s value or action-value estimate is unbiased if all its children have unbiased estimates.
+Base case: leaf state nodes are unbiased.
+As stated in the tree construction, $\hat{V}(s)$ when $s$ is a leaf node is the mean of the rollout values, therefore:
+$\mathbb{E}_{s\sim\pi_{rollout}}\left[\hat{V}(s)\right]=\mathbb{E}_{s\sim\pi_{rollout}}\left[\frac{1}{n(s)_{+1}}\sum_{i=1}^{n(s)_{+1}}v^{i}\right]$ (76)
+$=\frac{1}{n(s)_{+1}}\sum_{i=1}^{n(s)_{+1}}\mathbb{E}_{s\sim\pi_{rollout}}\left[v^{i}\right]$ (77)
+$=V^{\pi_{rollout}}(s),$ (78)
+hence it is unbiased.
+Inductive step 1: Value estimates $\hat{V}$ are unbiased.
+Induction hypothesis: Assume that $\hat{Q}(s,a)$ is unbiased for all $(s,a)\in C(s)$.
+The value estimate is defined as the mean of the action-value estimates weighted by the number of times each action was selected.
+Therefore, for the policy representing the current visitation counters, i.e.
+defined by $\pi_{visits}(s)\stackrel{{\scriptstyle\triangle}}{{\sim}}n(s,a)/n(s)$, the value estimate satisfies:
+$\mathbb{E}\left[\hat{V}(s)\right]=\mathbb{E}\left[\sum_{a\in C(s)}\frac{n(s,a)}{n(s)}\hat{Q}(s,a)\right]$ (79)
+$=\sum_{a\in C(s)}\frac{n(s,a)}{n(s)}\mathbb{E}\left[\hat{Q}(s,a)\right]$ (80)
+$=\sum_{a\in C(s)}p(a\mid\pi_{visits}(s))Q(s,a)$ (81)
+$=\mathbb{E}_{a\sim\pi_{visits}(s)}\left[V(s)\right]$ (82)
+$=V^{\pi_{visits}}(s),$ (83)
+where we used the definition of $\pi_{visits}$ and the induction hypothesis $\mathbb{E}\left[\hat{Q}(s,a)\right]=Q(s,a)$.
+This explains the importance of "given tree structure" in the theorem statement - when changing the visitation counters and the according action selection policy at each state node, it results in a different weighting of action-value estimates, and hence could introduce a bias when considering the action-value function of a different policy at the root node of the search tree.
+Inductive step 2: Action-value estimates $\hat{Q}$ are unbiased.
+Induction hypothesis: Assume that $\hat{V}(s^{\prime})$ is unbiased for all $s^{\prime}\in C(s,a)$.
+We note that from our construction, $s^{\prime}$ was sampled from $p_{T}(s^{\prime}\mid s,a_{prop}(s^{\prime}))$, it holds that $p_{T}(s^{\prime}\mid s,a_{prop}(s^{\prime}))>0$ for all $s^{\prime}\in C(s,a)$.
+Therefore, from assumption (MIS-II), we have that $w_{i}\neq 0$ for all $i$.
+By using the tower property and the induction hypothesis, the expected future-value estimate is:
+$\mathbb{E}\left[\hat{V}_{f}(s,a)\right]=\mathbb{E}\left[\sum_{s^{\prime}\in C(s,a)}w_{i}\frac{p_{T}(s^{\prime}\mid s,a)}{p_{T}(s^{\prime}\mid s,a_{prop}(s^{\prime}))}\hat{V}(s^{\prime})\right]$ (84)
+=$\stackrel{{\scriptstyle(1)}}{{=}}\mathbb{E}\left[\mathbb{E}\left[\sum_{s^{\prime}\in C(s,a)}w_{i}\frac{p_{T}(s^{\prime}\mid s,a)}{p_{T}(s^{\prime}\mid s,a_{prop}(s^{\prime}))}\hat{V}(s^{\prime})\;\bigg{|}\;s^{\prime}\right]\right]$ (85)
+=$\stackrel{{\scriptstyle(2)}}{{=}}\mathbb{E}\left[\sum_{s^{\prime}\in C(s,a)}w_{i}\frac{p_{T}(s^{\prime}\mid s,a)}{p_{T}(s^{\prime}\mid s,a_{prop}(s^{\prime}))}\mathbb{E}\left[\hat{V}(s^{\prime})\;\bigg{|}\;s^{\prime}\right]\right]$ (86)
+=$\stackrel{{\scriptstyle(3)}}{{=}}\mathbb{E}\left[\sum_{s^{\prime}\in C(s,a)}w_{i}\frac{p_{T}(s^{\prime}\mid s,a)}{p_{T}(s^{\prime}\mid s,a_{prop}(s^{\prime}))}V(s^{\prime})\right]$ (87)
+[…]
+$=\int_{s^{\prime}}p_{T}(s^{\prime}\mid s,a)V(s^{\prime})\mathop{}\!\mathrm{d}s^{\prime}$ (89)
+$=\mathbb{E}\left[V(s^{\prime})\right]$ (90)
+where transitions (1) and (2) are due to the tower property of expectations, (3) is due to the induction hypothesis, and (4) follows from the linearity of the expectation/integral.
+Similarly, we can show that:
+$\mathbb{E}\left[\hat{r}(s,a)\right]=\mathbb{E}\left[\sum_{s^{\prime}\in C(s,a)}w_{i}\frac{p_{T}(s^{\prime}\mid s,a)}{p_{T}(s^{\prime}\mid s,a_{prop}(s^{\prime}))}r(s,a,s^{\prime})\right]$ (91)
+$=\mathbb{E}\left[r(s,a,s^{\prime})\right].$ (92)
+We can now combine the two results to obtain:
+$\mathbb{E}\left[\hat{Q}(s,a)\right]=\mathbb{E}\left[\hat{r}(s,a)+\gamma\hat{V}_{f}(s,a)\right]$ (93)
+$=\mathbb{E}\left[r(s,a,s^{\prime})+\gamma V(s^{\prime})\right]$ (94)
+$=Q(s,a).$ (95)
+In POMDPs, when using particle-belief approximations instead of exact belief representation, this approximation may incur a bias in the action-value estimate, that can be bounded with finite-sample concentration inequalities.
+For an in-depth analysis of this bias we refer the reader to [39].
+∎
+A.5 Derivations of MIS Tree Update Equations
+In this subsection we show the derivation of the MIS tree update equation, and also show their numerically stable forms using log likelihoods for probability values.
+We denote the regular and the weighted log-sum-exp function as $LSE$, differentiated by having one or two vector inputs, and they are defined as
+$LSE(\boldsymbol{x}_{1,\ldots,n})\triangleq\log\left(\sum_{i=1}^{n}\exp(\boldsymbol{x}_{i})\right),$ (96)
+$LSE(\boldsymbol{x}_{1,\ldots,n},\boldsymbol{a}_{1,\ldots,n})\triangleq\log\left(\sum_{i=1}^{n}\exp(\boldsymbol{x}_{i})\cdotp\boldsymbol{a}_{i}\right).$ (97)
+To handle negative $\boldsymbol{a}_{1,\ldots,n}$, (97) can be defined to compute the log of absolute value and sign separately of the log argument, such that the sign is multiplied back after exponentiation if required.
+We also denote a vector with brackets []delimited-[][$\cdot$][ ⋅ ].
+We slightly modify the notations to an enumerated form in this subsection to later match vector notations. In these notations, the visitation counter relationships are:
+$n(s)=\sum_{C(s)}n(s,a^{i}),\;n(s,a)=\sum_{C(s,a)}n(s^{\prime,i})_{+1},$ (98)
+the state value estimator:
+$\hat{V}(s)\triangleq\sum_{C(s)}\frac{n(s,a^{i})}{n(s)}\hat{Q}(s,a^{i}),$ (99)
+and the self normalized MIS estimator:
+$\eta(s,a)\triangleq\sum_{C(s,a)}\omega_{a}(s^{\prime,i})n(s^{\prime,i})_{+1},$ (100)
+$\hat{V}_{f}(s,a)=\eta(s,a)^{-1}\sum_{C(s,a)}\omega_{a}(s^{\prime,i})n(s^{\prime,i})_{+1}\hat{V}(s^{\prime,i}).$ (101)
+Additionally we denote $M\triangleq\lvert{C(s,a)}\rvert$.
+We define the following vectors for each a given action node $(s,a)$ in the MDP case, or $(\bar{b},a)$ in the particle-belief MDP (PB-MDP) case:
+  1. 1.
+    Child state visitation counters.
+    MDP:
+    $\boldsymbol{n}\triangleq(n(s^{\prime,i})_{+1})_{i=1}^{M},$ (102)
+    PB-MDP:
+    $\boldsymbol{n}\triangleq(n(\bar{b}^{\prime,i})_{+1})_{i=1}^{M}.$ (103)
+  2. 2.
+    Target transition likelihoods.
+    MDP:
+    $\boldsymbol{p}=(p_{T}(s^{\prime,i}\mid s,a))_{i=1}^{M},$ (104)
+    PB-MDP:
+    $\boldsymbol{p}=(p_{T}(\bar{b}^{-,i}\mid\bar{b},a))_{i=1}^{M}.$ (105)
+  3. 3.
+    Proposal transition likelihoods.
+    MDP:
+    $\boldsymbol{q}=(p_{T}(s^{\prime,i}\mid s,a_{prop}(s^{\prime,i})))_{i=1}^{M},$ (106)
+    PB-MDP:
+    $\boldsymbol{q}=(p_{T}(\bar{b}^{-,i}\mid\bar{b},a_{prop}(\bar{b}^{-,i})))_{i=1}^{M}.$ (107)
+  4. 4.
+    Immediate rewards.
+    MDP:
+    $\boldsymbol{r}=(r(s,a,s^{\prime,i}))_{i=1}^{M}$ (108)
+    PB-MDP:
+    $\boldsymbol{r}=(r(\bar{b},a,\bar{b}^{\prime,i}))_{i=1}^{M}.$ (109)
+  5. 5.
+    Future value estimates.
+    MDP:
+    $\boldsymbol{v}=(\hat{V}(s^{\prime,i}))_{i=1}^{M}$ (110)
+    PB-MDP:
+    $\boldsymbol{v}=(\hat{V}(\bar{b}^{\prime,i}))_{i=1}^{M}$ (111)
+From here on, the PB-MDP case is the same up to different notations as the MDP case, and we will show only the latter.
+Using these notations the immediate reward and future value estimates can be compactly written:
+$\eta(s,a)=\boldsymbol{1}^{T}(\boldsymbol{n}\odot\boldsymbol{p}\oslash\boldsymbol{q}),$ (112)
+$\hat{r}(s,a)=\eta(s,a)^{-1}\boldsymbol{1}^{T}(\boldsymbol{n}\odot\boldsymbol{p}\odot\boldsymbol{r}\oslash\boldsymbol{q}),$ (113)
+$\hat{V}_{f}(s,a)=\eta(s,a)^{-1}\boldsymbol{1}^{T}(\boldsymbol{n}\odot\boldsymbol{p}\odot\boldsymbol{v}\oslash\boldsymbol{q}),$ (114)
+[…]
+We store the following values:
+  1. 1.
+    $l(\eta)$.
+  2. 2.
+    Log likelihoods $l(\boldsymbol{p})$ and $l(\boldsymbol{q})$.
+  3. 3.
+    $\boldsymbol{n}^{\prime}$ and previous visitation counters $\boldsymbol{n}$.
+  4. 4.
+    $\boldsymbol{v}^{\prime}$ and previous future value estimates $\boldsymbol{v}$.
+The equations for $\hat{r}(s,a)$ are analogous to $\hat{V}_{f}(s,a)$, and we will not show their derivation explicitly.
+Action Backpropagation.
+Let $s^{\prime,j}$ be an updated node.
+Hence, we updated $\boldsymbol{n}_{j}$ to $\boldsymbol{n}^{\prime}_{j}$ and $\boldsymbol{v}_{j}$ to $\boldsymbol{v}^{\prime}_{j}$.
+We update $n(s,a)$ by (98) and perform
+$\eta^{\prime}(s,a)=\eta(s,a)+\frac{\boldsymbol{p}_{j}}{\boldsymbol{q}_{j}}\left(\boldsymbol{n}_{j}^{\prime}-\boldsymbol{n}_{j}\right),$ (115)
+$\hat{V}_{f}^{\prime}(s,a)=\frac{1}{\eta^{\prime}(s,a)}\left(\eta(s,a)\hat{V}_{f}(s,a)+\frac{\boldsymbol{p}_{j}}{\boldsymbol{q}_{j}}(\boldsymbol{n}_{j}^{\prime}\boldsymbol{v}_{j}^{\prime}-\boldsymbol{n}_{j}\boldsymbol{v}_{j})\right).$ (116)
+Therefore, the log-form update equations are:
+$l(\eta^{\prime}(s,a))=LSE\left([l(\eta(s,a)),l(\boldsymbol{p}_{j})-l(\boldsymbol{q}_{j})+l(\boldsymbol{n}_{j}^{\prime}-\boldsymbol{n}_{j})]\right)$ (117)
+$\hat{V}_{f}^{\prime}(s,a)=\exp\Big{(}LSE\big{(}[l(\eta(s,a)),l(\boldsymbol{p}_{j})-l(\boldsymbol{q}_{j})],[\hat{V}_{f}(s,a),\boldsymbol{n}_{j}^{\prime}\boldsymbol{v}_{j}^{\prime}-\boldsymbol{n}_{j}\boldsymbol{v}_{j}]\big{)}-l(\eta^{\prime}(s,a))\Big{)}.$ (118)
+Proof.
+We start with the update for $\eta^{\prime}(s,a)$:
+$\eta^{\prime}(s,a)-\eta(s,a)$ (119)
+$=\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}^{\prime}\boldsymbol{p}_{i}}{\boldsymbol{q}_{i}}-\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}}{\boldsymbol{q}_{i}}$ (120)
+$=\frac{\boldsymbol{p}_{j}}{\boldsymbol{q}_{j}}(\boldsymbol{n}_{j}^{\prime}-\boldsymbol{n}_{j}),$ (121)
+due to all terms being equal except for $\boldsymbol{n}_{j}^{\prime}\neq\boldsymbol{n}_{j}$. Similarly, for $\hat{V}_{f}(s,a)$:
+$\hat{V}_{f}^{\prime}(s,a)-\frac{\eta(s,a)}{\eta^{\prime}(s,a)}\hat{V}_{f}(s,a)$ (122)
+$=\frac{1}{\eta^{\prime}(s,a)}\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}^{\prime}\boldsymbol{p}_{i}\boldsymbol{v}_{i}^{\prime}}{\boldsymbol{q}_{i}}-\frac{\eta(s,a)}{\eta^{\prime}(s,a)}\frac{1}{\eta(s,a)}\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}\boldsymbol{v}_{i}}{\boldsymbol{q}_{i}}$ (123)
+$=\frac{1}{\eta^{\prime}(s,a)}\left(\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}^{\prime}\boldsymbol{p}_{i}\boldsymbol{v}_{i}^{\prime}}{\boldsymbol{q}_{i}}-\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}\boldsymbol{v}_{i}}{\boldsymbol{q}_{i}}\right)$ (124)
+$=\frac{1}{\eta^{\prime}(s,a)}\frac{\boldsymbol{p}_{j}}{\boldsymbol{q}_{j}}(\boldsymbol{n}_{j}^{\prime}\boldsymbol{v}_{j}^{\prime}-\boldsymbol{n}_{j}\boldsymbol{v}_{j}),$ (125)
+due to all terms being equal except for $\boldsymbol{v}^{\prime}_{j}\neq\boldsymbol{v}_{j}$, $\boldsymbol{n}^{\prime}_{j}\neq\boldsymbol{n}_{j}$. ∎
+State Expansion.
+This case is the same as action backpropagation, where the updated index is a new summand at index $j=M+1$. The same equations apply with $n(s^{\prime,j})=0$.
+Proof.
+We append new terms to all vectors: $\boldsymbol{n}_{M+1}$, $\boldsymbol{p}_{M+1}$, $\boldsymbol{v}_{M+1}$, $\boldsymbol{q}_{M+1}$. Update to $\eta(s,a)$:
+$\eta^{\prime}(s,a)-\eta(s,a)$ (126)
+$=\sum_{i=1}^{M+1}\frac{\boldsymbol{n}_{i}^{\prime}\boldsymbol{p}^{\prime}_{i}}{\boldsymbol{q}^{\prime}_{i}}-\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}}{\boldsymbol{q}_{i}}$ (127)
+$=\frac{\boldsymbol{n}_{M+1}^{\prime}\boldsymbol{p}^{\prime}_{M+1}}{\boldsymbol{q}^{\prime}_{M+1}},$ (128)
+due to all terms being equal except for index $M+1$.
+This corresponds to (121) with $\boldsymbol{n}_{M+1}=0$, after initializing $\boldsymbol{p}^{\prime}_{M+1}=\boldsymbol{p}_{M+1}$ and $\boldsymbol{q}^{\prime}_{M+1}=\boldsymbol{q}_{M+1}$ Similarly, for $\hat{V}_{f}(s,a)$:
+$\hat{V}_{f}^{\prime}(s,a)-\frac{\eta(s,a)}{\eta^{\prime}(s,a)}\hat{V}_{f}(s,a)$ (129)
+$=\frac{1}{\eta^{\prime}(s,a)}\sum_{i=1}^{M+1}\frac{\boldsymbol{n}_{i}^{\prime}\boldsymbol{p}_{i}\boldsymbol{v}_{i}^{\prime}}{\boldsymbol{q}_{i}}-\frac{\eta(s,a)}{\eta^{\prime}(s,a)}\frac{1}{\eta(s,a)}\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}\boldsymbol{v}_{i}}{\boldsymbol{q}_{i}}$ (130)
+$=\frac{1}{\eta^{\prime}(s,a)}\left(\sum_{i=1}^{M+1}\frac{\boldsymbol{n}_{i}^{\prime}\boldsymbol{p}_{i}\boldsymbol{v}_{i}^{\prime}}{\boldsymbol{q}_{i}}-\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}\boldsymbol{v}_{i}}{\boldsymbol{q}_{i}}\right)$ (131)
+$=\frac{1}{\eta^{\prime}(s,a)}\frac{\boldsymbol{p}^{\prime}_{M+1}}{\boldsymbol{q}^{\prime}_{M+1}}\boldsymbol{n}_{M+1}^{\prime}\boldsymbol{v}_{M+1}^{\prime},$ (132)
+which again corresponds to (125) with $\boldsymbol{n}_{M+1}=0$. ∎
+Action Update.
+Let $a^{\prime}$ be the new action. Here, we have to calculate $\omega_{a^{\prime}}(s^{\prime,i})$ for all $s^{\prime,i}\in C(s,a)$ in $O(M)$ time. We update
+$\omega_{a^{\prime}}(s^{\prime,i})=p_{T}(s^{\prime,i}\mid s,a^{\prime})(p_{T}(s^{\prime,i}\mid s,a(s^{\prime,i})))^{-1},$ (133)
+and recalculate (100) and (101) with the new $\omega_{a^{\prime}}(s^{\prime,i})$, which are also $O(M)$ time operations.
+For the immediate reward, new rewards must be computed for the new triplets $(s,a^{\prime},s^{\prime,i})$, and the update is analogous.
+Therefore, the log-form equations are:
+$l(\eta^{\prime}(s,a))=LSE(l(\boldsymbol{n})\oplus l(\boldsymbol{p}^{\prime})\ominus l(\boldsymbol{q}))$ (134)
+$\hat{V}_{f}^{\prime}(s,a)=\exp\left(LSE(l(\boldsymbol{p}^{\prime})\ominus l(\boldsymbol{q}),\boldsymbol{n}\odot\boldsymbol{v})-l(\eta^{\prime}(s,a))\right)$ (135)
+Proof.
+This update corresponds to computing a new target likelihood vector $\boldsymbol{p}$. Therefore, the update equations follow from the definitions:
+$\eta^{\prime}(s,a)=\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}^{\prime}}{\boldsymbol{q}_{i}},$ (136)
+$\hat{V}_{f}^{\prime}(s,a^{\prime})=\frac{1}{\eta^{\prime}(s,a)}\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}^{\prime}\boldsymbol{v}_{i}}{\boldsymbol{q}_{i}}.$ (137)
+Terminal State Backpropagation.
+Let $s$ be a terminal node. Its value estimate is based only on rollouts. Hence, for a new rollout value $v^{\prime}$, we perform a running average update
+$\hat{V}^{\prime}(s)=\hat{V}(s)+\frac{n^{\prime}(s)-n(s)}{n^{\prime}(s)_{+1}}(v^{\prime}-\hat{V}(s)),$ (138)
+where $n^{\prime}(s)$ is $n(s)$ plus the number of rollouts (usually 1).
+Proof.
+This is a simple running average update for the mean estimator:
+$\hat{V}(s)=\frac{1}{n(s)_{+1}}\sum_{i=1}^{n(s)_{+1}}v^{i},$ (139)
+for all $v^{i}$ rollout values from the terminal node $s$. ∎
+Non-Terminal State Backpropagation.
+Let $(s,a^{j})$ be an updated node. Hence, we updated $n(s,a^{j})$ to $n^{\prime}(s,a^{j})$ and $\hat{Q}(s,a^{j})$ to $\hat{Q}^{\prime}(s,a^{j})$. We update $n^{\prime}(s)$ by (98) and perform
+$\hat{V}^{\prime}(s)=(n^{\prime}(s))^{-1}\left(n(s)\hat{V}(s)+n^{\prime}(s,a^{j})\hat{Q}^{\prime}(s,a^{j})-n(s,a^{j})\hat{Q}(s,a^{j})\right).$ (141)
+Proof.
+Following the definition of the state value estimator (99), the updated estimator is:
+$\hat{V}^{\prime}(s)=\sum_{C(s)}\frac{n^{\prime}(s,a^{i})}{n^{\prime}(s)}\hat{Q}^{\prime}(s,a^{i}).$ (142)
+Therefore:
+$\hat{V}^{\prime}(s)-\frac{n(s)}{n^{\prime}(s)}\hat{V}(s)$ (143)
+$=\sum_{C(s)}\frac{n^{\prime}(s,a^{i})}{n^{\prime}(s)}\hat{Q}^{\prime}(s,a^{i})-\frac{n(s)}{n^{\prime}(s)}\sum_{C(s)}\frac{n(s,a^{i})}{n(s)}\hat{Q}(s,a^{i})$ (144)
+$=\frac{1}{n^{\prime}(s)}\sum_{C(s)}(n^{\prime}(s,a^{i})\hat{Q}^{\prime}(s,a^{i})-n(s,a^{i})\hat{Q}(s,a^{i})).$ (145)
+All terms are equal except for the updated node $j$. Hence, we obtain
+$\hat{V}^{\prime}(s)-\frac{n(s)}{n^{\prime}(s)}\hat{V}(s)$
+$=\frac{1}{n^{\prime}(s)}\left(n^{\prime}(s,a^{j})\hat{Q}^{\prime}(s,a^{j})-n(s,a^{j})\hat{Q}(s,a^{j})\right).$ (146)
+Appendix B AGMCTS Algorithm
+Algorithm 2 AGMCTS - Full Pseudocode
 procedure S IMULATE$(s,d)$
@@
 1:  $addSample\leftarrow$ FALSE
-2:  for all $k=1,\dots,K_{opt}$ do
-3:     $g_{a}^{q}=\hat{\nabla}_{a}\hat{Q}(s,a)$Eq. (19) or (17)}
-4:     Initialize $a^{acc}\leftarrow a$ for $(s,a)$ if required
-5:     aacc
+2:  if
+A
+CTION
+O
+PT
+R
+ULE(s,a,d) then
+3:     for all $k=1,\dots,K_{opt}$ do
+4:        $g_{a}^{q}=\hat{\nabla}_{a}\hat{Q}(s,a)$Eq. (149) or (148)}
+5:        Initialize $a^{acc}\leftarrow a$ for $(s,a)$ if required
+6:        aacc
 O
@@
 {Adam/other optimizer}
-6:     if $\lVert{a^{acc}-a}\rVert>T_{d_{a}}$ then
-7:        $addSample\;|\mathbin{=}\;$
+7:        if $\lVert{a^{acc}-a}\rVert>T_{d_{a}}^{\min}$ then
+8:           $a^{\prime}\;\mathbin{=}\;$
+C
+LIP
+N
+ORM$(a^{acc},T_{d_{a}}^{\max})$if $\lVert a^{\mathrm{acc}}\rVert>T_{d_{a}}^{\max}$, scale it down; else keep it}
+9:           $addSample\;|\mathbin{=}$
 A
@@
 U
-PDATE$(s,a,a^{acc})$Eqs. (11), (13); Return TRUE if all $\omega^{i}\leq T_{\omega}^{add}$.}
-8:  return $addSample$
-We introduce Action Gradient MCTS (AGMCTS), a novel online planning algorithm based on MCTS with DPW.
-It is applicable to POMDPs via ordered particle belief states, like PFT-DPW [?].
-It is described in high-level in Algorithm 1, with novel components highlighted in blue.
-It combines the MIS estimators for the action-value function, with gradient-based action optimization, and updates actions accordingly.
-AGMCTS applies action update before expansion.
-This allows us to pass the information from to determine whether to expand the current action node.
-Monte Carlo Gradient Estimation.
-While computing (14), (15) can be done based on all children of the current action node, we employ the idea of stochastic gradient optimization to MC estimate the gradient.
-The first estimate is to sample $K_{O}^{\nabla}$ observation branches with weights $\omega^{i}$.
-It is possible to uniformly sample and weight by $\omega^{i}$, but we found it to be less effective.
+PDATE$(s,a,a^{\prime})$Eqs. (11), (12), (16); Return TRUE if all $\omega_{a^{\prime}}(s^{\prime})\leq T_{\omega}^{add}$.}
+10:  return $addSample$
+The full AGMCTS algorithm description is outlined in Algorithm 2.
+It is based on MCTS with DPW, with the new components highlighted in blue.
+We discuss several of the heuristics and optimizations that we found to be crucial for the performance of AGMCTS.
+B.1 Monte Carlo Gradient Estimation
+While computing (19), (49) can be done based on all children of the current action node, we borrow from stochastic gradient optimization the idea to use an MC estimate the gradient.
+The first MC estimate in AGMCTS is done for POMDPs in the computation of the transition probability log-gradient. We estimate (6) with $K_{b}^{\nabla}$ state particles via
+^$\hat{\nabla}_{a}\log p_{T}(\bar{b}^{-}\mid\bar{b},a)=\frac{J}{K_{b}^{\nabla}}\sum_{l=1}^{K_{b}^{\nabla}}\nabla_{a}\log p_{T}(s^{-,j_{l}}\mid s^{j_{l}},a),$ (147)
+where indices $j_{l}$ are sampled uniformly from $[1,\dots,J]$. Due to (64), this estimator is unbiased.
+In the case we perform exact importance weight update (133), we also sample a number of observation branches to compute the gradient.
+It is possible to uniformly sample and weight by $\omega_{a}(s^{\prime})$, but we found it to be less effective.
 The MC approximation of (19) becomes
-^$\textstyle\hat{\nabla}_{a}\hat{Q}(s,a)=\frac{1}{K_{O}^{\nabla}}\sum_{k=1}^{K_{O}^{\nabla}}\nabla_{a}\log p_{T}(s^{\prime,i_{k}}\mid s,a)$
-$\textstyle\cdot(r(s,a,s^{\prime,i_{k}})+\gamma\hat{V}(s^{\prime,i_{k}}))+\nabla_{a}r(s,a,s^{\prime,i_{k}}).$ (17)
-For POMDPs, we estimate (16) with $K_{b}^{\nabla}$ state particles via
-$\textstyle\nabla_{a}\log p_{T}(\bar{b}^{-}\mid\bar{b},a)\approx\frac{J}{K_{b}^{\nabla}}\sum_{l=1}^{K_{b}^{\nabla}}\nabla_{a}\log p_{T}(s^{-,j_{l}}\mid s^{j_{l}},a),$ (18)
-where indices $j_{l}$ are sampled uniformly from $[1,\dots,J]$.
-In state-reward POMDPs, we estimate the immediate reward separately.
-We generate $K_{r}^{\nabla}$ posterior states by sampling $s_{new}^{j}\sim p_{T}(\cdot\mid\bar{b},a)$.
+^$\hat{\nabla}_{a}\hat{Q}(s,a)=\frac{1}{K_{O}^{\nabla}}\sum_{k=1}^{K_{O}^{\nabla}}\nabla_{a}\log p_{T}(s^{\prime,i_{k}}\mid s,a)\cdot(r(s,a,s^{\prime,i_{k}})+\gamma\hat{V}(s^{\prime,i_{k}}))+\nabla_{a}r(s,a,s^{\prime,i_{k}}).$ (148)
+In some of the domains we used direct state sampling to compute the immediate reward’s contribution to the gradient based on Theorem 1 Extended (Immediate Reward Gradient), which we found to be more effective.
+In this form, we generate $K_{r}^{\nabla}$ posterior states by sampling $s_{new}^{j}\sim p_{T}(\cdot\mid\bar{b},a)$.
 The MC estimate of (49) is then
-^$\textstyle\hat{\nabla}_{a}\hat{Q}(\bar{b},a)=\frac{1}{K_{r}^{\nabla}}\big{(}\sum_{i=j}^{K_{r}^{\nabla}}\nabla_{a}\log p_{T}(s_{new}^{j}\mid s,a)\cdot$
-$\textstyle r(s,a,s_{new}^{j})+\nabla_{a}r(s,a,s_{new}^{j})\big{)}+\big{(}\frac{\gamma}{K_{O}^{\nabla}}\sum_{k=1}^{K_{O}^{\nabla}}(\frac{J}{K_{b}^{\nabla}}\sum_{l=1}^{K_{b}^{\nabla}}$
-$\textstyle\nabla_{a}\log p_{T}(s^{-,i_{k},j_{l}}\mid s^{i_{k},j_{l}},a))\textstyle\cdot\hat{V}(\bar{b}^{\prime,i_{k}})\big{)}.$ (19)
-Gradient Optimization.
-We’ve found throughout our experiments that the Adam optimizer, utilizing step size normalization, was crucial for consistent behavior due to the high variance of the gradient estimates caused by MC estimation and extremely high variance of the propagated belief transition probability.
-A slight step size decay was beneficial as well for convergence, allowing to choose larger Adam step sizes.
+^$\hat{\nabla}_{a}\hat{Q}(\bar{b},a)=\frac{1}{K_{r}^{\nabla}}\big{(}\sum_{i=j}^{K_{r}^{\nabla}}\nabla_{a}\log p_{T}(s_{new}^{j}\mid s,a)\cdot r(s,a,s_{new}^{j})+\nabla_{a}r(s,a,s_{new}^{j})\big{)}$
+$+\big{(}\frac{\gamma}{K_{O}^{\nabla}}\sum_{k=1}^{K_{O}^{\nabla}}(\frac{J}{K_{b}^{\nabla}}\sum_{l=1}^{K_{b}^{\nabla}}\nabla_{a}\log p_{T}(s^{-,i_{k},j_{l}}\mid s^{i_{k},j_{l}},a))\cdot\hat{V}(\bar{b}^{\prime,i_{k}})\big{)}.$ (149)
+In the case of linearizing the importance weights update as in equation (12), the summation is over all observation children $s^{\prime}\in C(s,a)$.
+Since we do not sample them, we weight the summation by the importance weights $\omega_{a}(s^{\prime})$.
+In this case, the gradient estimator becomes:
+^$\hat{\nabla}_{a}\hat{Q}(s,a)=\eta(s,a)^{-1}\sum_{s^{\prime}\in C(s,a)}\omega_{a}(s^{\prime})n(s^{\prime})_{+1}$
+$\cdot\left(\nabla_{a}\log p_{T}(s^{\prime}\mid s,a)\cdot(r(s,a,s^{\prime})+\gamma\hat{V}(s^{\prime}))+\nabla_{a}r(s,a,s^{\prime})\right).$ (150)
+Similar estimators can be derived for the POMDP case, and for the MDP or POMDP cases when using Theorem 1 Extended (Immediate Reward Gradient).
+B.2 Action Update Rule
+In some domains, we found that adding a threshold on the number of children $\left|{C(s,a)}\right|\geq K_{\text{child}}^{visits}$ for action update to be performed was beneficial.
+This improves the estimate of the gradient, since it is the MIS estimator based on the sampled children.
+In order to control computational complexity, and also ensure in some domains that enough samples are added to a subtree before performing consecutive action updates, we set a hyperparameter $K_{\text{child}}^{\min}$ for action update to be performed only whenever $n(s,a)\text{ mod }K_{\text{child}}^{\min}=0$.
+B.3 Gradient Optimization
+We’ve found throughout our experiments that the Adam optimizer [27] that normalizes the step size was crucial for consistent behavior due to the high variance of the gradient estimates.
 We’ve fixed as a hyperparameter $K_{opt}$ the number of consecutive gradient optimization iterations before simulation.
 Together with the step size, this controls a trade-off between accuracy and computational complexity.
-Lastly, since action updates are significantly more costly than computing the gradient estimates, we set a threshold action distance for action update.
-After each gradient iteration, we update an accumulator of the gradient steps $a_{t}^{acc}$.
-Only after passing a threshold distance $\lVert{a_{t}^{acc}-a_{t}}\rVert>T_{d_{a}}$, we update $a_{t}\leftarrow a_{t}^{acc}$.
-This limits the number of action updates, and acts like a convergence criterion.
-Thresholds For Adding/Deleting State Nodes.
-Although the SN-MIS estimator normalizes the contribution of each posterior node, we would like to ensure that we have relevant samples after several action updates.
-As a relevance heuristic, we delete child state $s^{\prime,i}$ from $C(s,a)$ if $\omega^{i}<T_{\omega}^{del}$ after action update.
-Additionally, if $\omega^{i}<T_{\omega}^{add}$ for all $s^{\prime,i}\in C(s,a)$, we force to sample a new posterior node, overriding the progressive widening limitation.
-This ensures that we do not waste computational resources on states with low contribution to the estimator $\hat{Q}(s,a)$.
-Due to the possibility of deleting posterior children, care must be taken when updating visitation counters since they do not monotonically increment anymore.
-For this reason equations (9), (10), (11), (12), (13) are written using previous and updated visitation counts.
-4 Experiments
-Numerical simulations were carried out to evaluate the performance of AGMCTS in a POMDP setting, compared to state-of-the-art algorithms POMCPOW and PFT-DPW, whose implementations were taken as the standard in the POMDPs.jl framework [?].
-AGMCTS was implemented using POMDPs.jl as well.
-For gradient computations we used automatic differentiation with Enzyme.jl [?].
-The analysis and comparison was based on the code of [?], and similarly we used rollout-first-then-random action sampling.
-4.1 D-Continuous Light Dark
-(a) Middle of scenario, $t=5$.
-(b) End of scenario, $t=6$.
-Figure 3: AGMCTS at 2-Continuous Light Dark.
-The agent’s current state is the red dot, the current belief particles in orange, the next state in black, and the next belief particles in gray, the next observation is the green dot.
-The goal is the blue ring centered at $(0,2.5)$, $b_{0}$ is the red ring centered at $(0,0)$, and the beacon is the green star at $(2.5,0)$.
-The reward function for the posterior state is the blue-red heatmap drawn in the background.
-The shaded area is where the observation noise is $\sigma_{O}>5$.
-The chosen action branch of AGMCTS is drawn as the blue arrows, getting darker in hue with each action update.
-We can see that the optimized actions point closer to the goal from the current belief’s mean.
-This is a continuous extrapolation of the well-known Light-Dark POMDP.
-In this domain, $\mathcal{S}=\mathcal{O}=\mathbb{R}^{D}$.
-The agent starts at a random position on the sphere centered at the origin $S^{D-1}_{0.5}$.
-The goal position is located at $s_{g}=(\boldsymbol{0}_{D-1},2.5)$.
-A beacon is located to the side at $s_{b}=(2.5,\boldsymbol{0}_{D-1})$.
-$T=0.2$ is the goal tolerance, and the POMDP terminates if $\lVert{s-s_{g}}\rVert<T$.
-The horizon is $L=6$ and $\gamma=0.99$.
-The action space is the ball $B_{1.5}(\boldsymbol{0}_{D})$.
-The transition model is a simple Gaussian with added noise: $s^{\prime}\sim\mathcal{N}(s+a,(0.025)^{2}\mathbf{I})$.
-The observations are the agent’s relative position to the beacon, with a Gaussian added noise that increases with distance: $o\sim\mathcal{N}(s-s_{b},(\sigma_{O}(\lVert{s-s_{b}}\rVert))^{2}\mathbf{I})$, with noise function $\sigma_{O}(x)=\min\{15,0.01(x+x^{8})\}$.
-The reward function is $r(s,a,s^{\prime})=10\cdot g_{1}(d)-2\cdot g_{2}(d)-0.02d^{2}$, where $d\mathbin{=}\lVert{s^{\prime}-s_{g}}\rVert$.
-$g_{1}$ is a Gaussian centered at 00 with width $\frac{T}{2}$, $g_{2}$ is a Gaussian centered at $5T$ with width $T$.
-This reward function only depends on the posterior state.
-The agent needs to carefully target the goal as the reward decreases quickly from the goal, and advancing at random towards it will incur a penalty due to the moat $g_{2}$.
-To simulate more difficult domains, we include a noisy rollout policy.
-The rollout heads directly to $s_{g}$, but the computed action has an added noise sampled from $\mathcal{N}(\mathbf{0},\sigma_{r}^{2}\mathbf{I})$.
-4.2 Performance Evaluation
-[…]
-D $\sigma_{r}$ POMCPOW PFT-DPW AGMCTS
-2 0.1 $\boldsymbol{6.62}\pm 0.07$ $5.71\pm 0.08$ $\boldsymbol{6.67}\pm 0.06$
-2 0.2 $\boldsymbol{6.23}\pm 0.07$ $4.74,0.08$ $5.96\pm 0.08$
-2 0.3 $\boldsymbol{5.61}\pm 0.08$ $4.01\pm 0.09$ $\boldsymbol{5.49}\pm 0.09$
-3 0.1 $3.42\pm 0.08$ $3.63\pm 0.08$ $\boldsymbol{5.47}\pm 0.07$
-3 0.2 $1.87\pm 0.08$ $1.79\pm 0.08$ $\boldsymbol{3.82}\pm 0.1$
-3 0.3 $1.06\pm 0.07$ $0.29\pm 0.06$ $\boldsymbol{1.58}\pm 0.1$
-4 0.1 $2.58\pm 0.07$ $2.36\pm 0.07$ $\boldsymbol{4.26},0.08$
-4 0.2 $0.53\pm 0.05$ $0.2\pm 0.05$ $\boldsymbol{1.53}\pm 0.09$
-4 0.3 $\boldsymbol{-0.36}\pm 0.03$ $-0.86\pm 0.02$ $-1.13\pm 0.04$
-Table 1: Performance results for the different algorithms at D-Continuous Light Dark for $D=2,3,4$ dimensions.
-AGMCTS and PFT-DPW were run for 500 simulations, and POMCPOW for an equivalent time to PFT-DPW.
-The rollout policy’s standard deviation was $\sigma_{r}=0.1$ during hyperparameter tuning, and then the algorithms were also tested for $\sigma_{r}=0.2,0.3$.
-The reported results are mean returns over 1000 scenarios $\pm$ standard error.
-AGMCTS achieved the best performance overall in the different domains.
-5 Conclusions
-We introduce a novel methodology for considering dynamic action branches in MCTS in continuous (PO)MDP settings.
-The MIS tree allows action updates while keeping action-value estimates consistent.
-The new gradient derivation allows computing the action-value gradient in continuous POMDPs, and is supported by the MIS tree for its practical estimation.
-Our novel planning algorithm AGMCTS incorporates these two novelties together, and with additional algorithmic heuristics, is shown to achieve improved performance on a per-simulation budget.
-Our experiments suggest that gradient-based action updates may be more effective at higher dimensions, and where domain knowledge is less accurate.
-Practical Considerations.
-AGMCTS requires a transition likelihood model, both for MIS updates and for the gradient computation.
-It is application-dependent whether the computational cost of the gradient computation is worth the improvement in performance, compared to using more simulations.
-In our experiments, AGMCTS had significantly slower planning sessions compared to PFT-DPW.
-Each algorithm was run in a single thread, yet particle belief algorithms such as AGMCTS and PFT-DPW are highly parallelizable.
-Furthermore, AGMCTS could be combined with more sophisticated heuristics, like Voronoi-based action sampling and Bayesian optimization methods, which might boost performance.
-Finally, it is important to note that many algorithmic choices can be tailored depending on the application, like the MIS heuristic and the gradient optimizer, which may significantly affect run times.
-Future Work.
-The gradient computation and reuse that we’ve shown for POMDPs are based on the propagated belief, which requires full belief trajectories that might be computationally demanding in some domains.
-We plan to extend our methods efficiently for state trajectories as well.
-For the action selection we’ve used the standard UCT, however it is not obvious that it remains theoretically correct for the MIS estimator.
-It is likely that the exploration bonus varies at action updates, and further investigation is required.
-Finally, AGMCTS requires significant hyperparameter tuning, including DPW parameters and parameters related to action optimization.
-It is future research to find algorithms that are more general across different domains.
-Acknowledgments
-This research was partially supported by the Israel Science Foundation (ISF).
-Appendix A Proofs and Derivations
-A.1 Lemma 1
-Lemma 1.
-Let $\bar{b}^{-}=((s^{-,j},w^{j}))_{j=1}^{J}$ be generated from $\bar{b}=((s^{j},w^{j}))_{j=1}^{J}$ by SIS or SIR.
-Hence, the $j$th state particle was generated via $s^{-,j}\sim p_{T}(\cdot\mid s^{j},a)$, and the weight was copied, i.e.
-$w^{-,j}=w^{j}$.
-The ordered particle belief likelihood is then:
-$\textstyle p_{T}(\bar{b}^{-}\mid\bar{b},a)=\prod_{i=1}^{J}p_{T}(s^{-,j}\mid s^{j},a).$ (20)
-Proof.
-In the SIS and SIR algorithms, each particle is sampled $s^{-,j}\sim p_{T}(\cdot\mid s^{j},a)$ independently for each $j=1,\ldots,J$. Therefore:
-$p(\bar{b}^{-}\mid\bar{b},a)$ (21)
-$=p(((s^{-,j},w^{-,j}))_{j=1}^{J}\mid((s^{l},w^{l}))_{l=1}^{J},a),$ (22)
-due to the beliefs being ordered, and the independent sampling, we can factorize as:
-$=\prod_{i=1}^{J}p(((s^{-,j},w^{-,j}))\mid((s^{l},w^{l}))_{l=1}^{J},a)$ (23)
-$=\prod_{i=1}^{J}p(((s^{-,j},w^{-,j}))\mid(s^{j},w^{j}),a).$ (24)
-Finally, since $w^{-,j}=w^{j}$, we can eliminate the weights to obtain
-$p(\bar{b}^{-}\mid\bar{b},a)=\prod_{i=1}^{J}p_{T}(s^{-,j}\mid s^{j},a),$ (25)
-and we denote $p_{T}(\bar{b}^{-}\mid\bar{b},a)\triangleq p(\bar{b}^{-}\mid\bar{b},a)$. ∎
-A.2 Derivations of MIS Tree Update Equations
-Here we show the derivation of the update equation, and also show their numerically stable forms using log likelihoods for probability values.
-We denote the regular and the weighted log-sum-exp function as $LSE$, differentiated by having one or two vector inputs, and they are defined as
-$LSE(\boldsymbol{x}_{1,\ldots,n})\triangleq\log\left(\sum_{i=1}^{n}\exp(\boldsymbol{x}_{i})\right),$ (26)
-$LSE(\boldsymbol{x}_{1,\ldots,n},\boldsymbol{a}_{1,\ldots,n})\triangleq\log\left(\sum_{i=1}^{n}\exp(\boldsymbol{x}_{i})\cdotp\boldsymbol{a}_{i}\right).$ (27)
-To handle negative weights, (27) can be defined to compute the log of absolute value and sign separately of the log argument, such that the sign is multiplied back after exponentiation if required.
-We also denote a vector with brackets []delimited-[][$\cdot$][ ⋅ ].
-We remind the visitation counters relationships:
-$n(s)=\sum_{C(s)}n(s,a^{i}),\;n(s,a)=\sum_{C(s,a)}n(s^{\prime,i})_{+1},$ (28)
-the state value estimator:
-$\hat{V}(s)\triangleq\sum_{C(s)}\frac{n(s,a^{i})}{n(s)}\hat{Q}(s,a^{i}),$ (29)
-the self normalized MIS estimator:
-$\eta\triangleq\sum_{C(s,a)}\omega^{i}n(s^{\prime,i})_{+1},$ (30)
-$\hat{V}_{f}(s,a)=\eta^{-1}\sum_{C(s,a)}\omega^{i}n(s^{\prime,i})_{+1}\hat{V}(s^{\prime,i}),$ (31)
-and we also remind that we denoted $M\triangleq\lvert{C(s,a)}\rvert$.
-We define the following vectors for each a given action node $(s,a)$ or $(\bar{b},a)$:
-  1. 1.
-    Child state visitation counters.
-    MDP:
-    $\boldsymbol{n}\triangleq(n(s^{\prime,i})_{+1})_{i=1}^{M},$ (32)
-    PB-MDP:
-    $\boldsymbol{n}\triangleq(n(\bar{b}^{\prime,i})_{+1})_{i=1}^{M}.$ (33)
-  2. 2.
-    Target transition likelihoods.
-    MDP:
-    $\boldsymbol{p}=(p_{T}(s^{\prime,i}\mid s,a))_{i=1}^{M},$ (34)
-    PB-MDP:
-    $\boldsymbol{p}=(p_{T}(\bar{b}^{-,i}\mid\bar{b},a))_{i=1}^{M}.$ (35)
-  3. 3.
-    Proposal transition likelihoods.
-    MDP:
-    $\boldsymbol{q}=(p_{T}(s^{\prime,i}\mid s,a_{prop}(s^{\prime,i})))_{i=1}^{M},$ (36)
-    PB-MDP:
-    $\boldsymbol{q}=(p_{T}(\bar{b}^{-,i}\mid\bar{b},a_{prop}(\bar{b}^{-,i})))_{i=1}^{M}.$ (37)
-  4. 4.
-    Immediate rewards.
-    MDP:
-    $\boldsymbol{r}=(r(s,a,s^{\prime,i}))_{i=1}^{M}$ (38)
-    PB-MDP:
-    $\boldsymbol{r}=(r(\bar{b},a,\bar{b}^{\prime,i}))_{i=1}^{M}.$ (39)
-  5. 5.
-    Future value estimates.
-    MDP:
-    $\boldsymbol{v}=(\hat{V}(s^{\prime,i}))_{i=1}^{M}$ (40)
-    PB-MDP:
-    $\boldsymbol{v}=(\hat{V}(\bar{b}^{\prime,i}))_{i=1}^{M}$ (41)
-From here on the PB-MDP is the same up to different notations as the MDP case, and we will show only the latter.
-Using these notations the immediate reward and future value estimates can be compactly written:
-$\eta=\boldsymbol{1}^{T}(\boldsymbol{n}\odot\boldsymbol{p}\oslash\boldsymbol{q}),$ (42)
-$\hat{r}(s,a)=\eta^{-1}\boldsymbol{1}^{T}(\boldsymbol{n}\odot\boldsymbol{p}\odot\boldsymbol{r}\oslash\boldsymbol{q}),$ (43)
-$\hat{V}_{f}(s,a)=\eta^{-1}\boldsymbol{1}^{T}(\boldsymbol{n}\odot\boldsymbol{p}\odot\boldsymbol{v}\oslash\boldsymbol{q}),$ (44)
-[…]
-We store the following values:
-  1. 1.
-    $l(\eta)$.
-  2. 2.
-    Log likelihoods $l(\boldsymbol{p})$ and $l(\boldsymbol{q})$.
-  3. 3.
-    $\boldsymbol{n}^{\prime}$ and previous visitation counters $\boldsymbol{n}$.
-  4. 4.
-    $\boldsymbol{v}^{\prime}$ and previous future value estimates $\boldsymbol{v}$.
-The equations for $\hat{r}(s,a)$ are analogous to $\hat{V}_{f}(s,a)$, and we will not show their derivation explicitly.
-Action Backpropagation.
-Let $s^{\prime,j}$ be an updated node.
-Hence, we updated $\boldsymbol{n}_{j}$ to $\boldsymbol{n}^{\prime}_{j}$ and $\boldsymbol{v}_{j}$ to $\boldsymbol{v}^{\prime}_{j}$.
-We update $n(s,a)$ by (28) and perform
-$\eta^{\prime}=\eta+\frac{\boldsymbol{p}_{j}}{\boldsymbol{q}_{j}}\left(\boldsymbol{n}_{j}^{\prime}-\boldsymbol{n}_{j}\right),$ (45)
-$\hat{V}_{f}^{\prime}(s,a)=\frac{1}{\eta^{\prime}}\left(\eta\hat{V}_{f}(s,a)+\frac{\boldsymbol{p}_{j}}{\boldsymbol{q}_{j}}(\boldsymbol{n}_{j}^{\prime}\boldsymbol{v}_{j}^{\prime}-\boldsymbol{n}_{j}\boldsymbol{v}_{j})\right).$ (46)
-Therefore, the log-form update equations are:
-$l(\eta^{\prime})=LSE\left([l(\eta),l(\boldsymbol{p}_{j})-l(\boldsymbol{q}_{j})+l(\boldsymbol{n}_{j}^{\prime}-\boldsymbol{n}_{j})]\right)$ (47)
-$\hat{V}_{f}^{\prime}(s,a)=\exp\Big{(}LSE\big{(}[l(\eta),l(\boldsymbol{p}_{j})-l(\boldsymbol{q}_{j})],$
-$[\hat{V}_{f}(s,a),\boldsymbol{n}_{j}^{\prime}\boldsymbol{v}_{j}^{\prime}-\boldsymbol{n}_{j}\boldsymbol{v}_{j}]\big{)}-l(\eta^{\prime})\Big{)}.$ (48)
-Proof.
-We start with the update for $\eta^{\prime}$:
-$\eta^{\prime}-\eta$ (49)
-$=\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}^{\prime}\boldsymbol{p}_{i}}{\boldsymbol{q}_{i}}-\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}}{\boldsymbol{q}_{i}}$ (50)
-$=\frac{\boldsymbol{p}_{j}}{\boldsymbol{q}_{j}}(\boldsymbol{n}_{j}^{\prime}-\boldsymbol{n}_{j}),$ (51)
-due to all terms being equal except for $\boldsymbol{n}_{j}^{\prime}\neq\boldsymbol{n}_{j}$. Similarly, for $\hat{V}_{f}(s,a)$:
-$\hat{V}_{f}^{\prime}(s,a)-\frac{\eta}{\eta^{\prime}}\hat{V}_{f}(s,a)$ (52)
-$=\frac{1}{\eta^{\prime}}\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}^{\prime}\boldsymbol{p}_{i}\boldsymbol{v}_{i}^{\prime}}{\boldsymbol{q}_{i}}-\frac{\eta}{\eta^{\prime}}\frac{1}{\eta}\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}\boldsymbol{v}_{i}}{\boldsymbol{q}_{i}}$ (53)
-$=\frac{1}{\eta^{\prime}}\left(\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}^{\prime}\boldsymbol{p}_{i}\boldsymbol{v}_{i}^{\prime}}{\boldsymbol{q}_{i}}-\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}\boldsymbol{v}_{i}}{\boldsymbol{q}_{i}}\right)$ (54)
-$=\frac{1}{\eta^{\prime}}\frac{\boldsymbol{p}_{j}}{\boldsymbol{q}_{j}}(\boldsymbol{n}_{j}^{\prime}\boldsymbol{v}_{j}^{\prime}-\boldsymbol{n}_{j}\boldsymbol{v}_{j}),$ (55)
-due to all terms being equal except for $\boldsymbol{v}^{\prime}_{j}\neq\boldsymbol{v}_{j}$, $\boldsymbol{n}^{\prime}_{j}\neq\boldsymbol{n}_{j}$. ∎
-State Expansion.
-This case is the same as action backpropagation, where the updated index is a new summand at index $j=M+1$. The same equations apply with $n(s^{\prime,j})=0$.
-Proof.
-We append new terms to all vectors: $\boldsymbol{n}_{M+1}$, $\boldsymbol{p}_{M+1}$, $\boldsymbol{v}_{M+1}$, $\boldsymbol{q}_{M+1}$. Update to $\eta$:
-$\eta^{\prime}-\eta$ (56)
-$=\sum_{i=1}^{M+1}\frac{\boldsymbol{n}_{i}^{\prime}\boldsymbol{p}^{\prime}_{i}}{\boldsymbol{q}^{\prime}_{i}}-\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}}{\boldsymbol{q}_{i}}$ (57)
-$=\frac{\boldsymbol{n}_{M+1}^{\prime}\boldsymbol{p}^{\prime}_{M+1}}{\boldsymbol{q}^{\prime}_{M+1}},$ (58)
-due to all terms being equal except for index $M+1$.
-This corresponds to (51) with $\boldsymbol{n}_{M+1}=0$, after initializing $\boldsymbol{p}^{\prime}_{M+1}=\boldsymbol{p}_{M+1}$ and $\boldsymbol{q}^{\prime}_{M+1}=\boldsymbol{q}_{M+1}$ Similarly, for $\hat{V}_{f}(s,a)$:
-$\hat{V}_{f}^{\prime}(s,a)-\frac{\eta}{\eta^{\prime}}\hat{V}_{f}(s,a)$ (59)
-$=\frac{1}{\eta^{\prime}}\sum_{i=1}^{M+1}\frac{\boldsymbol{n}_{i}^{\prime}\boldsymbol{p}_{i}\boldsymbol{v}_{i}^{\prime}}{\boldsymbol{q}_{i}}-\frac{\eta}{\eta^{\prime}}\frac{1}{\eta}\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}\boldsymbol{v}_{i}}{\boldsymbol{q}_{i}}$ (60)
-$=\frac{1}{\eta^{\prime}}\left(\sum_{i=1}^{M+1}\frac{\boldsymbol{n}_{i}^{\prime}\boldsymbol{p}_{i}\boldsymbol{v}_{i}^{\prime}}{\boldsymbol{q}_{i}}-\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}\boldsymbol{v}_{i}}{\boldsymbol{q}_{i}}\right)$ (61)
-$=\frac{1}{\eta^{\prime}}\frac{\boldsymbol{p}^{\prime}_{M+1}}{\boldsymbol{q}^{\prime}_{M+1}}\boldsymbol{n}_{M+1}^{\prime}\boldsymbol{v}_{M+1}^{\prime},$ (62)
-which again corresponds to (55) with $\boldsymbol{n}_{M+1}=0$. ∎
-Action Update.
-Let $a^{\prime}$ be the new action. Here, we have to calculate $\omega^{i}$ for all $s^{\prime,i}\in C(s,a)$ in $O(M)$ time. We update
-$(\omega^{i})^{\prime}=p_{T}(s^{\prime,i}\mid s,a^{\prime})(p_{T}(s^{\prime,i}\mid s,a(s^{\prime,i})))^{-1},$ (63)
-and recalculate (30) and (31) with the new ($\omega^{i})^{\prime}$, which are also $O(M)$ time operations.
-For the immediate reward, new rewards must be computed for the new triplets $(s,a^{\prime},s^{\prime,i})$, and the update is analogous.
-Therefore, the log-form equations are:
-$l(\eta^{\prime})=LSE(l(\boldsymbol{n})\oplus l(\boldsymbol{p}^{\prime})\ominus l(\boldsymbol{q}))$ (64)
-$\hat{V}_{f}^{\prime}(s,a)=\exp\left(LSE(l(\boldsymbol{p}^{\prime})\ominus l(\boldsymbol{q}),\boldsymbol{n}\odot\boldsymbol{v})-l(\eta^{\prime})\right)$ (65)
-Proof.
-This update corresponds to computing a new target likelihood vector $\boldsymbol{p}$. Therefore, the update equations follow from the definitions:
-$\eta^{\prime}=\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}^{\prime}}{\boldsymbol{q}_{i}},$ (66)
-$\hat{V}_{f}^{\prime}(s,a^{\prime})=\frac{1}{\eta^{\prime}}\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}^{\prime}\boldsymbol{v}_{i}}{\boldsymbol{q}_{i}}.$ (67)
-Terminal State Backpropagation.
-Let $s$ be a terminal node. Its value estimate is based only on rollouts. Hence, for a new rollout value $v^{\prime}$, we perform a running average update
-$\hat{V}^{\prime}(s)=\hat{V}(s)+\frac{n^{\prime}(s)-n(s)}{n^{\prime}(s)_{+1}}(v^{\prime}-\hat{V}(s)),$ (68)
-where $n^{\prime}(s)$ is $n(s)$ plus the number of rollouts (usually 1).
-Proof.
-This is a simple running average update for the mean estimator:
-$\hat{V}(s)=\frac{1}{n(s)_{+1}}\sum_{i=1}^{n(s)_{+1}}v^{i},$ (69)
-for all $v^{i}$ rollout values from the terminal node $s$. ∎
-Non-Terminal State Backpropagation.
-Let $(s,a^{j})$ be an updated node. Hence, we updated $n(s,a^{j})$ to $n^{\prime}(s,a^{j})$ and $\hat{Q}(s,a^{j})$ to $\hat{Q}^{\prime}(s,a^{j})$. We update $n^{\prime}(s)$ by (28) and perform
-$\hat{V}^{\prime}(s)=(n^{\prime}(s))^{-1}\Big{(}n(s)\hat{V}(s)+n^{\prime}(s,a^{j})\hat{Q}^{\prime}(s,a^{j})$
-${}-n(s,a^{j})\hat{Q}(s,a^{j})\Big{)}.$ (71)
-Proof.
-Following the definition of the state value estimator (29), the updated estimator is:
-$\hat{V}^{\prime}(s)=\sum_{C(s)}\frac{n^{\prime}(s,a^{i})}{n^{\prime}(s)}\hat{Q}^{\prime}(s,a^{i}).$ (72)
-Therefore:
-$\hat{V}^{\prime}(s)-\frac{n(s)}{n^{\prime}(s)}\hat{V}(s)$ (73)
-$=\sum_{C(s)}\frac{n^{\prime}(s,a^{i})}{n^{\prime}(s)}\hat{Q}^{\prime}(s,a^{i})-\frac{n(s)}{n^{\prime}(s)}\sum_{C(s)}\frac{n(s,a^{i})}{n(s)}\hat{Q}(s,a^{i})$ (74)
-$=\frac{1}{n^{\prime}(s)}\sum_{C(s)}(n^{\prime}(s,a^{i})\hat{Q}^{\prime}(s,a^{i})-n(s,a^{i})\hat{Q}(s,a^{i})).$ (75)
-All terms are equal except for the updated node $j$. Hence, we obtain
-$\hat{V}^{\prime}(s)-\frac{n(s)}{n^{\prime}(s)}\hat{V}(s)$
-$=\frac{1}{n^{\prime}(s)}\left(n^{\prime}(s,a^{j})\hat{Q}^{\prime}(s,a^{j})-n(s,a^{j})\hat{Q}(s,a^{j})\right).$ (76)
-A.3 Theorem 1
-We first highlight that the required ”well-behaved” conditions for the theorem are for changing the order of derivatives and integrals by the Leibniz integral rule.
-A sufficient condition is that the integrand and all its partial derivatives are continuous in all variables, which is satisfied when both $p_{T},r$ satisfy these conditions.
-Stronger versions of this theorem exist, such as only requiring that the partial derivatives exist almost everywhere.
-As this is not the focus of our discussion, we leave that for future work.
-Additionally, as stated in the main paper, it must be satisfied that $p_{T}(s^{\prime}\mid s,a)=0$ implies $p_{T}(s^{\prime}\mid s,a^{\prime})=0$.
-Theorem 1.
-For well-behaved $p_{T}$, $r$, the MDP action gradient of the action-value function is given by the expectation
-$\nabla_{a^{\prime}}Q_{t}^{\pi}(s,a^{\prime})=\mathbb{E}_{s^{\prime}\mid s,a}\bigg{[}\frac{p_{T}(s^{\prime}\mid s,a^{\prime})}{p_{T}(s^{\prime}\mid s,a)}$
-$(\nabla_{a^{\prime}}\log p_{T}(s^{\prime}\mid s,a^{\prime})$
-$\cdot(r(s,a^{\prime},s^{\prime})+\gamma V_{t+1}^{\pi}(s^{\prime}))+\nabla_{a^{\prime}}r(s,a^{\prime},s^{\prime}))\bigg{]}.$ (77)
-Additionally, for well-behaved $p_{T}$, $p_{O}$, $r$, the POMDP action gradient of the action-value function can be computed by replacing posterior belief likelihood terms with propagated belief likelihood terms, i.e.
-$b_{t+1}^{-}$ instead of $b_{t+1}$ in all $p_{T}$ terms.
-Proof.
-We start with the MDP case.
-$\nabla_{a^{\prime}}Q_{t}^{\pi}(s,a^{\prime})$ (78)
-$=\nabla_{a^{\prime}}\mathbb{E}_{s{{}^{\prime}}\mid s,a}\bigg{[}\frac{p(s{{}^{\prime}}\mid s,a{{}^{\prime}})}{p(s{{}^{\prime}}\mid s,a)}(r(s,a{{}^{\prime}},s{{}^{\prime}})+\gamma V_{t+1}^{\pi}(s{{}^{\prime}}))\bigg{]},$ (79)
-Applying the Leibniz integral rule:
-$=\mathbb{E}_{s{{}^{\prime}}\mid s,a}\bigg{[}\nabla_{a^{\prime}}\frac{p(s{{}^{\prime}}\mid s,a{{}^{\prime}})}{p(s{{}^{\prime}}\mid s,a)}\cdot(r(s,a{{}^{\prime}},s{{}^{\prime}})+\gamma V_{t+1}^{\pi}(s{{}^{\prime}}))$
-$+\frac{p(s{{}^{\prime}}\mid s,a{{}^{\prime}})}{p(s{{}^{\prime}}\mid s,a)}\cdot\nabla_{a^{\prime}}(r(s,a{{}^{\prime}},s{{}^{\prime}})+\gamma V_{t+1}^{\pi}(s{{}^{\prime}}))\bigg{]}.$ (80)
-For the first term, we use the log-gradient substitution:
-$\nabla_{a^{\prime}}p(s{{}^{\prime}}\mid s,a{{}^{\prime}})=p(s{{}^{\prime}}\mid s,a{{}^{\prime}})\cdot\nabla_{a^{\prime}}\log p(s{{}^{\prime}}\mid s,a{{}^{\prime}}),$ (81)
-and for the second term, since $V_{t+1}^{\pi}(s^{\prime})$ is not a function of $a^{\prime}$, we obtain:
-$\nabla_{a^{\prime}}(r(s,a{{}^{\prime}},s{{}^{\prime}})+\gamma V_{t+1}^{\pi}(s{{}^{\prime}}))=\nabla_{a^{\prime}}r(s,a{{}^{\prime}},s{{}^{\prime}}).$ (82)
-By gathering terms we obtain the result (77).
-We now show the POMDP case. We assume that the same conditions are satisfied for $p_{T},p_{O},r$. We first show the propagated belief trick.
-$Q_{t}^{\pi}(b,a)=\mathbb{E}_{b^{\prime}\mid b,a}\left[r(b,a,b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime})\right]$ (83)
-$=\int_{b^{\prime}}p(b^{\prime}\mid b,a)(r(b,a,b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime}))\mathop{}\!\mathrm{d}b^{\prime},$ (84)
-and we continue by marginalizing over the propagated belief $b^{-}$ and the observation $o$:
-$=\int_{b^{-}}\int_{o}\int_{b^{\prime}}p(b^{-},o,b^{\prime}\mid b,a)$
-$(r(b,a,b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime}))\mathop{}\!\mathrm{d}b^{\prime}\mathop{}\!\mathrm{d}o\mathop{}\!\mathrm{d}b^{-}.$ (85)
-By its definition, the propagated belief $b^{-}$ is the generated belief by conditioning on the history without the last observation.
-Therefore, using the chain rule and removing variables that are conditionally independent, we obtain:
-[…]
-$p(b^{-}\mid b,a)(r(b,a,b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime}))\mathop{}\!\mathrm{d}b^{\prime}\mathop{}\!\mathrm{d}o\mathop{}\!\mathrm{d}b^{-}$ (86)
-$=\int_{b^{-}}\int_{o}\int_{b^{\prime}}p(b^{\prime}\mid o,b^{-})p(o\mid b^{-})p(b^{-}\mid b,a)$
-$(r(b,a,b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime}))\mathop{}\!\mathrm{d}b^{\prime}\mathop{}\!\mathrm{d}o\mathop{}\!\mathrm{d}b^{-}$ (87)
-$=\mathbb{E}_{b^{-}\mid b,a}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}\left[r(b,a,b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime})\right].$ (88)
-We introduce an importance ratio over the propagated belief to calculate the derivative for proposal action $a$ and target action $a^{\prime}$:
-$Q_{t}^{\pi}(b,a^{\prime})=\mathbb{E}_{b^{-}\mid b,a}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}\bigg{[}\frac{p_{T}(b^{-}\mid b,a^{\prime})}{p_{T}(b^{-}\mid b,a)}$
-$(r(b,a^{\prime},b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime}))\bigg{]},$ (89)
-where we denoted $p_{T}(b^{-}\mid b,a^{\prime})\triangleq p(b^{-}\mid b,a^{\prime})$ as the transition model for the belief MDP.
-Similarly to the derivation of the MDP case, we can apply the Leibniz integral rule to obtain the result:
-$\nabla_{a^{\prime}}Q_{t}^{\pi}(b,a^{\prime})=\mathbb{E}_{b^{-}\mid b,a}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}\bigg{[}\frac{p_{T}(b^{-}\mid b,a^{\prime})}{p_{T}(b^{-}\mid b,a)}$
-$(\nabla_{a^{\prime}}\log p_{T}(b^{-}\mid b,a^{\prime})$
-$\cdot(r(b,a^{\prime},b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime}))+\nabla_{a^{\prime}}r(b,a^{\prime},b^{\prime}))\bigg{]}.$ (90)
-A.4 Theorem 2
-We highlight that as an extension of Theorem 1, it requires the same conditions.
-Additionally, for the simplification of the immediate reward expectation, we make use of Fubini’s theorem for changing integration orders.
-Sufficient conditions are that there exists some $R\in\mathbb{R}$ such that $\sup_{s,a,s^{\prime}}\lvert{r(s,a,s^{\prime})}\rvert<R$, and that $r,p_{T},p_{O}$ are all continuous.
-Theorem 2.
-For a state-reward POMDP, the action gradient of the action-value function is given by:
-$\nabla_{a^{\prime}}Q_{t}^{\pi}(b,a)=\mathbb{E}_{s\mid b}\mathbb{E}_{s^{\prime}\mid s,a^{\prime}}[\nabla_{a^{\prime}}\log p_{T}(s^{\prime}\mid s,a^{\prime})\cdot r(s,a^{\prime},s^{\prime})$
-$+\nabla_{a^{\prime}}r(s,a^{\prime},s^{\prime})]+\gamma\mathbb{E}_{b^{-}\mid b,a}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}\bigg{[}\frac{p_{T}(b^{-}\mid b,a^{\prime})}{p_{T}(b^{-}\mid b,a)}$
-$\nabla_{a^{\prime}}\log p_{T}(b^{-}\mid b,a^{\prime})\cdot V_{t+1}^{\pi}(b^{\prime})\bigg{]}.$ (91)
-Proof.
-The central part of this proof is the decomposition of the belief action-value function to an expectation over state trajectories of immediate rewards, and an expectation over belief trajectories of future values.
-Afterwards, we apply the Leibniz integral rule similarly to Theorem 1.
-We start with the belief action-value function:
-$Q_{t}^{\pi}(b,a)=\mathbb{E}_{b^{\prime}\mid b,a}\left[r(b,a,b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime})\right]$ (92)
-$\mathbb{E}_{b^{\prime}\mid b,a}\left[r(b,a,b^{\prime})\right]+\gamma\mathbb{E}_{b^{\prime}\mid b,a}\left[V_{t+1}^{\pi}(b^{\prime})\right].$ (93)
-We expand the first expectation, using the chain rule and then Fubini’s theorem to cancel the expectation over the observation and posterior belief:
-$\mathbb{E}_{b^{\prime}\mid b,a}\left[r(b,a,b^{\prime})\right]$ (94)
-$=\mathbb{E}_{b^{-}\mid b,a}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}\left[\mathbb{E}_{s,s^{\prime}\mid b,a,b^{\prime}}[r(s,a,s^{\prime})]\right]$ (95)
-$=\int_{b^{-}}\int_{o}\int_{b^{\prime}}\int_{s}\int_{s^{\prime}}p(s,s^{\prime},o,b^{-},b^{\prime}\mid b,a)$
-$r(s,a,s^{\prime})\mathop{}\!\mathrm{d}s^{\prime}\mathop{}\!\mathrm{d}s\mathop{}\!\mathrm{d}b^{\prime}\mathop{}\!\mathrm{d}o\mathop{}\!\mathrm{d}b^{-}$ (96)
-$=\int_{b^{-}}\int_{o}\int_{b^{\prime}}\int_{s}\int_{s^{\prime}}p(o,b^{\prime},b^{-}\mid s,s^{\prime},b,a)p(s,s^{\prime}\mid b,a)$
-$r(s,a,s^{\prime})\mathop{}\!\mathrm{d}s^{\prime}\mathop{}\!\mathrm{d}s\mathop{}\!\mathrm{d}b^{\prime}\mathop{}\!\mathrm{d}o\mathop{}\!\mathrm{d}b^{-}$ (97)
-$=\int_{s}\int_{s^{\prime}}p(s,s^{\prime}\mid b,a)r(s,a,s^{\prime})$
-(∫$\cancel{\left(\int_{b^{-}}\int_{o}\int_{b^{\prime}}p(o,b^{\prime},b^{-}\mid s,s^{\prime},b,a)\mathop{}\!\mathrm{d}b^{\prime}\mathop{}\!\mathrm{d}o\mathop{}\!\mathrm{d}b^{-}\right)}\mathop{}\!\mathrm{d}s^{\prime}\mathop{}\!\mathrm{d}s$ (98)
-[…]
-$=\mathbb{E}_{s\mid b}\mathbb{E}_{s^{\prime}\mid s,a}[r(s,a,s^{\prime})].$ (100)
-Now, we compute $\nabla_{a^{\prime}}$ for the terms obtained.
-The gradient of the immediate reward expectation, when the proposal action is $a$ and the target action is $a^{\prime}$, is given by (skipping some steps similar to Theorem 1):
-$\nabla_{a^{\prime}}\mathbb{E}_{s\mid b}\mathbb{E}_{s^{\prime}\mid s,a}\bigg{[}\frac{p_{T}(s^{\prime}\mid s,a^{\prime})}{p_{T}(s^{\prime}\mid s,a)}r(s,a^{\prime},s^{\prime})\bigg{]}$ (101)
-$=\mathbb{E}_{s\mid b}\mathbb{E}_{s^{\prime}\mid s,a}\bigg{[}\frac{p_{T}(s^{\prime}\mid s,a^{\prime})}{p_{T}(s^{\prime}\mid s,a)}\big{(}\nabla_{a^{\prime}}\log p_{T}(s^{\prime}\mid s,a^{\prime})$
-$\cdot r(s,a^{\prime},s^{\prime})+\nabla_{a^{\prime}}r(s,a^{\prime},s^{\prime})\big{)}\bigg{]},$ (102)
-and the importance ratio cancels out for $a=a^{\prime}$.
-Next, the gradient of the future value expectation is simply:
-$\nabla_{a^{\prime}}\mathbb{E}_{b^{\prime}\mid b,a}\left[V_{t+1}^{\pi}(b^{\prime})\right]$ (103)
-$=\nabla_{a^{\prime}}\mathbb{E}_{b^{-}\mid b,a}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}\bigg{[}\frac{p_{T}(b^{-}\mid b,a^{\prime})}{p_{T}(b^{-}\mid b,a)}V_{t+1}^{\pi}(b^{\prime})\bigg{]}$ (104)
-$=\mathbb{E}_{b^{-}\mid b,a}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}\bigg{[}\frac{p_{T}(b^{-}\mid b,a^{\prime})}{p_{T}(b^{-}\mid b,a)}$
-$\nabla_{a^{\prime}}\log p_{T}(b^{-}\mid b,a^{\prime})\cdot V_{t+1}^{\pi}(b^{\prime})\bigg{]}.$ (105)
-Combining the two results, we obtain the final expression (91). ∎
-Appendix B Experiments
-B.1 AGMCTS Implementation Details
-Optimizer
-For the Adam optimizer, we used the hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$.
-The learning rate $\eta_{Adam}$ was determined by cross entropy optimization (CE), described later, in conjunction with the optimization over the other algorithm hyperparameters.
-In our implementation we used an exponential decay on the step size after the Adam step size. The calculation for the update of the accumulated action is done as:
+In some domains we used an exponential decay on the step size after the Adam step size. The calculation for the update of the accumulated action is done as:
 δAdam=
@@
 […]
-$\lambda=\max\{$0.999$^{T},$0.1$\}a^{acc}=a+\lambda\delta_{Adam},$ (107)
+$\lambda=\max\{$0.999$^{T},$0.1$\},$ (152)
+$a^{acc}=a+\lambda\delta_{Adam},$ (153)
 where $T$ is the number of gradient optimizations performed at this particular action node.
-B.2 D-Continuous Light Dark
+Since action updates are more costly than computing the gradient estimates, especially when not performing a linearized update of the importance weights, we set a threshold action distance for action update in some domains.
+After each gradient iteration, we update an accumulator of the gradient steps $a_{t}^{acc}$.
+Only after passing a threshold distance $\lVert{a^{acc}-a}\rVert>T_{d_{a}}^{\min}$, we update $a\leftarrow a^{acc}$.
+This limits the number of action updates, and acts like a convergence criterion.
+To ensure controllable step-sizes, we also set a max-norm threshold on the final action update $\lVert{a^{\prime}-a}\rVert<T_{d_{a}}^{\max}$.
+For the Adam optimizer, we used the standard hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$.
+The learning rate $\eta_{Adam}$, alongside $T_{d_{a}}^{\min}$ in domains it was not set to 00, were determined by cross entropy optimization (CE) described later, in conjunction with the optimization over the other algorithm hyperparameters.
+B.4 Thresholds For Adding/Deleting State Nodes
+To save computational overhead, we would like to ensure that we have relevant samples after several action updates.
+As a relevance heuristic, we delete child state $s^{\prime}$ from $C(s,a)$ if $\omega_{a}(s^{\prime})<T_{\omega}^{del}$ after action update.
+Additionally, if $\omega_{a}(s^{\prime})<T_{\omega}^{add}$ for all $s^{\prime}\in C(s,a)$, we force to sample a new posterior node, overriding the progressive widening limitation.
+This ensures that we do not waste computational resources on states with low contribution to the estimator $\hat{Q}(s,a)$.
+Due to the possibility of deleting posterior children, care must be taken when updating visitation counters since they do not monotonically increment anymore.
+For this reason, equations (13), (14), (11), (15), (16) are written using previous and updated visitation counts.
+Appendix C Experiments
+C.1 Experimental Method
+Table 1: Hyperparameters optimized by CE in the evaluations of each algorithm in the D-Continuous Light-Dark, Mountain Car, Hill Car and Lunar Lander domains.
+Parameter Notation Explanation Algorithms Tuning
+$c$ UCT exploration bonus All CE
+$k_{a}$ Action prog. widening factor All CE
+$\alpha_{a}$ Action prog. widening power All CE
+$k_{o}$ Obs. prog. widening factor All CE
+$\alpha_{o}$ Obs. prog. widening power All CE
+$\eta_{Adam}$ Adam step size (learning rate) AGMCTS CE
+$T_{d_{a}}^{\min}$ Dist. threshold AGMCTS Manual/CE
+$J$ No. particles AGMCTS/(PFT-)DPW Manual
+$K_{rollout}$ No. states for rollout AGMCTS/(PFT-)DPW Manual
+$K_{opt}$ No. grad. iterations AGMCTS Manual
+$T_{d_{a}}^{\max}$ Max. action update step AGMCTS Manual
+$T_{\omega}^{add}$ Imp. ratio add threshold AGMCTS Manual
+$T_{\omega}^{del}$ Imp. ratio delete threshold AGMCTS Manual
+$K_{\text{child}}^{\min}$ Action update min. children AGMCTS Manual
+$K_{\text{child}}^{visits}$ Action update every mod $\text{mod }k$ AGMCTS Manual
+$K_{b}^{\nabla}$ No. states for ^$\hat{\nabla}_{a}\log p_{T}$ AGMCTS Manual
+$K_{O}^{\nabla}$ No. obs. branches for ^$\hat{\nabla}_{a}\hat{Q}$ AGMCTS Manual
+$n^{sims}$ Simulation budget All Manual
+The different hyperparameters and their explanations defined by each algorithm are summarized in Table 1.
+In high-level, the experiments were conducted for each domain in the following manner:
+  1. 1.
+    Constant hyperparameters were defined for each algorithm, including simulation budget $n^{sims}$, particle count $J$, and rollout policy.
+  2. 2.
+    A cross-entropy (CE) optimization was performed over the UCB, DPW and step size (where applicable) for each algorithm.
+  3. 3.
+    For the best performing hyperparameters of each algorithm, we ran 1000 simulations of the algorithm with the same seeds between algorithms, at the multiples $[10^{-1},10^{-0.75},10^{-0.5},10^{-0.25},1.0]$ of the simulation budget $n^{sims}$.
+The simulation jobs were run on a distributed cluster of 8 machines with different cores and configurations, using the Julia programming language with Distributed.jl multi-processing tools.
+Each simulation was run on a single core, without any parallelization in the solver itself.
+The processors used were:
+  1. 1.
+    Ryzen 7 3700X 8-Core (16 threads) Processor.
+  2. 2.
+    Intel i7-6820HQ 4-Core (8 threads) Processor.
+  3. 3.
+    Intel i7-8550U 4-Core (8 threads) Processor.
+  4. 4.
+    Intel E3-1241 v3 4-Core (8 threads) Processor.
+  5. 5.
+    Intel E5-1620 v4 4-Core (8 threads) Processor.
+  6. 6.
+    Intel i9-7920X 12-Core (24 threads) Processor.
+  7. 7.
+    Intel i9-11900K 8-Core (16 threads) Processor.
+  8. 8.
+    Intel i9-12900K 12-Core (24 threads) Processor.
+C.1.1 Simulation Budget
+Since POMCPOW is a state-trajectory algorithm, in the POMDP scenarios we measured empirically the runtime of PFT-DPW at $n^{sims}$ and $J$, and set a larger simulation budget for POMCPOW $n^{sims,\text{POMCPOW}}$ to match the runtime of PFT-DPW.
+C.1.2 Rollout Policy
+In order to save computation time and not compute a partially observable rollout, the rollout computation for a new particle belief was done by drawing $K_{rollout}$ particles from the belief, and computing the rollout policy based on the mean state, while calculating the rollout estimate based on the mean of the returns of the $K_{rollout}$ particles.
+For POMCPOW, since every new node is expanded initially only with a single particle, the rollout was computed with that particle.
+This is a disadvantage for POMCPOW as it made the rollout returns less meaningful, however it seemed to have achieved consistently better results than PFT-DPW despite that.
+C.1.3 Cross-Entropy (CE) Optimization
+The algorithm hyperparameters were first determined by cross entropy optimization (CE) [40], largely following [41], using Gaussian distributions over the continuous DPW hyperparameters:
+  1. 1.
+    UCT exploration bonus $c$.
+  2. 2.
+    Action progressive widening parameters $k_{a}$ and $\alpha_{a}$.
+  3. 3.
+    Observation progressive widening parameters $k_{o}$ and $\alpha_{o}$.
+For AGMCTS, we also optimized over the Adam learning rate $\eta_{Adam}$, and the action update distance threshold $T_{d_{a}}^{\min}$ only in the Light-Dark domains.
+We used 150 parameter samples, 40 simulations for each parameter sample to determine its mean return, used 30 elite samples to fit the new distribution, with smoothing the new distribution’s parameters:
+$\mu_{\text{new}}=\alpha_{\mu}\mu_{\text{new}}+(1-\alpha_{\mu})\mu_{\text{old}},$ (154)
+$\Sigma_{\text{new}}=\alpha_{\Sigma}\Sigma_{\text{new}}+(1-\alpha_{\Sigma})\Sigma_{\text{old}},$ (155)
+where we chose $\alpha_{\mu}=0.8,\alpha_{\Sigma}=0.5$.
+When fine-tuning parameters based on previous runs, we used the previous mean and covariance as the starting point for the new CE optimization, with 80 parameter samples and 20 elite samples.
+We performed 50 CE iterations for each algorithm in each dimension, with reasonable initial guesses for the starting parameters based on the literature and manual experiments.
+For some domains, we tuned the parameters of AGMCTS based on the best performing parameters of PFT-DPW, and kept the shared parameters constant throughtout the CE iterations of AGMCTS to limit it search space.
+The hyperparameters chosen were those from the CE iteration achieving the maximal mean of mean returns over the 30 elite samples.
+After choosing the parameters, we reported the mean return over 1000 scenarios, with the standard error of the mean.
+C.2 D-Continuous Light-Dark
+(a) Middle of scenario, $t=5$.
+(b) End of scenario, $t=6$.
+Figure 4: AGMCTS at 2-Continuous Light-Dark.
+The agent’s current state is the red dot, the current belief particles in orange, the next state in black, and the next belief particles in gray, the next observation is the green dot.
+The goal is the blue ring centered at $(0,2.5)$, $b_{0}$ is the red ring centered at $(0,0)$, and the beacon is the green star at $(2.5,0)$.
+The reward function for the posterior state is the blue-red heatmap drawn in the background.
+The shaded area is where the observation noise is $\sigma_{O}>5$.
+The chosen action branch of AGMCTS is drawn as the blue arrows, getting darker in hue with each action update.
+We can see that the optimized actions point closer to the goal from the current belief’s mean.
+C.2.1 Problem Description
 In this domain, $\mathcal{X}=\mathcal{Z}=\mathbb{R}^{D}$.
@@
 The rollout heads directly to $x_{g}$, but the computed action has an added noise of $\mathcal{N}(\mathbf{0},\sigma_{r}^{2}\mathbf{I})$.
-In our experiments we used $\sigma_{r}=0.1$ for parameter tuning, and additionally evaluated for $\sigma_{r}=0.2,0.3$.
-The implementation of D-Continuous Light Dark was written using POMDPs.jl [?].
-The gradient function $\nabla_{a}r$ was equal 00 in our setting, and $\nabla_{a}\log p_{T}$ was calculated using Enzyme.jl [?].
+In our experiments we set $\sigma_{r}=0.1$.
+The implementation of D-Continuous Light-Dark was written using POMDPs.jl [32].
+The gradient function $\nabla_{a}r$ was equal to 00 in our setting, and $\nabla_{a}\log p_{T}$ was calculated using Enzyme.jl [33].
+In this domain, it is straightforward to compute the transition density:
+$p_{T}(x^{\prime}\mid x,a)=\frac{1}{(2\pi\sigma_{T}^{2})^{D/2}}\exp\left(-\frac{\lVert{x^{\prime}-(x+a)}\rVert^{2}}{2\sigma_{T}^{2}}\right).$ (156)
 Enzyme.jl provided extremely efficient gradients, achieving the execution times almost as close to the original functions.
-We did not care to optimize for making primal and dual calls, which might have allowed to save some computation time.
 However, it seems that it would have had a relatively minor effect as the heaviest computation was calculating the belief transition importance ratio, which due to it being a ratio of two products, cannot be MC approximated without introducing bias.
@@
 Yet, in more complex problem where the dynamics are more complex, it might prove beneficial to compute gradients in the same pass as the transition likelihood wherever it can be done.
-B.3 Evaluations
-Parameter Notation Explanation
-$c$ UCT exploration bonus
-$k_{a}$ Action progressive widening (factor)
-$\alpha_{a}$ Action progressive widening (power)
-$k_{o}$ Observation progressive widening (factor)
-$\alpha_{o}$ Observation progressive widening (power)
-$\eta_{Adam}$ Adam step size (i.e. learning rate)
-$T_{d_{a}}$ Action update distance threshold
-Table 2: Hyperparameters optimized by CE in the evaluations of each algorithm for the D-Continuous Light Dark domain, for $D$ dimensions.
+C.2.2 Evaluation
+Table 2: Hyperparameters used in the evaluations of each algorithm for the D-Continuous Light-Dark domain, for $D=2,3,4$ dimensions, rounded to $3$ significant digits after the decimal point.
 Parameter D POMCPOW PFT-DPW AGMCTS
@@
 […]
-Table 3: Hyperparameters used in the evaluations of each algorithm for the D-Continuous Light Dark domain, for $D$ dimensions, rounded to $3$ significant digits after the decimal point.
-We evaluated the algorithms on the D-Continuous Light Dark domain for $D=2,3,4$ dimensions.
-PFT-DPW and AGMCTS were given a budget of 500 simulations.
+We evaluated the algorithms on the D-Continuous Light-Dark domain for $D=2,3,4$ dimensions.
+PFT-DPW and AGMCTS were given a budget of $n^{sims}=500$ simulations.
 The particle count during planning was set to be dependent on the dimension:
-$J_{D=2}=256,\quad J_{D=3}=512,$
-$J_{D=4}=1024.$ (108)
+$J_{D=2}=256,\quad J_{D=3}=512,\quad J_{D=4}=1024.$ (157)
 The particle filter used for inference between planning steps had a particle count $J^{\text{PF}}$ given by:
-$J^{\text{PF}}_{D=2}=2048,\quad J^{\text{PF}}_{D=3}=4096,$
-$J^{\text{PF}}_{D=4}=8192.$ (109)
+$J^{\text{PF}}_{D=2}=2048,\quad J^{\text{PF}}_{D=3}=4096,\quad J^{\text{PF}}_{D=4}=8192.$ (158)
 POMCPOW was given a simulation budget according to the formula
-$n_{D}^{sims,POMCPOW}=500\cdot$0.08$\cdot J_{D},$ (110)
-which was empirically found to produce similar runtimes to PFT-DPW in the chosen domain.
-In order to save computation time and not compute a partially observable rollout, the rollout computation for a new particle belief was done by drawing $s_{0}$ and 10 additional particles from the belief.
-The rollout actions were computed based on $s_{0}$, but the return was the mean of returns from the 10 particles.
-This mechanism allows to approximate a partially observable rollout without receiving any future observations.
-For POMCPOW, since every new node is expanded initially only with a single particle, the rollout was computed with that particle.
-This is a disadvantage for POMCPOW as it made the rollout returns less meaningful, however it seemd to have achieved better results than PFT-DPW despite that.
-The algorithm hyperparameters were first determined by cross entropy optimization (CE) [?], largely following [?], using Gaussian distributions over the continuous DPW hyperparameters:
-  1. 1.
-    UCT exploration bonus $c$.
-  2. 2.
-    Action progressive widening parameters $k_{a}$ and $\alpha_{a}$.
-  3. 3.
-    Observation progressive widening parameters $k_{o}$ and $\alpha_{o}$.
-For AGMCTS, we also optimized over the Adam learning rate $\eta_{Adam}$, and the action update distance threshold $T_{d_{a}}$.
-We searched for hyperparameters for each dimension while using rollout policy noise $\sigma_{r}=0.1$, and the results for $\sigma_{r}=0.2,0.3$ were obtained using the same hyperparameters of the same dimension.
-We used 150 parameter samples, 40 simulations for each parameter sample to determine its mean return, used 30 elite samples to fit the new distribution, with smoothing the new distribution’s parameters:
-$\mu_{\text{new}}=\alpha_{\mu}\mu_{\text{new}}+(1-\alpha_{\mu})\mu_{\text{old}},$ (111)
-$\Sigma_{\text{new}}=\alpha_{\Sigma}\Sigma_{\text{new}}+(1-\alpha_{\Sigma})\Sigma_{\text{old}},$ (112)
-where we chose $\alpha_{\mu}=0.8,\alpha_{\Sigma}=0.5$.
-We performed 50 CE iterations for each algorithm in each dimension.
-The starting mean for $D=2$ was the same for all algorithms, and for $D=3,4$ we used the mean of the best hyperparameters found for $D-1$.
-The hyperparameters chosen were those from the CE iteration achieving the maximal mean of mean returns over the 30 elite samples.
-The parameters used for each algorithm are summarized in Table 3.
-After choosing the parameters, we reported the mean return over 1000 scenarios, with the standard error of the mean.
-Table 4 summarizes the results with the individual mean runtimes for each algorithm in each scenario.
-D $\sigma_{r}$ POMCPOW time (s) PFT-DPW time (s) AGMCTS time (s)
-2 0.1 $\boldsymbol{6.62}\pm 0.07$ 0.114 $5.71\pm 0.08$ 0.126 $\boldsymbol{6.67}\pm 0.06$ 1.665
-2 0.2 $\boldsymbol{6.23}\pm 0.07$ 0.122 $4.74,0.08$ 0.129 $5.96\pm 0.08$ 1.730
-2 0.3 $\boldsymbol{5.61}\pm 0.08$ 0.129 $4.01\pm 0.09$ 0.129 $\boldsymbol{5.49}\pm 0.09$ 1.838
-3 0.1 $3.42\pm 0.08$ 0.216 $3.63\pm 0.08$ 0.299 $\boldsymbol{5.47}\pm 0.07$ 2.442
-3 0.2 $1.87\pm 0.08$ 0.240 $1.79\pm 0.08$ 0.296 $\boldsymbol{3.82}\pm 0.1$ 2.416
-3 0.3 $1.06\pm 0.07$ 0.253 $0.29\pm 0.06$ 0.309 $\boldsymbol{1.58}\pm 0.1$ 2.389
-4 0.1 $2.58\pm 0.07$ 0.548 $2.36\pm 0.07$ 0.683 $\boldsymbol{4.26},0.08$ 10.576
-4 0.2 $0.53\pm 0.05$ 0.640 $0.2\pm 0.05$ 0.707 $\boldsymbol{1.53}\pm 0.09$ 8.303
-4 0.3 $\boldsymbol{-0.36}\pm 0.03$ 0.660 $-0.86\pm 0.02$ 0.736 $-1.13\pm 0.04$ 5.778
-Table 4: Performance results for the different algorithms at D-Continuous Light Dark for $D=2,3,4$ dimensions.
-AGMCTS and PFT-DPW were run for 500 simulations with particle counts given by (108), and POMCPOW for an approximately equivalent time to PFT-DPW according to (110).
-The rollout policy’s standard deviation was $\sigma_{r}=0.1$ during hyperparameter tuning, and then the algorithms were also tested for $\sigma_{r}=0.2,0.3$.
-The reported results are mean returns over 1000 scenarios $\pm$ standard error.
-AGMCTS achieved the best performance in most scenarios.
+$n_{pomcpow,D}^{sims}=500\cdot$0.08$\cdot J_{D}.$ (159)
+Additional parameters that were manually set in all dimensions were:
+[…]
+$T_{d_{a}}^{\max}=\infty,\quad K_{\text{child}}^{\min}=1,\quad K_{\text{child}}^{visits}=1,\quad K_{b}^{\nabla}=5.$ (160)
+In this domain, exponential step size decay was used.
+Table 3: Performance and mean time results at 2D-Continuous Light-Dark.
+$n^{sims}_{pomcpow}$ POMCPOW time(s) $n^{sims}$ PFT-DPW time(s) AGMCTS time(s)
+1024 $3.60\pm 0.11$ 0.018 50 $3.94\pm 0.09$ 0.014 $\boldsymbol{4.27}\pm 0.09$ 0.018
+1823 $4.63\pm 0.10$ 0.016 89 $4.62\pm 0.09$ 0.022 $\boldsymbol{5.28}\pm 0.09$ 0.041
+3236 $5.21\pm 0.10$ 0.033 158 $5.06\pm 0.09$ 0.041 $\boldsymbol{6.06}\pm 0.08$ 0.201
+5755 $\boldsymbol{6.28}\pm 0.08$ 0.058 281 $5.52\pm 0.08$ 0.073 $\boldsymbol{6.26}\pm 0.07$ 0.593
+10240 $\boldsymbol{6.62}\pm 0.07$ 0.114 500 $5.71\pm 0.08$ 0.126 $\boldsymbol{6.67}\pm 0.06$ 1.665
+Table 4: Performance and mean time results at 3D-Continuous Light-Dark.
+$n^{sims}_{pomcpow}$ POMCPOW time(s) $n^{sims}$ PFT-DPW time(s) AGMCTS time(s)
+2048 $\boldsymbol{2.37}\pm 0.09$ 0.025 50 $\boldsymbol{2.37}\pm 0.08$ 0.031 $2.01\pm 0.08$ 0.328
+3645 $2.36\pm 0.09$ 0.036 89 $\boldsymbol{2.78}\pm 0.08$ 0.052 $\boldsymbol{2.81}\pm 0.10$ 0.162
+6472 $\boldsymbol{3.39}\pm 0.09$ 0.065 158 $3.12\pm 0.08$ 0.089 $3.09\pm 0.10$ 0.432
+11510 $3.60\pm 0.08$ 0.116 281 $3.44\pm 0.08$ 0.164 $\boldsymbol{5.46}\pm 0.07$ 1.064
+20480 $3.42\pm 0.08$ 0.216 500 $3.63\pm 0.08$ 0.299 $\boldsymbol{5.47}\pm 0.07$ 2.442
+Table 5: Performance and mean time results at 4D-Continuous Light-Dark.
+$n^{sims}_{pomcpow}$ POMCPOW time(s) $n^{sims}$ PFT-DPW time(s) AGMCTS time(s)
+4096 $\boldsymbol{1.48}\pm 0.07$ 0.052 50 $0.85\pm 0.07$ 0.065 $0.65\pm 0.06$ 0.132
+7291 $\boldsymbol{2.13}\pm 0.07$ 0.089 89 $1.47\pm 0.07$ 0.116 $1.34\pm 0.09$ 0.318
+12943 $\boldsymbol{2.15}\pm 0.08$ 0.157 158 $1.98\pm 0.07$ 0.213 $1.79\pm 0.09$ 1.031
+23020 $2.44\pm 0.07$ 0.286 281 $2.13\pm 0.07$ 0.382 $\boldsymbol{3.75}\pm 0.09$ 3.024
+40960 $2.58\pm 0.07$ 0.548 500 $2.36\pm 0.07$ 0.683 $\boldsymbol{4.26}\pm 0.08$ 10.57
+C.3 Mountain Car and Hill Car
+C.3.1 Problem Description
+A car is placed randomly in a valley between two hills, and the goal is to reach the top of the right hill.
+The car’s action is its acceleration left or right, and must gain momentum in order to reach the goal.
+In this domain, the state $s=(x,v)\in\mathbb{R}^{2}$ is the car’s position and velocity, $\mathcal{A}=[-a_{\max},a_{\max}]\subset\mathbb{R}$ is the action space.
+In the POMDP version, $z\in\mathbb{R}$ is the observation of the car’s position with noise. It is given by $z\sim\mathcal{N}(x,\sigma_{O}^{2})$, where $\sigma_{O}=0.03$.
+The car’s position is bounded to $[x_{\min},x_{\max}]$, and the velocity is bounded to $[-v_{\max},v_{\max}]$.
+If the car reaches the goal $x\geq x_{\max}$, it receives a reward of $100$.
+If $x<x_{\min}$, or $\left|{v}\right|\geq v_{\max}$, a penalty reward of $-100$ is attained and the scenario terminates.
+Otherwise, the car obtains a reward of $-0.1$ at each time step that does not terminate.
+The reward function is a function of the posterior state, i.e.
+$r(s,a,s^{\prime})=r(s^{\prime})$.
+In Mountain Car, the scenario terminates after $L=200$ time steps, while in Hill Car it terminates after $L=30$ time steps.
+The discount in both domains is $\gamma=0.99$.
+The rollout policy is calculated by choosing $a_{\max}$ when $v>0$ and $-a_{\max}$ otherwise.
+In both domains, the generative transition model is given by an input action noise to the deterministic dynamics:
+$\nu\sim\mathcal{N}(0,\sigma_{\nu}^{2}),$ (161)
+$\tilde{a}=\text{clip}(a+w,-a_{\max},a_{\max}),$ (162)
+$s^{\prime}=f(s,\tilde{a},s^{\prime}),$ (163)
+where $\sigma_{\nu}=0.1$ for both domains.
+When calculating the transition density in this domain, we do it based on the input noise to a deterministic simulator case via the Area Formula.
+For each $(s,a^{\prime},s^{\prime})$ tuple we calculate the generating action noise $\nu$, depending on the scenario.
+If the obtained $\nu$ is outside of the bounds (beyond numerical errors) $[L,R]$ where $L=-a_{\max}-a$ and $R=a_{\max}-a$, then the density is 00.
+Otherwise, because of the clip operation, we calculate its density by the following mixture:
+$p(\nu\mid s,a)=\begin{cases}CDF(\mathcal{N}(0,\sigma_{\nu}^{2}),L),&\text{if }\nu\leq L\\ 1-CDF(\mathcal{N}(0,\sigma_{\nu}^{2}),R),&\text{if }\nu\geq R\\ PDF(\mathcal{N}(0,\sigma_{\nu}^{2}),\nu),&\text{otherwise}.\end{cases}$ (164)
+Afterwards, we calculate the transition model density by the Area Formula according to the case of input noise to a deterministic simulator.
+C.3.2 Mountain Car Transition Model
+In this domain, we set $x_{\min}=-1.5$, $x_{\max}=0.5$, $v_{\max}=0.05$, $a_{\max}=1.0$. The deterministic dynamics are calculated by:
+$v^{\prime}=v+0.001\cdot\tilde{a}-0.0025\cos(3x),$ (165)
+$x^{\prime}=x+v^{\prime}.$ (166)
+In order to calculate the transition density, for given $s,a^{\prime},s^{\prime}$ we invert the dynamics to obtain $\nu$.
+C.3.3 Hill Car Transition Model
+In this domain, we set $x_{\min}=-1.0$, $x_{\max}=1.0$, $v_{\max}=2.5$, $a_{\max}=4.0$.
+The deterministic dynamics are calculated by integrating the car’s continuous-time dynamics with an ODE solver over the time period $t_{0}$ to $t_{1}$.
+To limit the computational complexity, we chose crudely $\delta t=0.01$ for the integration, while $t_{1}-t_{0}=0.1$ was the time step between consecutive time steps of the MDP/POMDP scenario.
+The integration method chosen was Tsitouras 5/4 Runge-Kutta method.
+The continuous-time dynamics are given by:
+$\dot{v}=\left(\frac{a}{m}-g\cdot h^{\prime}(x)-v^{2}\cdot h^{\prime}(x)\cdot h^{\prime\prime}(x)\right)\frac{1}{1+h^{\prime}(x)^{2}},$ (167)
+$\dot{x}=v,$ (168)
+where $m=1$, $g=9.81$, and the hill-curve function $h(x)$ is given by:
+$h(x)=\begin{cases}x^{2}+x,&\text{if }x<0\\ \frac{x}{\sqrt{1.0+5.0*x^{2}}},&\text{otherwise}.\end{cases}$ (169)
+Because the input noise is only based on the action, we calculate $\nu$ by caching simulator outputs $(s,a,\nu,s^{\prime})$ tuples for which $s^{\prime}=f(s,\tilde{a})$.
+For a new tuple $(s,a^{\prime},s^{\prime})$, we calculate $\nu^{\prime}=\tilde{a}-a^{\prime}$ for the cached $\tilde{a}$.
+The jacobian of the dynamics is calculated by ForwardDiff.jl and SciMLSensitivity.jl overloaded gradient function for the ODE solver.
+C.3.4 Evaluation
+DPW, PFT-DPW and AGMCTS were given a simulation budget of $n^{sims}=500$ simulations.
+In the POMDP settings, the particle count was set to $J=30$, and we used $J^{PF}=200$ for inference between planning steps.
+POMCPOW simulation budget was set by the formula:
+$n_{pomcpow}^{sims}=500\cdot$0.08$\cdot J.$ (170)
+In this domain, we used linearized weight updates for AGMCTS MIS action update. Exponential step size decay was not used.
+Table 6: Hyperparameters used in evaluations of AGMCTS in the Mountain/Hill Car domains.
+Parameter Mountain/MDP Mountain/POMDP Hill/MDP Hill/POMDP
+$c$ 0.0 0.001 132.24 132.83
+$k_{a}$ 6.876 4.558 6.552 8.657
+$\alpha_{a}$ 0.619 0.698 0.532 0.490
+$k_{o}$ 0.292 0.379 5.375 6.050
+$\alpha_{o}$ 0.385 0.382 0.203 0.130
+[…]
+$T_{d_{a}}^{\min}$ 0.0 0.0 0.0 0.0
+$J$ 1 30 1 30
+$K_{rollout}$ 1 5 1 5
+$K_{opt}$ 3 3 3 3
+$T_{d_{a}}^{\max}$ 0.1 0.1 0.1 0.1
+$T_{\omega}^{add}$ 1.0 0.99 1.0 0.99
+[…]
+$K_{\text{child}}^{\min}$ 2 2 2 2
+$K_{\text{child}}^{visits}$ 1 1 1 1
+$K_{b}^{\nabla}$ 3 - 3
+$K_{O}^{\nabla}$ linearized linearized linearized linearized
+$n^{sims}$ 500 500 500 500
+Table 7: Hyperparameters used in evaluations of DPW/PFT-DPW in the Mountain/Hill Car domains.
+Parameter Mountain/MDP Mountain/POMDP Hill/MDP Hill/POMDP
+$c$ 92.148 146.08 132.24 119.28
+$k_{a}$ 6.672 5.625 6.552 7.386
+$\alpha_{a}$ 0.581 0.824 0.532 0.528
+$k_{o}$ 0.277 1.049 5.375 1.256
+$\alpha_{o}$ 0.454 0.415 0.203 0.588
+$J$ 1 30 1 30
+$K_{rollout}$ 1 5 1 5
+$n^{sims}$ 500 500 500 500
+Table 8: Hyperparameters used in evaluations of POMCPOW in the Mountain/Hill Car domains.
+Parameter Mountain/POMDP Hill/POMDP
+$c$ 59.585 90.960
+$k_{a}$ 4.082 6.944
+$\alpha_{a}$ 0.640 0.209
+$k_{o}$ 0.520 5.052
+$\alpha_{o}$ 0.197 0.845
+$n^{sims}$ 2739 2739
+Table 9: Performance and mean time results at Mountain Car MDP.
+$n^{sims}$ DPW time(s) AGMCTS time(s)
+50 $\boldsymbol{-52.00}\pm 0.17$ 0.001 $\boldsymbol{-52.00}\pm 0.17$ 0.0011
+89 $\boldsymbol{-51.66}\pm 0.25$ 0.002 $\boldsymbol{-51.66}\pm 0.25$ 0.0030
+158 $\boldsymbol{-48.00}\pm 0.58$ 0.004 $-51.05\pm 0.31$ 0.0053
+281 $-45.43\pm 0.66$ 0.008 $\boldsymbol{29.43}\pm 0.07$ 0.0378
+500 $24.94\pm 0.30$ 0.012 $\boldsymbol{29.52}\pm 0.066$ 0.0865
+Table 10: Performance and mean time results at Mountain Car POMDP.
+$n^{sims}_{pomcpow}$ POMCPOW time(s) $n^{sims}$ PFT-DPW time(s) AGMCTS time(s)
+274 $24.52\pm 0.33$ 0.007 50 $\boldsymbol{24.95}\pm 0.33$ 0.002 $\boldsymbol{25.29}\pm 0.32$ 0.003
+487 $\boldsymbol{25.26}\pm 0.32$ 0.014 89 $24.45\pm 0.37$ 0.005 $24.11\pm 0.42$ 0.006
+865 $\boldsymbol{25.78}\pm 0.28$ 0.025 158 $23.85\pm 0.41$ 0.009 $24.45\pm 0.41$ 0.011
+1539 $25.00\pm 0.36$ 0.044 281 $23.29\pm 0.44$ 0.017 $\boldsymbol{26.46}\pm 0.24$ 0.022
+2739 $24.24\pm 0.43$ 0.080 500 $23.16\pm 0.44$ 0.030 $\boldsymbol{26.65}\pm 0.19$ 0.040
+Table 11: Performance and mean time results at Hill Car MDP.
+$n^{sims}$ DPW time(s) AGMCTS time(s)
+50 $\boldsymbol{-92.87}\pm 1.05$ 0.029 $\boldsymbol{-92.87}\pm 0.09$ 0.003
+89 $\boldsymbol{-90.11}\pm 1.23$ 0.005 $\boldsymbol{-90.44}\pm 0.09$ 0.018
+158 $-92.22\pm 1.10$ 0.009 $\boldsymbol{-90.60}\pm 0.08$ 0.016
+281 $-89.30\pm 1.27$ 0.015 $\boldsymbol{17.25}\pm 0.07$ 0.060
+500 $-88.66\pm 1.31$ 0.029 $\boldsymbol{51.82}\pm 0.06$ 0.260
+Table 12: Performance and mean time results at Hill Car POMDP.
+$n^{sims}_{pomcpow}$ POMCPOW time(s) $n^{sims}$ PFT-DPW time(s) AGMCTS time(s)
+274 $\boldsymbol{52.54}\pm 1.09$ 0.031 50 $-82.91\pm 1.56$ 0.030 $-77.49\pm 1.76$ 0.055
+487 $\boldsymbol{53.98}\pm 1.00$ 0.054 89 $-75.58\pm 1.83$ 0.055 $-69.48\pm 1.99$ 0.063
+865 $\boldsymbol{54.15}\pm 1.0$ 0.097 158 $-65.22\pm 2.09$ 0.098 $-67.16\pm 2.05$ 0.106
+1539 $\boldsymbol{54.55}\pm 0.98$ 0.178 281 $-64.86\pm 2.10$ 0.183 $3.75\pm 2.42$ 0.227
+2739 $\boldsymbol{55.65}\pm 0.90$ 0.311 500 $-36.53\pm 2.46$ 0.329 $51.79\pm 1.12$ 0.641
+C.4 Lunar Lander
+C.4.1 Problem Description
+For this domain, we took the Julia implementation provided by Mern et al. [8].
+The vehicle state is represented by a six dimensional tuple $(x,y,\theta,\dot{x},\dot{y},\omega)$, where $x$ and $y$ are the horizontal and vertical positions, $\theta$ is the orientation angle, $\dot{x}$ and $\dot{y}$ are the horizontal and vertical speeds, and $\omega$ is the angular rate.
+The action space $\mathcal{A}\subset\mathbb{R}^{3}$ is a three-dimensional continuous space defined by the tuple $(F_{x},T,\delta)$.
+$T$ is the main thrust which acts along the vehicle’s vertical axis through its center of mass, and is in the range $[0,15]$.
+$F_{x}$ is the corrective thrust, which acts along a horizontal axis offset from the center of mass by a distance $\delta$.
+$F_{x}$ is in the range $[-5,5]$ and $\delta\in[-1,1]$.
+The transition model is computed as an additive Gaussian noise to the output of a deterministic transition function.
+The deterministic transition function is given by the following set of computations:
+$f_{x}=\cos(\theta)\cdot F_{x}-\sin(\theta)\cdot T,$ (171)
+$f_{z}=\cos(\theta)\cdot T+\sin(\theta)\cdot F_{x},$ (172)
+$\tau=-\delta\cdot F_{x},$ (173)
+$a_{x}=\frac{f_{x}}{m},\quad a_{z}=\frac{f_{z}}{m},\quad\dot{\omega}=\frac{\tau}{I},$ (174)
+$v_{x}^{\prime}=v_{x}+a_{x}\cdot\Delta t+\epsilon_{1}\cdot Q_{4},$ (175)
+$v_{z}^{\prime}=v_{z}+(a_{z}-9.0)\cdot\Delta t+\epsilon_{2}\cdot Q_{5},$ (176)
+$\omega^{\prime}=\omega+\dot{\omega}\cdot\Delta t+\epsilon_{3}\cdot Q_{6},$ (177)
+$x^{\prime}=x+v_{x}\cdot\Delta t,\quad z^{\prime}=z+v_{z}\cdot\Delta t,\quad\theta^{\prime}=\theta+\omega\cdot\Delta t,$ (178)
+$s^{\prime}=\begin{bmatrix}x^{\prime}\\ z^{\prime}\\ \theta^{\prime}\\ v_{x}^{\prime}\\ v_{z}^{\prime}\\ \omega^{\prime}\end{bmatrix},$ (179)
+where the noise variables $\epsilon_{i}$ are sampled from a standard normal distribution, $Q_{4}=0.1$, $Q_{5}=0.1$, $Q_{6}=0.01$.
+We took the values $m=1$, $I=10$, and $\delta t=0.4$.
+For computing the transition density, we used the Area Formula by solving the dynamics for the noise variables $\epsilon_{i}$ for given $(s,a^{\prime},s^{\prime})$ tuples.
+In the POMDP variant, the vehicle makes noisy observations of its angular rate, horizontal speed, and above-ground height as measured by a distance sensor.
+The noise is Gaussian with a diagonal covariance matrix $\Sigma_{O}=\text{diag}(1.0^{2},0.01^{2},0.1^{2})$.
+The reward function is defined as:
+$r(s,a,s^{\prime})=\begin{cases}-1000,&\text{if }x\geq 15\vee\theta\geq 0.5\\ 100-x-\dot{y}^{2},&\text{if }y\leq 1\\ -1,&\text{otherwise}\end{cases}$ (180)
+The initial vehicle state is sampled from a multivariate Gaussian with mean $\mu=(x=0,y=50,\theta=0,\dot{x}=0,\dot{y}=-10,\omega=0)$.
+The rollout policy was computed as a simple proportional rule: $a_{rollout}=(-0.1\cdot\dot{x},-0.1\cdot\dot{y},0)$.
+C.4.2 Evaluation
+DPW, PFT-DPW and AGMCTS were given a simulation budget of $n^{sims}=1000$ simulations.
+In the POMDP settings, the particle count was set to $J=150$, and we used $J^{PF}=2000$ for inference between planning steps.
+POMCPOW simulation budget was set by the formula:
+$n_{pomcpow}^{sims}=1000\cdot$0.035$\cdot J.$ (181)
+In this domain, we used linearized weight updates for AGMCTS MIS action update. Exponential step size decay was not used.
+Specifically in the Lunar Lander MDP/POMDP domains, we observed that the planner was converging to a setting in which the gradients were practically never used.
+We tried repeating the optimization process via a constrained CE, in which rejection-sampling is used to constrain the parameter samples satisfy the following condition:
+$\eta_{Adam}\cdot\frac{K_{opt}}{K_{\text{visits}}}\left(N-\text{ceil}\left(\left(\frac{K_{\text{child}}^{\min}}{k_{o}}\right)^{1/\alpha_{o}}\right)\right)\geq 0.01\cdot T_{d_{a}}^{\max}$ (182)
+It seemed that forcing the planner to use the gradients was not beneficial in this domain, however in the unconstrained CE setting, the hyperparameters converged to a setting in which the gradients were not used, and its performance seemed to match that of PFT-DPW.
+Table 13: Hyperparameters used in evaluations of AGMCTS in the Lunar Lander domains.
+Parameter Lunar Lander MDP Lunar Lander POMDP
+$c$ 61.55 112.14
+$k_{a}$ 3.052 4.846
+$\alpha_{a}$ 0.377 0.320
+$k_{o}$ 0.114 1.347
+$\alpha_{o}$ 0.047 0.273
+[…]
+$T_{d_{a}}^{\min}$ 0.0 0.0
+$J$ 1 150
+$K_{rollout}$ 1 5
+$K_{opt}$ 1 5
+$T_{d_{a}}^{\max}$ 0.1 0.1
+$T_{\omega}^{add}$ 0.9 0.99
+[…]
+$K_{\text{child}}^{\min}$ 1 2
+$K_{\text{child}}^{visits}$ 1 2
+$K_{b}^{\nabla}$ 3
+$K_{O}^{\nabla}$ linearized linearized
+$n^{sims}$ 500 500
+Table 14: Hyperparameters used in evaluations of DPW/PFT-DPW in the Lunar Lander domains.
+Parameter Lunar Lander MDP Lunar Lander POMDP
+$c$ 60.49 60.22
+$k_{a}$ 1.421 2.687
+$\alpha_{a}$ 0.595 0.436
+$k_{o}$ 0.082 0.274
+$\alpha_{o}$ 0.726 0.575
+$J$ 1 150
+$K_{rollout}$ 1 5
+$n^{sims}$ 500 500
+Table 15: Hyperparameters used in evaluations of POMCPOW in the Lunar Lander POMDP domain.
+Parameter Lunar Lander POMDP
+$c$ 66.04
+$k_{a}$ 3.309
+$\alpha_{a}$ 0.393
+$k_{o}$ 0.354
+$\alpha_{o}$ 0.433
+$n^{sims}$ 12247
+Table 16: Performance and mean time results at Lunar Lander MDP.
+$n^{sims}$ DPW time(s) AGMCTS time(s)
+100 $\boldsymbol{43.55}\pm 1.44$ 0.002 $\boldsymbol{44.15}\pm 1.06$ 0.023
+178 $\boldsymbol{50.01}\pm 0.98$ 0.002 $\boldsymbol{48.49}\pm 1.03$ 0.010
+316 $\boldsymbol{56.19}\pm 0.39$ 0.002 $53.88\pm 0.46$ 0.022
+562 $\boldsymbol{58.29}\pm 1.02$ 0.004 $\boldsymbol{57.22}\pm 0.40$ 0.046
+1000 $\boldsymbol{63.09}\pm 0.29$ 0.007 $60.02\pm 0.46$ 0.096
+Table 17: Performance and mean time results at Lunar Lander POMDP.
+$n^{sims}_{pomcpow}$ POMCPOW time(s) $n^{sims}$ PFT-DPW time(s) AGMCTS time(s)
+1225 $\boldsymbol{50.96}\pm 3.02$ 0.015 100 $15.33\pm 5.09$ 0.0082 $-25.86\pm 6.38$ 0.026
+2180 $\boldsymbol{53.06}\pm 2.88$ 0.023 178 $33.12\pm 3.95$ 0.0145 $0.872\pm 5.32$ 0.041
+3870 $\boldsymbol{57.58}\pm 2.30$ 0.042 316 $40.44\pm 3.40$ 0.0262 $25.56\pm 4.12$ 0.078
+6883 $\boldsymbol{58.48}\pm 2.28$ 0.075 562 $41.86\pm 3.55$ 0.0444 $37.68\pm 3.54$ 0.147
+12247 $\boldsymbol{55.19}\pm 2.83$ 0.141 1000 $48.03\pm 2.95$ 0.0764 $42.98\pm 3.54$ 0.298